### v1

# Distributionally Robust Markov Games with Average Reward

Abstract
This paper introduces the formulation of a distributionally robust Markov game (DR-MG) with average rewards, a crucial framework for multi-agent decision-making under uncertainty over extended horizons. Unlike finite-horizon or discounted models, the average-reward criterion naturally captures long-term performance for systems designed for continuous operation, where sustained reliability is paramount. We account for uncertainty in transition kernels, with players aiming to optimize their worst-case average reward. We first establish a connection between the multi-agent and single agent settings, and derive the solvability of the robust Bellman equation under the average-reward formulation. We then rigorously prove the existence of a robust Nash Equilibrium (NE), offering essential theoretical guarantees for system stability. We further develop and analyze an algorithm named robust Nash-Iteration to compute the robust Nash Equilibria among all agents, providing practical tools for identifying optimal strategies in complex, uncertain, and long-running multi-player environments. Finally, we demonstrate the connection between the average-reward NE and the well-studied discounted NEs, showing that the former can be approximated as the discount factor approaches one. Together, these contributions provide a comprehensive theoretical and algorithmic foundation for identifying optimal strategies in complex, uncertain, and long-running multi-player environments, which allow for the future extension of robust average-reward single-agent problems to the multi-agent setting.
1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments Shapley (1953); Filar and Vrieze (2012); Littman (1994). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the "Sim-to-Real" gap McMahan, Artiglio, and Xie (2024), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks Bukharin et al. (2023); Ma et al. (2023). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
To address this vulnerability, the framework of distributionally robust Markov Games (DR-MGs) has been developed Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011), extending the principles of distributionally robust Markov Decision Processes (MDPs) Bagnell, Ng, and Schneider (2001); Nilim and El Ghaoui (2004); Iyengar (2005) to the multi-agent setting. Instead of relying on a single, fixed MG model, the robust approach seeks to find an equilibrium that optimizes the worst-case performance over a predefined uncertainty set of plausible game models under potential model mismatches. The resulting robust equilibrium provides a performance guarantee across all models within this set, thereby ensuring each agent’s resilience against model mismatch.
The nature of a DR-MG problem is fundamentally shaped by the optimality criterion used. Much of the prior work has concentrated on the finite-horizon Ma et al. (2023); Shi et al. (2024a, b); Jiao and Li (2024); Li et al. (2025); Blanchet et al. (2023) or discounted-reward setting Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011), where the objective is to maximize the expected reward over a finite number of steps or a discounted sum of rewards.
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics Krnjaic et al. (2024), communication networks Qu et al. (2020), autonomous vehicle coordination Peake et al. (2020), financial markets Vadori et al. (2024), or even peer-to-peer energy trading Qiu et al. (2021). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate Hernández-Lerma and Lasserre (2000). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium Filar and Vrieze (2012); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions Puterman (1994); Atia et al. (2021). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
Moreover, none of the methods for DR-MGs under the discounted reward or finite horizon can be extended to the average-reward setting. The existence of (non-stationary) Nash Equilibrium under finite horizon is derived through backward induction Blanchet et al. (2023), which does not apply to infinite-horizon settings. The existence of stationary Nash Equilibrium under discounted reward is derived in Kardeş, Ordóñez, and Hall (2011), which however heavily relies on the uniqueness of the solution to the discounted robust Bellman equation Iyengar (2005) and hence cannot be applied to the average reward. Another natural attempt is to use the discounted reward as an intermediate step to study the robust Nash Equilibrium under the average reward by setting the discount factor approaching to $1$. Although this limit method is effective in the single-agent setting Wang et al. (2023a), it cannot solve the DR-MGs, due to the complicated structure (also see our detailed discussion in Section 3). To date, the literature on average-reward DR-MGs remains sparse, leaving fundamental questions about the structure of robust equilibria and the design of convergent solution algorithms largely unanswered.
1.1 Contributions
In this paper, we develop a comprehensive study of DR-MGs under the average-reward formulation. Our contributions are summarized as follows. All of our studies are developed under a standard irreducibility assumption (Assumption 4.1), and can be further extended to unichain settings (see Section F in the Appendix).
Solvability of the Robust Bellman Equation: As a foundational step, we first revisit the single-agent distributionally robust average-reward MDPs. Despite existing studies all utilizing the robust Bellman equation Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025), its solvability is not proved, but rather implicitly assumed. We establish the well-posedness of the robust average-reward Bellman equation, study the structure of its solutions, and connect them with the optimal robust policies. This analysis provides structural insights and tools that are essential for solving multi-agent Markov games when long-term performance is of interest.
Existence and Algorithmic Identification of Robust Nash Equilibrium: We further introduce a formulation of induced distributionally robust MDPs to reduce the Nash Equilibrium identification in comprehensive multi-agent games to optimizing single-agent distributionally robust MDPs. We then utilize the connection revealed to derive the convexity and the semi-continuity of the optimal policy set, based on which we further establish the existence of a stationary Nash Equilibrium. This result addresses a fundamental open question, demonstrating that stable solutions exist in these complex, robust, long-horizon competitive environments.
Convergence Algorithm: We moreover propose a practical, iterative algorithm—a robust Nash iteration scheme—designed to compute the robust NE in the average-reward setting. We provide a rigorous proof of the algorithm’s convergence to the robust equilibrium policies, offering the first provably convergent method for solving average-reward DR-MGs.
Connection to Discounted Equilibria: We also aim to bridge the gap between DR-MGs under the average reward and the well-studied discounted-reward settings. We demonstrate the closeness between the robust average reward NE and robust discounted NEs as the discount factor approaches one. This connection not only provides deeper insight into the nature of the average reward equilibrium, but also suggests potential approximation methods based on the discounted framework.
1.2 Related Works
We discuss the most related works in this section.
Non-robust Markov Games. Markov Games Littman (1994) provide the foundational mathematical framework for multi-agent sequential decision-making. The majority of early work focused on the discounted-reward setting, where the existence of a stationary Nash Equilibrium was established in Shapley (1953). The average-reward criterion, while more suitable for systems with long operational horizons, presents greater analytical challenges. Unlike the discounted case, a stationary NE is not guaranteed to exist in a general average-reward MG. Its existence has only been proven under additional structural assumptions on the dynamics of the game Guo and Yang (2008); Hernández-Lerma and Lasserre (2000); Guo and Hernández-Lerma (2003); Zheng and Guo (2024); Nowak (1999); Iwase, Tanaka, and Wakuta (1976); Küenle and Schurath (2003); Tanaka and Wakuta (1977); Jaśkiewicz and Nowak (2001); Filar and Vrieze (2012).
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, the distributionally robust MDPs are first developed in Iyengar (2005); Nilim and El Ghaoui (2004) for both finite and discounted settings. The studies under the average reward setting are recently developed in Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025); Grand-Clement, Petrik, and Vieille (2023); Grand-Clément and Petrik (2023); Chatterjee et al. (2023). While the robust Bellman equation is central to these methods, its solvability was often implicitly assumed without rigorous proof.
Distributionally Robust Markov Games. Extending distributional robustness to the multi-agent setting is a recent and active area of research. However, the literature has overwhelmingly concentrated on the finite-horizon Ma et al. (2023); Shi et al. (2024a, b); Jiao and Li (2024); Li et al. (2025); Blanchet et al. (2023) and discounted-reward criteria Zhang et al. (2020); Kardeş, Ordóñez, and Hall (2011). These settings are often more analytically tractable because the influence of future rewards diminishes, simplifying the analysis. However, none of these methods can be extended to the average-reward setting, as we shall discuss later.
2 Preliminaries and Problem Formulation
[…]
A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is a distribution over $\mathcal{A}$ for any given state $s$, meaning the agent takes action $a$ at state $s$ with probability $\pi(a|s)$. The goal of an MDP is to maximize the accumulative reward. Since different reward criteria give rise to markedly distinct problem settings, the resulting objective functions can vary greatly, modifying the key properties necessary to analyze the underlying system. The more extensively studied criterion is the discounted reward, whose objective function is the discounted value function: $V^{\pi}_{\mathsf{P}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right]$, where $\gamma\in[0,1)$ is some discount factor. When long-term performance is of interest, the average reward is generally adopted as the criterion of choice. By effectively setting $\gamma=1$, it does not discount the reward over time, but considers the behavior of the underlying Markov process under the steady-state distribution. Under this formulation, the average reward induced by following policy $\pi$ starting from state $s\in\mathcal{S}$ is defined as $g_{\mathsf{P}}^{\pi}(s)\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\big{[}\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\big{]}.$
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (Littman, 1994), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\otimes_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}^{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the actions of all agents. At some state $s$, all agents act simultaneously by taking joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on the Nash Equilibrium (NE). More specifically, a NE is a product policy $\pi^{*}=(\pi^{*}_{1},...,\pi^{*}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{*}$:
$V^{\pi^{*}}_{\mathsf{P},i}(s)\geq V^{(\pi^{*}_{-i},\pi_{i})}_{\mathsf{P},i}(s),\forall i\in\mathcal{N},\forall\pi_{i}\in\Pi_{i},s\in\mathcal{S}$ (1)
where $\Pi_{i}$ denotes the set of stationary policies for agent $i$, $\pi^{*}_{-i}$ denotes the joint policy of all other agents except $i$ from $\pi^{*}$, and $(\pi^{*}_{-i},\pi_{i})$ is the joint policy of $\pi^{*}_{-i}$ and an arbitrary policy $\pi_{i}$ of agent $i$. The existence of stationary NE is derived for MGs with discounted reward Shapley (1953) and average-reward MGs with additional structural assumptions Filar and Vrieze (2012).
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\otimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ Iyengar (2005); Nilim and El Ghaoui (2004); Zhang et al. (2020); Blanchet et al. (2023); Shi et al. (2024b), where the uncertainty set is independently defined over all $(s,a)$ pairs.
After all agents take their actions $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}=\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}$, and robust average reward $g^{\pi}_{\mathcal{P},i}=\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{*}$ such that
$g^{\pi^{*}}_{\mathcal{P},i}(s)\geq g^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i}(s),\forall i\in\mathcal{N},\forall\pi_{i},\forall s\in\mathcal{S}.$ (2)
3 Hardness in Average Reward Nash Equilibrium Identification
Extending existing methods for discounted or finite-horizon robust MGs to the average-reward setting is prevented by two major challenges:
Additional hidden environment player. The first major difference between a vanilla non-robust MG and a DR-MG is the introduction of an additional player, which selects the transition kernel or the underlying game model to diminish the gains of the other players Zhang et al. (2020). This additional player fundamentally changes the structure of non-robust MGs. For example, in the extensively studied special case of two-player zero-sum MGs, where $r_{1}=-r_{2}=r$, one fundamental structure of standard MGs is the Min-Max duality:
$\max_{\pi_{1}}\min_{\pi_{2}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}=\min_{\pi_{2}}\max_{\pi_{1}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}$, (3)
which directly implies the existence of a Nash Equilibrium (NE) Shapley (1953). Such results are further extended to non-robust average-reward MG studies through the limit method, which sets the discount factor $\gamma\to 1$ Filar and Vrieze (2012); Sennott (1994); Rieder (1995).
However, when looking at the distributionally robust setting for MGs, this notion fails due to consideration of the worst-case scenario. Since agents have different rewards, their corresponding worst-case transition kernels are also different. Hence, the duality fails (note that from agent $2$’s perspective, the worst case is $\max_{\mathsf{P}}V^{\pi}_{\mathsf{P},r}$ as it aims to minimize w.r.t. $r$):
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}$. (4)
Consequently, even through the uniform convergence of the discounted robust value functions enables the interchanging of the order of $\max,\min,\lim$ Wang et al. (2023a), it does not imply the existence of NE.
Complicated structure of the average reward. The second major challenge stems from the inherent complexity of the robust average reward. When compared to the Bellman equations for the robust discounted reward Iyengar (2005); Nilim and El Ghaoui (2004); Xu and Mannor (2010) and the non-robust average reward, the robust average reward has a more complicated Bellman equation Wang et al. (2023a); Wang, Zou, and Wang (2024). Primarily, its solvability is not yet justified (an incorrect proof is given in Wang et al. (2023a)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Moreover, even if it is solvable, its solution is not unique, even up to constant vectors Wang et al. (2023b). Whereas, in the non-robust setting, the average-reward Bellman equation has a unique solution up to constant vectors Puterman (1994). Therefore, despite the existence of the NE being proved for both discounted robust MGs Kardeş, Ordóñez, and Hall (2011); Zhang et al. (2020) and average-reward non-robust MGs Guo and Yang (2008); Hernández-Lerma and Lasserre (2000); Guo and Hernández-Lerma (2003); Zheng and Guo (2024); Nowak (1999); Iwase, Tanaka, and Wakuta (1976); Küenle and Schurath (2003); Tanaka and Wakuta (1977); Jaśkiewicz and Nowak (2001), the existence of the NE for average-reward robust MGs remains to be seen.
By identifying these challenges, we aim to develop a novel method to identify the Nash Equilibrium of average-reward distributionally robust MGs.
4 Solvability of the Robust Bellman Equation
In this section, we develop a deeper study of the single-agent robust Bellman equation, which will serve as an essential role in our following studies of distributionally robust games. Despite the extensive study of the robust Bellman equation, e.g., Wang et al. (2023a, b); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025); Roch et al. (2025); Roch, Atia, and Wang (2025); Chen, Wang, and Si (2025), whether this equation is solvable is in question. Hence, in this section, we aim to close this gap by examining the solvability of the robust Bellman equation.
To facilitate our analysis, we adopt the following standard irreducible assumption from (robust) average-reward studies Filar and Vrieze (2012); Xu, Mondal, and Aggarwal (2025); Xu, Ganesh, and Aggarwal (2025). Under it, the average reward is a constant and is independent from the initial state.
Assumption 4.1.
For any deterministic policy $\pi$ and any transition kernel $\mathsf{P}\in\mathcal{P}$, the induced Markov chain $(\mathcal{S},\mathcal{A},\mathsf{P}^{\pi})$ is irreducible, i.e., all states belong to a single recurrent class [^2] We naturally extend this assumption to the multi-agent setting: $\mathsf{P}^{\pi}$ is irreducible for any joint deterministic policy and any $\mathsf{P}\in\mathcal{P}$.. Moreover, $\mathcal{P}$ is compact and convex.
Remark 4.2.
For convenience and to simplify the problem, we adopt the irreducibility assumption, which is slightly stronger than the standard unichain assumption Wang et al. (2023a, b). However, our results can be extended to the unichiain cases (see Section F in the Appendix).
We then derive the solvability of the robust Bellman equation under Assumption 4.1 as follows.
Theorem 4.3.
(1). (Robust Bellman Equation)
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 4.1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the gain/bias of $(\pi,\mathsf{P}_{w}):$ $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution to the equation:
$V(s)=\sum_{a}\pi(a|s)(r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)),$ (5)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$.
(2). (Optimal Robust Bellman Equation) The optimal robust Bellman equation also has a solution:
$V(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}^{a}_{s}}(V)\}$ (6)
Our results therefore prove the solvability of the robust Bellman equation, solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (5) can have non-unique solutions even up to constant vectors. Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward Puterman (1990), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.
These results and connections are of fundamental importance. As we will later show, solving the single-agent Bellman equations will be an essential step of showing the existence of a Nash Equilibrium.
5 Average-Reward Nash Equilibrium
In this section, we study the existence of a Nash Equilibrium in a distributionally robust Markov game under the average-reward criterion. As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs, hence we propose a direct approach by studying the robust best response policy.
Definition 5.1.
(Best Response Policy) In an average-reward DR-MG, fix agent $i$ and the joint policy $\pi_{-i}$ of all other agents. The set of best response policies of agent $i$ w.r.t. $\pi_{-i}$ is defined as
$\text{BR}_{i}(\pi_{-i})\triangleq\left\{\pi_{i}:g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}=\max_{\mu_{i}}g^{(\pi_{-i},\mu_{i})}_{\mathcal{P},i}\right\}$. (7)
That is, the best response policy is agent $i$’s optimal policy when other agents fix their policies. This notion and the following results are standard and extensively studied in game theory literature Osborne and Rubinstein (2004); Fudenberg and Tirole (1991):
Lemma 5.2.
$\pi^{*}$ is a Nash Equilibrium if $\pi^{*}_{i}\in\text{BR}_{i}(\pi^{*}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{*}$ is a fixed point of the BR mapping: $\pi^{*}\in\text{BR}(\pi^{*})$, where $\text{BR}(\pi)=\otimes_{i}\text{BR}_{i}(\pi^{*}_{i})$.
Lemma 5.3.
Kakutani (1941)(Kakutani’s Fixed Point Theorem) Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) for all $x\in X$, $f(x)$ is nonempty and convex; (2) $f$ is upper semi-continuous (i.e., for any converging sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$ $\forall n$, it holds that $y\in f(x)$). Then there exists some $x^{*}\in X$ such that $x^{*}\in f(x^{*})$.
These two results imply that it is sufficient to show that the best response policy mapping, BR, satisfies the conditions in Lemma 5.3. We further show that these conditions can be further reduced to single-agent cases.
Lemma 5.4.
If for any agent $i\in\mathcal{N}$ and any stationary joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper semi-continuous, then the BR also satisfies the conditions in Lemma 5.3.
It is hence sufficient to fix an agent $i$ and study the best response policy set $\text{BR}_{i}(\cdot)$, which is a single-agent problem.
5.1 Induced Distributionally Robust MDPs
We then develop a framework of induced distributionally robust MDPs to provide a framework for multi-agent studies.
Definition 5.5.
(Induced Distributionally Robust MDP) Given a distributionally robust Markov game $\mathcal{MG}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{P},r)$ and an agent $i\in\mathcal{N}$. The induced distributionally robust MDP of the $\mathcal{MG}$ by a joint policy $\pi_{-i}$ is defined as $\mathcal{M}(\mathcal{MG},i,\pi_{-i})\triangleq(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, where
$(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}\triangleq\Big{\{}\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)\mathsf{P}\Big{|}\mathsf{P}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}\Big{\}}$,
$r^{\pi_{-i}}(s,a_{i})=\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)r_{i}(s,(a_{i},a_{-i})).$ (8)
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, the irreducibility, convexity, and compactness are all preserved through inducing. Namely, if the distributionally robust Markov game $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 4.1, then the induced distributionally robust MDP $\mathcal{M}(\mathcal{MG},i,\pi_{-i})$ satisfies Assumption 4.1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
5.2 Existence of a Nash Equilibrium
We first show the following convexity result.
Theorem 5.6.
For any distributionally robust MDP, its optimal robust policy set $\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$ is convex.
As $\text{BR}_{i}(\pi_{-i})$ is the set of optimal robust policies of the (induced) distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, this result hence directly implies the convexity of $\text{BR}_{i}(\pi_{-i})$. We highlight that this proof is not a direct extension from its counterpart under the discounted reward as we need to tackle the non-uniqueness of solutions, whereas the discounted Bellman equation admits a unique solution.
We further show the semi-continuity of the optimal policy set of the induced distributionally robust MDP.
Theorem 5.7.
For convergent policy sequences $\{\pi^{n}_{-i}\}\to\pi_{-i}$ and $\{\mu^{n}_{i}\}\to\mu_{i}$, if $\mu^{n}_{i}$ is an optimal robust policy of the induced distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi^{n}_{-i}},r^{\pi^{n}_{-i}})$ for agent $i\in\mathcal{N}$, then $\mu_{i}$ is also optimal for $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$.
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, thus resulting in a time-varying single-agent distributionally robust MDP. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
Combining all of the above results, we can finally show the existence of a Nash Equilibrium.
Theorem 5.8.
(Existence of a Nash Equilibrium) For an average-reward distributionally robust Markov game satisfying Assumption 4.1, a stationary Nash equilibrium exists.
Our result herein implies the solvability of DR-MGs in the average-reward setting. The existence of an equilibrium is the most fundamental property of a game, thus proving its existence is a necessary prerequisite for any further analysis or algorithmic design.
6 Algorithmic Solutions
We then develop an algorithmic solution for average-reward DR-MGs. Notably, most of the existing algorithms for non-robust average-reward games are developed for two-player zero-sum MGs Filar and Vrieze (2012), which cannot be extended to the robust setting due to the existence of the additional hidden environment player Zhang et al. (2020). For example, standard policy iteration can not guarantee a monotonous performance due to the robustness aspect. We propose a value iteration based algorithm in Algorithm 1, named robust Nash-Iteration.
Algorithm 1 Robust Nash-Iteration
Initialize: $h_{i}=0,\forall i\in\mathcal{N}$
while TRUE do
  for all $i\in\mathcal{N}$ do
   $h^{0}_{i}\leftarrow h_{i}$
   for all $s,a\in\mathcal{S}\times\mathcal{A}$ do
    $Q_{i}(s,a,h_{i}^{0})\leftarrow r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i}^{0})$
    $\pi(s)\leftarrow\textbf{NE}(Q_{i}(s,a,h^{0}_{i}))$
    $h_{i}(s)\leftarrow\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i}^{0})]$
   end for
  end for
  if $\max_{i}\{\text{sp}(h_{i}-h^{0}_{i})\}=0$ then
   BREAK
  end if
end while
return  $\pi$
Our algorithm maintains an estimation, $h_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, achieving such a Nash Equilibrium can be PPAD-complete Deng et al. (2023). Moreover, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies Li (2003); Hu and Wellman (2003); Littman et al. (2001); Bowling (2001, 2000). This assumption is not merely a technical convenience; it is the conceptual linchpin that stabilizes the learning dynamics of the system. By placing constraints on which Nash Equilibrium the oracle returns, we ensure that the sequence of generated policies and value functions behaves predictably, guiding the algorithm steadily toward a solution instead of allowing it to wander chaotically or diverge. We then derive the convergence of our algorithm.
Theorem 6.1.
With a specific equilibrium selection oracle NE (see Assumption D.1 in the Appendix), the robust Nash-Iteration algorithm converge to a Nash Equilibrium of the average-reward distributionally robust game.
This theorem provides the algorithmic contribution of our work, demonstrating that the proposed Robust Nash-Iteration algorithm is a principled method with guaranteed convergence. Given the known difficulties associated with multi-agent learning algorithms in general-sum games under the long-term performance criterion, our algorithm stands as a significant contribution. Taken together, our results provide a complete theoretical foundation for this problem. From establishing the existence of a robust equilibrium in Theorem 5.8 to now providing a concrete algorithm to find it, this work makes the class of average-reward DR-MGs amenable to computational analysis. It provides a blueprint for developing practical, reliable multi-agent systems that can find stable operating points even when faced with significant environmental uncertainty.
7 Connections with Discounted Nash Equilibrium
In this section, we study the connection between distributionally robust games with discounted and average rewards. Our motivation is to utilize the extensively studied and mathematical elegance of the discounted setting to solve the more challenging average-reward setting. Recall that a policy $\pi^{*}_{\gamma}$ is a Nash Equilibrium of the $\gamma$-discounted DR-MG if
$V^{\pi^{*}}_{\mathcal{P},i}\geq V^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (9)
where $V^{\pi^{*}}_{\mathcal{P},i}$ is the $\gamma$-discounted robust value function. A policy $\pi^{*}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{*}}_{\mathcal{P},i}+\epsilon\geq g^{(\pi^{*}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (10)
We then derive the following results.
Theorem 7.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
This theorem establishes a crucial bridge between the discounted and average-reward criteria in DR-MGs. Specifically, we showed that the worst-case long-run average behavior of a multi-agent system can be approximated by its discounted behavior when the future is valued almost as much as the present (i.e., as $\gamma\to 1$).
The importance of this theorem lies in its computational implications. The analysis of $\gamma$-discounted games is often more tractable Kardeş, Ordóñez, and Hall (2011), benefiting from properties like the Banach fixed-point theorem which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as value iteration and policy iteration, are naturally formulated for the discounted setting. Our results hence provide a rigorous justification for using these well-understood tools to solve the more challenging average reward problem.
Remark 7.2.
Part (2) of Theorem 7.1 offers a more practical and constructive guarantee. In practice (like data-driven learning), it may be sufficient to find some approximation of the NE instead of the exact NE. This result therefore assures us that for any desired level of accuracy $\epsilon$, we can simply select a single discount factor $\gamma$ sufficiently close to 1. Any $\epsilon$-Nash Equilibrium found for this $\gamma$-discounted game will be an $\epsilon$-Nash Equilibrium for the average-reward game. This provides a clear and direct pathway for finding near-equilibrium policies for the average-reward criterion. Moreover, we can derive the following result to specify the reduction to discounted DR-MGs.
Lemma 7.3.
Define the robust diameter Auer, Jaksch, and Ortner (2008) of the distributionally robust Markov game by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
The result hence specifies a reduction factor whose Nash Equilibrium is an approximated equilibrium under the average reward criterion. Thus, we can apply existing algorithms for discounted DR-MGs Zhang et al. (2020) to approximately solve the average reward ones.
8 Numerical Experiments
In this section, we develop numerical experiments to validate the advantages of our formulation and methods. We mainly consider two aspects: the long-term performance and the robustness to environment uncertainty.
8.1 Long-Term Performance
To empirically validate our approach, we perform a series of experiments designed to compare the performance of our robust Nash-Iteration algorithm with average reward against the standard discounted-reward robust Nash-Iteration Zhang et al. (2020); Shi et al. (2024b) under DR-MGs. We aim to show that our algorithm will yield a joint policy that has better long-term performance than those derived from the discounted criterion.
We design a "Structured Random Environment" to create a complex strategic challenge with meaningful trade-offs between short-term gains and long-term consequences. The environment is a DR-MG with 20 states and 5 actions. The state space $S$ is partitioned into two clusters:
  • Prosperous States ($|S_{p}|=5$): A small set of states characterized by a higher mean reward (drawn from $\mathsf{N}(2,1)$).
  • Deprived States ($|S_{d}|=15$): A larger set of states with a lower mean reward (drawn from $\mathsf{N}(-2,1)$),
where $\mathsf{N}(\cdot,\cdot)$ is the Gaussian distribution. The nominal transition kernel $\mathsf{P}_{0}$ is also structured to create persistence within these clusters. Transitions originating from a state in a given cluster are five times more likely to terminate in a state within the same cluster. This design creates a strategic trap: an agent in a prosperous state might be tempted by an action with a high immediate reward that carries a significant risk of transitioning it to the deprived cluster, from which escape is difficult.
We further set the uncertainty set using a Kullback-Leibler (KL) divergence ball around the nominal kernel $\mathsf{P}_{0}$:
$\mathcal{P}^{a}_{s}=\{q\in\Delta(\mathcal{S})\ |\ D_{KL}(q||(\mathsf{P}_{0})^{a}_{s})\leq\theta\}.$
Based on the scale of our environment, the uncertainty budget $\theta$ is set to a modest value of 0.01.
We compare our average-reward Robust Nash-Iteration algorithm (Algorithm 1) and the Discounted-Reward Robust Nash-Iteration. We implement the two algorithms and evaluate their output policies under the robust average-reward criterion. The experimental procedure is as follows:
(1). The optimal average-reward policy, $\pi_{avg}$, is computed using the average-reward Robust Nash-Iteration algorithm. We evaluate the worst-case average reward of Player 1 under the policy, using the robust average reward single-agent RL method in Wang et al. (2023a).
(2). For a range of discount factors $\gamma$ from 0.5 to 0.99, the corresponding optimal discounted policy, $\pi_{disc}(\gamma)$, is computed using the Discounted-Reward Robust Nash-Iteration algorithm. And for each $\pi_{disc}(\gamma)$, we similarly evaluate its worst-case long-term performance.
For both algorithms, we plot the performance of Player 1 as a performance metric. Results are presented in Figure 1 (under different random environments). Policies learned under the discounted criterion consistently underperform the dedicated average-reward policy. This gap is most pronounced for smaller values of $\gamma$, where the myopic nature of the discounted algorithm leads it to fall into the strategic trap of the environment by choosing actions that yield high immediate rewards but lead to the persistent, low-reward "deprived" cluster of states. As $\gamma$ approaches 1, the discounted policy becomes more far-sighted, and its performance converges toward the optimal baseline.
Our results hence provide strong evidence that, for DRMGs under the average reward criterion with long-term performance optimization goals, our robust Nash-Iteration algorithm achieves a better performance, compared to the common practice of approximating this objective with a discounted formulation using $\gamma\approx 1$.
Figure 1: Performance Comparisons of Average and Discounted Robust Nash-Iteration
Figure 2: Performance Comparisons of Robust/Non-Robust Nash-Iteration
8.2 Robustness
We then test the robustness of our method by comparing it with the non-robust average-reward Nash-Iteration algorithm Li (2003), to test their ability to handle model uncertainty. Using the same environment, we evaluate both learners against a worst-case adversary throughout training to measure their inherent robustness. The results are presented in Figure 2. The policy learned by the robust algorithm consistently and monotonically improves its worst-case performance, rapidly converging to a high, stable value. This is because the algorithm directly optimizes for resilience against the adversary at each step. In contrast, the policy learned by the non-robust algorithm exhibits lower performance when evaluated against a worst-case adversary. Ultimately, it does not have the ability to reliably account for perturbations in the underlying environment, as opposed to its robust counterpart, which aligns with our theoretical results.
These numerical results therefore verify that our method is able to achieve superior long-term performance with increased stability when in the face of model uncertainty in multi-agent systems, validating our theoretical results.
9 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We established the well-posedness of the robust average-reward Bellman equation, proved the existence of a stationary Nash Equilibrium, and proposed the Robust Nash-Iteration algorithm as the first provably convergent method for this game class. We also connected average-reward equilibria to their discounted counterparts, and our numerical experiments validated the superior long-term performance and stability of our approach. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
Appendix A Preliminaries
In this section we briefly review previous studies of single-agent distributionally robust MDPs.
A.1 Discounted Setting
Under the robust setting, the worst-case performance over the uncertainty set of MDPs is defined through the discounted reward. More specifically, the robust discounted value function of a policy $\pi$ for a discounted MDP is defined as
$V^{\pi}_{\mathcal{P},\gamma}(s)\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\mathbb{E}_{\pi,\kappa}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right],$ (11)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{t\geq 0}\mathcal{P}$.
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator Nilim and El Ghaoui (2004); Iyengar (2005); Puterman (1994):
$\mathbf{T}_{\pi}V(s)\triangleq\sum_{a\in\mathcal{A}}\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right),$ (12)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{p\in\mathcal{P}^{a}_{s}}p^{\top}V$ is the support function of $V$ on $\mathcal{P}^{a}_{s}$. Based on the contraction of $\mathbf{T}_{\pi}$, robust dynamic programming approaches, e.g., robust value iteration, can be designed Nilim and El Ghaoui (2004); Iyengar (2005).
The goal is to find the optimal robust policy:
$\pi^{*}=\arg\max_{\pi}V^{\pi}_{\mathcal{P},\gamma}(s),$ (13)
and it is shown Iyengar (2005) that the optimal robust value function $V^{\pi^{*}}_{\mathcal{P},\gamma}(s)$ satisfies the optimal robust Bellman equation
$V(s)\triangleq\max_{a}\left\{\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right)\right\}.$ (14)
A.2 Average Reward Setting
For a fixed kernel $\mathsf{P}_{0}$, the average reward (or gain) and the bias are defined as
$g_{\mathsf{P}_{0}}^{\pi}(s)$ $\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right]$ (15)
and
$h_{\mathsf{P}_{0}}^{\pi}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\sum^{\infty}_{t=0}(r_{t}-g_{\mathsf{P}_{0}}^{\pi})|s_{0}=s\right].$ (16)
We can view equation (16) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by Puterman (1994), thus it holds the following non-robust Bellman equation:
$h^{\pi}_{\mathsf{P}_{0}}(s)=\mathbb{E}_{\pi}\left[r(s,a)-g^{\pi}_{\mathsf{P}}(s)+\sum_{s^{\prime}\in\mathcal{S}}p^{a}_{s,s^{\prime}}h_{\mathsf{P}_{0}}^{\pi}(s^{\prime})\right].$ (17)
In the robust setting, we consider the robust average reward, defined as
$g_{\mathcal{P}}^{\pi}(s)$ $\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\liminf_{n\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right].$ (18)
It is shown that under the unichain assumption, the robust average reward is independent from the initial state: $g_{\mathcal{P}}^{\pi}(s_{1})=g_{\mathcal{P}}^{\pi}(s_{2})$, for any $s_{1},s_{2}$. The robust Bellman equation (for a policy $\pi$) and the robust optimal Bellman equation for the average reward setting are derived in Wang et al. (2023a):
$V(s)+g$ $=\sum_{a}\pi(a|s)\big{(}r(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{)},$ (19)
$V(s)$ $=\max_{a}\big{\{}r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big{\}},\quad\forall s\in\mathcal{S}.$ (20)
It is shown Wang et al. (2023b) that if the equation (19) ((20)) has a solution $(g,V)$, then the solution $g$ is the robust average reward $g^{\pi}_{\mathcal{P}}$ (the optimal robust average reward $g^{*}_{\mathcal{P}}=\max_{\pi}g^{\pi}_{\mathcal{P}}$). However, whether these two equations are solvable is not studied but directly assumed.
Appendix B Proof of Section 4
Theorem B.1.
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 4.1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the robust Bellman equation:
$V=r^{\pi}-g+\sigma^{\pi}(V)$ (21)
has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$.
Proof.
Note that for $\mathsf{P}_{w}$, it holds the corresponding non-robust Bellman equation:
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (22)
We further denote that
$\mathsf{P}_{V}=\arg\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}},$ (23)
then it also holds that
$V^{\pi}_{\mathsf{P}_{V}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{V}}.$ (24)
Combining (22) and (24) implies that
$(\mathsf{P}^{\pi}_{w}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}}.$ (25)
By the definition of $\mathsf{P}_{V}$, it holds that $\mathsf{P}_{V}^{\pi}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}_{w}^{\pi}V^{\pi}_{\mathsf{P}_{w}}$, thus
$(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}},$ (26)
and it holds that
$\mathsf{P}^{\pi}_{V}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\leq(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}).$ (27)
Recursively applying it further implies that
$(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\geq(\mathsf{P}^{\pi}_{V})^{k}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}),\forall k.$ (28)
By denoting $V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}$ by $x$, the inequality above can be rewritten as
$x\geq(\mathsf{P}^{\pi}_{V})^{k}x,\forall k.$ (29)
We denote $s^{*}\triangleq\arg\min_{s\in\mathcal{S}}x(s)$, then it holds that
$x(s^{*})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s).$ (30)
By our assumption, for any $s^{\prime}\in\mathcal{S}$, there exists some $k$ such that $((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}>0$. Then it holds that
$x(s^{*})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s^{*})=x(s^{*}),$ (31)
which implies that
$x(s^{*})=((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime})+(1-((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}})x(s^{*}).$ (32)
We then have that
$((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{*})\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{*},s^{\prime}}x(s^{\prime}),$ (33)
which implies $x(s^{*})=x(s^{\prime}),\forall s^{\prime}\in\mathcal{S}$, and thus $x=me$ for some constant $m$, i.e., $V^{\pi}_{\mathsf{P}_{w}}=me+V^{\pi}_{\mathsf{P}_{V}}$.
Now we plug it in (22) and (24), and it holds that
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$ (34)
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}},$ (35)
which implies that
$g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (36)
Note that $g^{\pi}_{\mathsf{P}_{w}}\leq g^{\pi}_{\mathsf{P}_{V}}$, due to $\mathsf{P}_{w}$ is the worst-case kernel w.r.t. $g^{\pi}_{\mathsf{P}}$, and $\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}$, due to the definition of $\mathsf{P}_{V}$, thus
$0\leq g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}\leq 0.$ (37)
It further implies that
$\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$
i.e.,
$\sigma^{\pi}({V^{\pi}_{\mathsf{P}_{w}}})=\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (38)
Combine it with (22) and we have that
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\sigma^{\pi}(V^{\pi}_{\mathsf{P}_{w}}).$ (39)
Hence the Bellman equation (5) has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$, which completes the proof. ∎
Corollary B.2.
The optimal robust Bellman equation
$V(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}^{a}_{s}}(V)\}$ (40)
is solvable.
Proof.
Denote the optimal robust policy by $\pi^{*}$, and the corresponding worst-case kernel by $\mathsf{P}^{*}$. By Theorem 4.3, $(g^{\pi^{*}}_{\mathcal{P}},h^{\pi^{*}}_{\mathsf{P}^{*}})$ is the solution to the robust Bellman equation of $\pi^{*}$:
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)=r^{\pi^{*}}(s)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{\pi^{*}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})=\sum_{a}\pi^{*}(a|s)\left(r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right).$ (41)
We now set $\pi^{\prime}(s)=\arg\max_{a}\{r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\},\forall s$, then we have that
$r^{\pi^{\prime}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})=g^{\pi^{\prime}}_{\mathcal{P}}\leq g^{\pi^{*}}_{\mathcal{P}}$ $=\sum_{a}\pi^{*}(a|s)\left(r(s,a)-h^{\pi^{*}}_{\mathsf{P}^{*}}(s)+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right)$
$\leq r^{\pi^{\prime}}(s)-h^{\pi^{*}}_{\mathsf{P}^{*}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}}),$ (42)
which further implies that
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)\leq\sigma^{\pi^{\prime}}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})-\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})\leq\mathsf{P}^{\prime}(h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}).$ (43)
Now denote $x=h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}$, it holds that
$x\leq\mu x,$ (44)
for some positive distribution matrix $\mu$. Similarly, it implies $x(s)=x(s^{\prime})$, and thus $h^{\pi^{*}}_{\mathsf{P}^{*}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}=ce$. This further implies that all inequalities in (42) are equations, thus
$h^{\pi^{*}}_{\mathsf{P}^{*}}(s)=\max_{a}\left\{r(s,a)-g^{\pi^{*}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{*}}_{\mathsf{P}^{*}})\right\},$ (45)
hence $(h^{\pi^{*}}_{\mathsf{P}^{*}},g^{\pi^{*}}_{\mathsf{P}^{*}})$ is a solution to (6), completing the proof of the first part. The remaining proof follows in a similar way from Wang et al. (2023b). ∎
Appendix C Proof of Section 5
Lemma C.1.
If for any agent $i$ and any joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper semi-continuous, then BR also satisfies the conditions in Lemma 5.3
Proof.
(1). If for any $i$ and $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, then $\text{BR}(\pi)=(\text{BR}_{1}(\pi_{-1}),...,\text{BR}_{N}(\pi_{-N}))$ is clearly nonempty.
(2). Consider two policies $\pi^{1},\pi^{2}\in\text{BR}(\pi)$. To show the convexity of $\text{BR}(\pi)$, we need to show $\pi^{\prime}=\alpha\pi^{1}+(1-\alpha)\pi^{2}=(\alpha\pi^{1}_{1}+(1-\alpha)\pi^{2}_{1},...,\alpha\pi^{1}_{N}+(1-\alpha)\pi^{2}_{N})\in\text{BR}(\pi)$. Note that for any $i$, $\pi^{1}_{i},\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, and due to the convexity of $\text{BR}_{i}(\pi_{-i})$, $\alpha\pi^{1}_{i}+(1-\alpha)\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, which completes the proof.
(3). Consider two convergent policy sequences $\{\pi^{n}\}\to\pi$ and $\{\mu^{n}\}\to\mu$, with $\mu^{n}\in\text{BR}(\pi^{n})$, we need to show that $\mu\in\text{BR}(\pi)$. Note that $\pi^{n}\to\pi$ implies that $\pi^{n}_{i}\to\pi_{i}$ for any $i$, and similarly $\mu^{n}\to\mu$. Since $\mu_{i}^{n}\in\text{BR}_{i}(\pi_{-i}^{n})$, by the semi-continuity of $\text{BR}_{i}$, we have $\mu_{i}\in\text{BR}_{i}(\pi_{-i})$. Moreover note that
$\mu\in\text{BR}(\pi)\Leftrightarrow\mu_{i}\in\text{BR}_{i}(\pi_{-i}),\forall i,$ (46)
which hence completes the proof. ∎
C.1 Convexity
Lemma C.2.
Consider the optimal robust Bellman equation
$g+h(s)=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (47)
and denote any of its solutions by $(g,h)$. Additionally, consider an optimal robust policy $\pi$. Then $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\pi$:
$g+h(s)=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ (48)
Proof.
From Theorem 4.3, (48) is solvable, and we denote one of its solutions by $(g^{\pi},h^{\pi})$. Then we have that
$g+h(s)$ $=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (49)
$g^{\pi}+h^{\pi}(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}.$ (50)
Now from Theorem 4.3, $g=g^{\pi}=g^{*}_{\mathcal{P}}$, thus it holds that
$h^{\pi}(s)-h(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\sum_{a}\mu(a|s)\{\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})-\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq q_{s}^{\mu}(h^{\pi}-h),$ (51)
for some policy $\mu$ and some kernel $q\in\mathcal{P}$. By the irreducibility of $q^{\mu}$, $h^{\pi}-h=ce$. Thus $(g,h)=(g^{\pi},h^{\pi}+ce)$ is also a solution to (48). ∎
Based on the equivalence between the best response policy set and the optimal policy set of the induced robust MDP, the convexity can be derived from the following result.
Lemma C.3.
Define $\Pi^{*}\triangleq\{\pi\ |\ \text{optimal policies of }g^{\pi}_{\mathcal{P}}\}$. Then $\Pi^{*}$ is convex.
Proof.
Let $\pi_{1},\pi_{2}\in\Pi^{*}$ and define $\tilde{\pi}=\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1].$ Then by Theorem 4.3,
[…]
Alternatively, let $(g,h)$ be a solution to the optimal equation, i.e. $h(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ From Lemma C.2, this $(g,h)$ also solves the equation w.r.t. $\pi_{1},\pi_{2}$, i.e.
$g+h=r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\quad\text{and}\quad g+h=r^{\pi_{2}}+\sigma^{\pi_{2}}(h).$ (53)
It then follows that,
$g+h$ $=\alpha\big{(}r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\big{)}+(1-\alpha)\big{(}r^{\pi_{2}}+\sigma^{\pi_{2}}(h)\big{)}$ (54)
$=\sum_{a}\big{(}\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big{)}r(s,a)+\sum_{a}\big{(}\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big{)}\big{(}\sigma_{\mathcal{P}^{a}_{s}}(h)\big{)}$ (55)
$=r^{\tilde{\pi}}+\sigma^{\tilde{\pi}}(h).$ (56)
Consequently, $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\tilde{\pi}$. This implies that $g^{\tilde{\pi}}=g=g^{*}$, and thus that $\tilde{\pi}$ is also optimal.
Hence the optimal policy set is convex, which completes the proof. ∎
C.2 Semi-continuity
Lemma C.4.
Consider a policy sequence $\{\pi_{n}\}\rightarrow\pi$, and the sequence of induced transition matrices $\{\mathsf{P}_{\pi_{n}}\}$. It holds that $\mathsf{P}_{\pi_{n}}^{*}\rightarrow\mathsf{P}_{\pi}^{*}$ uniformly on $\mathcal{P}$.
Proof.
See Prop 8.4.6 in Puterman (1994) for the proof. ∎
Lemma C.5.
Let $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ be a joint policy of all other agents except agent $i$. Define the operator $\mathcal{T}^{\pi}_{\mathcal{P}}:\mathbb{R}^{\mathcal{S}}\rightarrow\mathbb{R}^{\mathcal{S}\mathcal{A}_{i}}$ as $\mathcal{T}^{\pi}_{\mathcal{P}^{a}_{s}}(h)(s,a)=r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)$, where $(\mathcal{P}^{\pi})^{a}_{s}=\big{\{}q(\cdot)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)\mathsf{P}(\cdot|s,(a,b))\ :\ \mathsf{P}(\cdot|s,(a,b))\in\mathcal{P}^{(a,b)}_{s}\big{\}}$ and $r^{\pi}(s,a)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)r(s,a,b).$ Consider the sequences $\{\pi_{n}\}\rightarrow\pi$ and $\{h_{n}\}\rightarrow h$, then $\mathcal{T}^{\pi_{n}}_{\mathcal{P}}(h_{n})\rightarrow\mathcal{T}^{\pi}_{\mathcal{P}}(h)$ uniformly on $\mathcal{S}\times\mathcal{A}_{i}$.
Proof.
Recalling the definition of $\mathcal{T}_{\mathcal{P}}$ we can obtain the difference,
[…]
$=\max_{s,a}\big{|}r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-r^{\pi}(s,a)\big{|}$
$\leq\max_{s,a}\big{\{}|r^{\pi_{n}}(s,a)-r^{\pi}(s,a)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|\big{\}}$
[…]
$B$ $=|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|$
$=|\min_{q\in(\mathcal{P}^{\pi_{n}})^{a}_{s}}qh-\min_{q\in(\mathcal{P}^{\pi})^{a}_{s}}qh|$
$=|\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi_{n}}h-\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi}h|$
$=\bigg{|}\min_{q\in\mathcal{P}^{a}_{s}}\Big{(}\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi_{n}(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big{)}-\min_{q\in\mathcal{P}^{a}_{s}}\Big{(}\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big{)}\bigg{|}$
$\leq\bigg{|}\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\Big{(}\pi_{n}(b|s)-\pi(b|s)\Big{)}q(s^{\prime}|s,a,b)h(s^{\prime})\bigg{|},\quad\quad\text{for some }q\in\mathcal{P}^{a}_{s},$
$\leq||\pi_{n}(\cdot|s)-\pi(\cdot|s)||_{1}\cdot H,$
[…]
Note that $\pi_{n}(s)\rightarrow\pi(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},$ $\pi_{n}(s)-\pi(s)||_{1}\leq\epsilon$. Similarly with $h_{n}(s)\rightarrow h(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},\ |h_{n}(s)-h(s)|\leq\epsilon.$ Now by setting $N=\max_{s\in\mathcal{S}}\{N_{s}\}$, then $\forall n>N,$
[…]
Therefore, when $n>N,$ $\pi_{n}(s)-\pi(s)||_{1}+A+B\leq\epsilon+\epsilon+H\epsilon\lesssim\epsilon$. This implies that $\mathcal{T}^{n}_{\mathcal{P}}\rightarrow\mathcal{T}_{\mathcal{P}}$ uniformly on $\mathcal{S}\times\mathcal{A}$, which completes the proof. ∎
Theorem C.6.
(Semi-continuity) Let $\pi_{n}\rightarrow\pi$ and $x_{n}\rightarrow x$, where $\pi_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ and $x_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$. If $x_{n}\in\text{BR}_{i}(\pi_{n}),$ then $x\in\text{BR}_{i}(\pi).$
Proof.
To show that $x\in\text{BR}_{i}(\pi),$ it suffices to show that $x$ is an optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi},r^{\pi}).$ Recalling the robust Bellman equation, it is also sufficient to show that $g^{(x,\pi)}_{\mathcal{P}}$ satisfies the following:
$g(s)=\max_{a\in\mathcal{A}_{i}}\{r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-h(s)\}=\max_{a\in\mathcal{A}_{i}}\{\mathcal{T}^{\pi}_{\mathcal{P}_{s}^{a}}(h)\}.$ (58)
See that since $x_{n}\in\text{BR}_{i}(\pi_{n})$, then $x_{n}$ is the optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{n}},r^{\pi_{n}}).$ Thus, there exists $h_{n}$, s.t. $(h_{n},g_{n}=g^{x_{n}}_{\mathcal{P}^{\pi_{n}}})$ is a solution to,
$g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g_{n}(s)=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}.$
See that $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=\min_{\mathsf{P}^{\pi_{n}}\in\mathcal{P}^{\pi_{n}}}g^{x_{n}}_{{}_{\mathsf{P}^{\pi_{n}}},r^{\pi_{n}}}=\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n},\pi_{n})}_{\mathsf{P},n}$, and so $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g^{\pi_{n},x_{n}}_{\mathcal{P}}$. Therefore, for any $(x_{n},\pi_{n})$, there exists some $h_{n}\in\mathbb{R}^{\mathcal{S}}$ such that for all $s\in\mathcal{S}$,
$g^{(x_{n},\pi_{n})}_{\mathcal{P}_{s}}(s)$ $=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}$
$=\max_{a}\{\mathcal{T}^{\pi_{n}}_{\mathcal{P}_{s}^{a}}(h_{n})\}.$ (59)
Moreover, since $h_{n}$ is some relative function Wang et al. (2023a), $h_{n}||_{\infty}\leq H.$ Then, there exists some subsequence $\{n_{m}\},$ s.t. $h_{n_{m}}\rightarrow h$ for some $h$, and $g^{(x_{n},\pi_{n})}_{\mathcal{P}}\rightarrow g$ for some $g$. Then, we can take the limit as $m\rightarrow\infty$ in (C.2) to find:
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\max_{a}\big{\{}\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big{\}}$
$\overset{(i)}{=}\max_{a}\big{\{}\lim_{m\rightarrow\infty}\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big{\}}$
$=\max_{a}\big{\{}\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big{\}},$ (60)
where $(i)$ holds by the uniform convergence in Lemma C.5, and the final equality by $\pi_{n_{m}}\rightarrow\pi$, and $h_{n_{m}}\rightarrow h$. Looking at the left-hand side of equation (C.2) we find
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}^{*}_{(x_{n_{m}},\pi_{n_{m}})}\cdot r^{(x_{n_{m}},\pi_{n_{m}})}.$
By Lemma C.4 we have $\mathsf{P}^{*}_{(x_{n_{m}},\pi_{n_{m}})}\rightarrow\mathsf{P}^{*}_{(x,\pi)}$ uniformly. Thus, $\lim\min=\min\lim$ implies,
$\lim_{m\rightarrow\infty}g_{\mathcal{P}}^{(x_{n_{m}},\pi_{n_{m}})}$ $=\min_{\mathsf{P}\in\mathcal{P}}\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\min_{\mathsf{P}\in\mathcal{P}}g^{(x,\pi)}_{\mathsf{P}}$
$=g^{(x,\pi)}_{\mathcal{P}}$.
Therefore, $g^{(x,\pi)}_{\mathcal{P}}=\max_{a}\big{\{}\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big{\}}$ for some $h\in\mathbb{R}^{\mathcal{S}}$, which completes the proof. ∎
Appendix D Proof of Section 6
Assumption D.1.
In Algorithm 1, at each step, the Nash Equilibrium $\pi(s)=\textbf{NE}(Q_{i}(s,a,h^{i}))$ satisfies one of the following conditions:
  • $\pi$ is global optimal: $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\geq\mathbb{E}_{\pi^{\prime}}[Q_{i}(s,a,h^{i})]$ for all joint policy $\pi^{\prime}$,$i,s,a$.
  • $\pi$ is a saddle point such that $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\leq\mathbb{E}_{(\pi^{\prime}_{-i},\pi_{i})}[Q_{i}(s,a,h^{i})]$, for all $i,s,a,$ and joint policy $\pi^{\prime}_{-i}$.
Such an assumption is necessary for convergence Li (2003); Hu and Wellman (2003); Littman et al. (2001); Bowling (2001, 2000).
Theorem D.2.
Additionally assume Assumption D.1. The robust Nash-Iteration algorithm converge to an average reward Nash Equilibrium of the distributionally robust game.
Proof.
We first show that, if the algorithm converges to $\pi$, then $\pi$ is an average reward Nash Equilibrium. To show this, we note that the convergence of Alg 1 implies that there exists some constant $c_{1},...,c_{N}$, such that
$\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i})]-h_{i}(s)=c_{i},\forall i,s.$ (61)
This implies that
$h_{i}(s)+c_{i}=\sum_{a}\pi(a|s)(r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})),$ (62)
which is the robust Bellman equation w.r.t. $\pi$. Hence by Theorem 4.3, $(c_{i},h_{i})=(g^{\pi}_{\mathcal{P},i},h^{\pi}_{\mathcal{P},i})$ for some vector $h^{\pi}_{\mathcal{P},i}$, i.e.,
$h^{\pi}_{\mathcal{P},i}+g^{\pi}_{\mathcal{P},i}=r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i}).$ (63)
Now assume that $\pi$ is not an average reward Nash Equilibrium, then there exists some agent $i$ and some policy of it $\mu_{i}$, such that
$g^{(\pi_{-i},\mu)}_{\mathcal{P},i}>g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}$. (64)
Similarly, there exists some $h^{\mu}_{i}$ satisfying
$h^{\mu}_{i}+g^{(\pi_{-i},\mu)}_{\mathcal{P},i}=r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i}).$ (65)
Combining (63),(64),(65) implies that
$r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}=g^{(\pi_{-i},\mu)}_{\mathcal{P},i}>g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}=r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-h^{\pi}_{\mathcal{P},i}.$ (66)
Moreover, since $\pi$ is the Nash Equilibrium of $Q_{i}$, i.e.,
$r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i})\leq r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i}),$ (67)
hence
$r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-h^{\pi}_{\mathcal{P},i}<r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}\leq\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}+r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i}),$ (68)
i.e.,
$h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i}<\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i})\leq P(h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i})$ (69)
for some $P\in\mathcal{P}$. Due to the irreducibility, it holds that
$h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i}=c_{i}e$ (70)
for some constant $c_{i}$. Thus $g^{(\pi_{-i},\mu)}_{\mathcal{P},i}=g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}$, which is a contradiction to (64), and hence $\pi$ is an average reward Nash Equilibrium.
It then remains to show the convergence of Algorithm 1, and it is sufficient to show that for any $i$, the operator $\mathbf{T}=(\mathbf{T}_{1},...,\mathbf{T}_{N}):\mathbb{R}^{NS}\to\mathbb{R}^{NS}$ is a contraction under Span semi-norm, where
$\mathbf{T}_{i}:\mathbb{R}^{NS}\to\mathbb{R}^{S}:\mathbf{T}_{i}(h=(h_{1},...,h_{N}))(s)=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})],$ (71)
here $\pi$ is the Nash Equilibrium specified by Algorithm 1. Consider $h,h^{\prime}$, and we consider the two cases from Assumption D.1. If $\pi$ is global optimal, then
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)$ $=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$=\max_{\pi(s)}\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\max_{\pi^{\prime}(s)}\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\mathbb{E}_{\pi(s)}[\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\max_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i}));$ (72)
And if the second case from Assumption D.1 holds, then
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)$ $=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\overset{(a)}{\leq}\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\overset{(b)}{\leq}\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\max_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})),$ (73)
where $(a)$ is due to $\pi^{\prime}$ is the Nash Equilibrium, and $(b)$ is due to the saddle point assumption. Similarly, we can show that
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)\geq\min_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})).$ (74)
Thus Lemma G.2 implies that
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq\text{sp}(\sigma_{\mathcal{P}}(h_{i})-\sigma_{\mathcal{P}}(h^{\prime}_{i}))=\text{sp}(\mathsf{P}(h_{i}-h^{\prime}_{i})),$ (75)
for some irreducible transition matrix $\mathsf{P}$. Applying Lemma G.1 implies
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq(1-\alpha(\mathsf{P}))\text{sp}(h\mathsf{P}_{0}i-h_{i}^{\prime}).$ (76)
Now since for irreducible transition matrix $\mathsf{P}$, $\alpha(\mathsf{P})<1$ Puterman (1994), then
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq\gamma_{i}\text{sp}(h_{i}-h_{i}^{\prime})$ (77)
for some constant $\gamma_{i}<1$. Taking the maximum over $i\in\mathcal{N}$ hence implies the Span-contraction, which completes the proof. ∎
Appendix E Proof of Section 7
Theorem E.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
Proof.
(1). By the uniform convergence of the robust discounted value function derived in Wang et al. (2023a), we have that
$\lim_{\gamma\to 1}(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}=g^{\pi}_{\mathcal{P},i},\text{ uniformly on }\Pi.$ (78)
Thus for any $\epsilon$, there exists some $\gamma_{e}$, such that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,$ (79)
for any $\pi\in\Pi$ and $\gamma>\gamma_{e}$. Moreover, there exists some constant $T$ such that for any $t>T$,
$\gamma_{t}>\gamma_{e},\|\pi_{t}-\pi\|\leq\epsilon.$ (80)
Since $\pi_{t}$ is the Nash Equilibrium of the discounted robust MG, it holds that
$V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}\geq V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i},\forall i\in\mathcal{N},\forall\mu^{i}.$ (81)
Thus we have that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$
$=-g^{\pi}_{\mathcal{P},i}+(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}$
$\quad-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}$. (82)
Consider any $t>T$, it first holds that due to (79):
$(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$ $\leq\epsilon,$ (83)
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}$ $\leq\epsilon.$ (84)
Moreover,
$(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}\leq(1-\gamma_{t})(V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-V^{\pi}_{\gamma_{t},\mathcal{P},i})\leq\epsilon,$ (85)
$(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq\epsilon,$ (86)
which is due to the $\frac{1}{1-\gamma}$-Lipschitz of the robust discounted value function in $\pi$, and the closeness of $\pi_{t}$ and $\pi$ in (80). Finally,
$-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq 0$ (87)
due to (81). Combining all results hence implies that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq\epsilon,$ (88)
for any $i$ and $\mu^{i}$ uniformly. Thus let $\epsilon\to 0$ we complete the proof.
(2). Set $\gamma$ to be any factor that is larger than $\gamma_{e}$. Then
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}=g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}.$ (89)
Note that for $\gamma>\gamma_{e}$, it holds that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq 2\epsilon+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}\leq 3\epsilon,$ (90)
due to $\pi$ is an $\epsilon$-NE for the discounted robust MG, which hence completes the proof. ∎
Lemma E.2.
The discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
Proof.
The proof is a direct corollary of Theorem 7.1 and results from Roch, Atia, and Wang (2025). Specifically, it is shown in Theorem 3.4 in Roch, Atia, and Wang (2025) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^3] The results in Roch, Atia, and Wang (2025) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ Roch, Atia, and Wang (2025)., it holds that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,\forall\pi,\forall i.$ (91)
Combining with Theorem 7.1 further completes the proof. ∎
Appendix F Extension to Unichain Cases
In this section, we mainly discuss the extension of the solvability results to the unichain cases. The remaining results can be derived following a similar way, by re-deriving the results based on the multi-step operator introduced below.
Definition F.1 (Robust Bellman Operators).
For a fixed policy $\pi$, the robust transition operator $\mathcal{T}^{\pi}_{\mathcal{P}}:\mathbb{R}^{|\mathcal{S}|}\to\mathbb{R}^{|\mathcal{S}|}$ is defined as:
$(\mathcal{T}_{\mathcal{P}_{s}^{\pi(s)}}V)(s):=\min_{\mathsf{P}\in\mathbb{P}_{s}^{\pi(s)}}\sum_{s^{\prime}\in\mathcal{S}}\mathsf{P}(s^{\prime}|s,\pi(s))V(s^{\prime})$
The robust Bellman evaluation operator $\hat{\mathcal{T}}^{\pi}_{\mathcal{P}}:\mathbb{R}^{|\mathcal{S}|}\to\mathbb{R}^{|\mathcal{S}|}$ is defined as:
$(\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s)}_{s}}V)(s):=r(s,\pi(s))+(\mathcal{T}_{\mathcal{P}^{\pi(s)}_{s}}V)(s)$
where $r(s,\pi(s))$ is the reward and $\mathbb{P}^{a}_{s}$ is the uncertainty set of transition kernels for state-action pair $(s,a)$.
We first derive the following lemma by extending from the non-robust results from Puterman (1990).
Lemma F.2 (Multi-Step Span-Contraction).
Assume the robust unichain assumption holds for the policy $\pi$. Then the operator $\mathcal{T}_{\mathcal{P}}^{\pi}$ is a multi-step span-contraction under some additional technique assumptions. That is, there exists an integer $k\geq 1$ and a modulus $\gamma\in[0,1)$ such that for any vectors $V,U\in\mathbb{R}^{|\mathcal{S}|}$:
$\text{Span}((\mathcal{T}_{\mathcal{P}}^{\pi})^{k}(V)-(\mathcal{T}_{\mathcal{P}}^{\pi})^{k}(U))\leq\gamma\cdot\text{Span}(V-U)$
Proof.
The proof is derived similarly to the Theorem 8.5.2 of Puterman (1994). ∎
Corollary F.3.
The operator $\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}$ inherits the same multi-step span-contraction property, since for any $V,U$:
$\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V)-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(U)=(r^{\pi}+\mathcal{T}_{\mathcal{P}}^{\pi}(V))-(r^{\pi}+\mathcal{T}_{\mathcal{P}}^{\pi}(U))=\mathcal{T}_{\mathcal{P}}^{\pi}(V)-\mathcal{T}_{\mathcal{P}}^{\pi}(U)$
Therefore, $\text{Span}((\hat{\mathcal{T}}_{\mathcal{P}}^{\pi})^{k}(V)-(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi})^{k}(U))\leq\gamma\cdot\text{Span}(V-U)$.
We now state and prove the main theorem.
Theorem F.4.
Under the robust unichain assumption for a fixed policy $\pi$, the robust average-reward Bellman equation (5) has a solution.
Proof.
We establish existence by analyzing the convergence of the Relative Value Iteration (RVI) algorithm. Let $V_{0}\in\mathbb{R}^{|\mathcal{S}|}$ be an arbitrary initial vector, and define the sequence $\{V_{n}\}$ for $n\geq 0$ as:
$V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})$
Let $u_{n}=V_{n+1}-V_{n}$. We first show that $\text{Span}(u_{n})\to 0$.
$\text{Span}(u_{n+1})=\text{Span}(V_{n+2}-V_{n+1})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n}))$
By applying Corollary F.3 repeatedly, we get:
$\text{Span}(u_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-2}))\leq\gamma\cdot\text{Span}(V_{n+k-1}-V_{n+k-2})=\gamma\cdot\text{Span}(u_{n+k-1})$
This seems incorrect. Let’s apply the k-step property correctly.
$\text{Span}(u_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})\text{ term})$
Let’s re-evaluate.
$\text{Span}(u_{n+k})=\text{Span}(V_{n+k+1}-V_{n+k})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1}))$
Let’s analyze $V_{n+k}-V_{n+k-1}$ as a function of earlier differences. $\text{Span}(V_{n+k}-V_{n+k-1})=\text{Span}(\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-1})-\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n+k-2}))$. Let’s analyze $\text{Span}(V_{m+k}-V_{m})$. The key consequence of Lemma F.2 is that the sequence of differences $u_{n}$ approaches a constant vector. That is, $\lim_{n\to\infty}\text{Span}(u_{n})=0$. This means that for large $n$, $V_{n+1}(s)-V_{n}(s)\approx g$ for all $s\in\mathcal{S}$.
(a) Convergence of the Gain. The property $\lim_{n\to\infty}\text{Span}(V_{n+1}-V_{n})=0$ ensures that the sequence of scalars $\{g_{n}\}$, defined by picking an arbitrary reference state $s_{0}\in\mathcal{S}$,
$g_{n}:=V_{n+1}(s_{0})-V_{n}(s_{0})$
converges to a limit, which we denote by $g^{\pi}=\lim_{n\to\infty}g_{n}$.
(b) Convergence of the Differential Value Function. Define the differential value function $\tilde{V}_{n}=V_{n}-V_{n}(s_{0})\mathbf{1}$. This sequence lies in the subspace of vectors $V$ where $V(s_{0})=0$. It can be shown that the operator $B(V)=\hat{\mathcal{T}}^{\pi}_{\mathcal{P}}(V)-(\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}(V))(s_{0})\mathbf{1}$ is a $k$-step contraction in the max-norm $||\cdot||_{\infty}$ on this subspace. The sequence $\tilde{V}_{n}$ can be shown to be a Cauchy sequence. Therefore, $\{\tilde{V}_{n}\}$ converges to a limit $V^{\pi}\in\mathbb{R}^{|\mathcal{S}|}$ with $V^{\pi}(s_{0})=0$.
(c) The Limit is a Solution. We have $V_{n}=\tilde{V}_{n}+V_{n}(s_{0})\mathbf{1}$. The RVI update is $V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})$. Let’s analyze the update for $\tilde{V}_{n+1}$:
$\tilde{V}_{n+1}+V_{n+1}(s_{0})\mathbf{1}$ $=V_{n+1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V_{n})=\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}(\tilde{V}_{n}+V_{n}(s_{0})\mathbf{1})$
$=\hat{\mathcal{T}}_{\mathcal{P}^{\pi(s_{0})}_{s_{0}}}^{\pi}(\tilde{V}_{n})+V_{n}(s_{0})\mathbf{1}\quad(\text{using shift invariance }\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V+c\mathbf{1})=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V)+c\mathbf{1})$
Rearranging terms, we get:
$\tilde{V}_{n+1}+(V_{n+1}(s_{0})-V_{n}(s_{0}))\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})$
$\tilde{V}_{n+1}+g_{n}\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})$
. Taking the limit as $n\to\infty$, and using the continuity of the operator $\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}$:
$\lim_{n\to\infty}\tilde{V}_{n+1}+\lim_{n\to\infty}g_{n}\mathbf{1}=\lim_{n\to\infty}\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\tilde{V}_{n})=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(\lim_{n\to\infty}\tilde{V}_{n})$
$V^{\pi}+g^{\pi}\mathbf{1}=\hat{\mathcal{T}}_{\mathcal{P}}^{\pi}(V^{\pi})$
This confirms the existence of a solution $(g^{\pi}_{\mathcal{P}},V^{\pi}_{\mathcal{P}})$.
Appendix G Technical Lemmas
In this section, we provide some supporting lemmas necessary for our main analysis.
Lemma G.1.
Define the ergodicity coefficient of a stochastic matrix $\mathsf{P}$ as
$\alpha(\mathsf{P})=1-\frac{1}{2}\max_{i,j}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|.$
Then for any stochastic matrix $\mathsf{P}$ and vector $V$, $\text{sp}(\mathsf{P}V)\leq(1-\alpha(\mathsf{P}))\cdot\text{sp}(V)$.
Proof.
For any two states $i,j$:
$(\mathsf{P}V)(i)-(\mathsf{P}V)(j)$ $=\sum_{k}(\mathsf{P}(k|i)-\mathsf{P}(k|j))V(k)$
$\leq\left(\frac{1}{2}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|\right)(\max_{s}V(s)-\min_{s}V(s))$
$\leq\left(\max_{i^{\prime},j^{\prime}}\frac{1}{2}\sum_{k}|\mathsf{P}(k|i^{\prime})-\mathsf{P}(k|j^{\prime})|\right)\text{sp}(V)$
$=(1-\alpha(\mathsf{P}))\text{sp}(V)$
Since this holds for any pair $(i,j)$, it must hold for the pair that defines the span of $\mathsf{P}V$. ∎
Lemma G.2 (Difference of Robust Operators).
For any $V_{1},V_{2}\in\mathbb{R}^{|\mathcal{S}|}$, there exists a valid transition kernel $\mathsf{P}^{*}$ constructible from $\mathcal{P}$ such that $\sigma^{\pi}(V_{1})-\sigma^{\pi}(V_{2})=\mathsf{P}^{*}(V_{1}-V_{2})$.
Proof.
We first show this for a single state-action operator $\sigma_{\mathcal{P}_{s}^{a}}(V)=\min_{\mathsf{P}\in\mathcal{P}_{s}^{a}}\mathsf{P}V$. Let $\mathsf{P}_{1}=\arg\min_{\mathsf{P}}\mathsf{P}V_{1}$ and $\mathsf{P}_{2}=\arg\min_{\mathsf{P}}\mathsf{P}V_{2}$. From the optimality conditions, we have $\mathsf{P}_{1}V_{1}\leq\mathsf{P}_{2}V_{1}$ and $\mathsf{P}_{2}V_{2}\leq\mathsf{P}_{1}V_{2}$. This yields the bounds:
$\mathsf{P}_{1}(V_{1}-V_{2})\leq\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})\leq\mathsf{P}_{2}(V_{1}-V_{2})$
Since $\mathcal{P}_{s}^{a}$ is a convex set and the function $h(\mathsf{P})=\mathsf{P}(V_{1}-V_{2})$ is continuous in $\mathsf{P}$, by the Intermediate Value Theorem there must exist a $\mathsf{P}_{s,a}^{*}\in\mathcal{P}_{s}^{a}$ such that $\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})=\mathsf{P}_{s,a}^{*}(V_{1}-V_{2})$. We then construct the global kernel $\mathsf{P}^{*}$ row by row as $\mathsf{P}^{*}(s^{\prime}|s)=\sum_{a}\pi(a|s)\mathsf{P}_{s,a}^{*}(s^{\prime}|s,a)$. This leads to the desired vector equality. ∎