### v2

# Distributionally Robust Markov Games with Average Reward

Abstract
We study distributionally robust Markov games (DR‑MGs) with the average‑reward criterion, a crucial framework for multi-agent decision-making under uncertainty over extended horizons. We first establish a connection between the best-response policies and the optimal policies for the induced single‑agent problems. Under a standard irreducible assumption, we derive a correspondence between the optimal policies and the solutions of the robust Bellman equation, and derive the existence of stationary Nash Equilibrium (NE) based on these results. We also study a more general weakly communicating setting. We construct a set-valued map and show its value is a subset of the best-response policies, convex and upper hemi-continuous, which imply the existence of NE. We then introduce Robust Nash‑Iteration, and provide convergence guarantees. Finally, we connect average‑reward NE to discounted robust equilibria, showing approximation as the discount factor approaches one. Our studies provide comprehensive theoretical and algorithmic foundation for decision-making in complex, uncertain, and long-running multi-player environments.
1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments (shapley1953stochastic; filar2012competitive; littman1994markov). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the "Sim-to-Real" gap (mcmahan2024roping), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks (bukharin2023robust; ma2023decentralized). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
To address this vulnerability, the framework of distributionally robust Markov Games (DR-MGs) has been developed (zhang2020robust; kardecs2011discounted), extending the principles of distributionally robust Markov Decision Processes (MDPs) (bagnell2001solving; nilim2004robustness; iyengar2005robust) to the multi-agent setting. Instead of relying on a single, fixed MG model, the robust approach seeks to find an equilibrium that optimizes the worst-case performance over a predefined uncertainty set of plausible game models under potential model mismatches. The resulting robust equilibrium provides a performance guarantee across all models within this set, thereby ensuring each agent’s resilience against model mismatch.
The nature of a DR-MG problem is fundamentally shaped by the optimality criterion used. Much of the prior work has concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) or discounted-reward setting (zhang2020robust; kardecs2011discounted), where the objective is to maximize the expected reward over a finite number of steps or a discounted sum of rewards.
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics (krnjaic2024scalable), communication networks (qu2020scalable), autonomous vehicle coordination (peake2020multi), financial markets (vadori2024towards), or even peer-to-peer energy trading (qiu2021multi). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate (hernandez2000zero). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium (filar2012competitive); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions (puterman1994markov; atia2021steady). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
Moreover, none of the methods for DR-MGs under the discounted reward or finite horizon can be extended to the average-reward setting. The existence of (non-stationary) Nash Equilibrium under finite horizon is derived through backward induction (blanchet2023double), which does not apply to infinite-horizon settings. The existence of stationary Nash Equilibrium under the discounted reward criterion is derived in (kardecs2011discounted), which however heavily relies on the uniqueness of the solution to the discounted robust Bellman equation (iyengar2005robust) and hence cannot be applied to the average reward setting. Another natural attempt is to use the discounted reward as an intermediate step to study the robust Nash Equilibrium under the average reward by setting the discount factor approaching to $1$. Although this limit method is effective in the single-agent setting (wang2023robust), it cannot solve the DR-MGs, due to the complicated structure (also see our detailed discussion in Section 3). To date, the literature on average-reward DR-MGs remains sparse, leaving fundamental questions about the structure of robust equilibria and the design of convergent solution algorithms largely unanswered.
1.1 Contributions
In this paper, we develop a comprehensive study of DR-MGs under the average-reward formulation. Our contributions are summarized as follows.
Solvability of the Robust Bellman Equation under Irreducible Setting: As a foundational step, we first revisit the single-agent distributionally robust average-reward MDPs with the irreducible assumption. Despite existing studies largely utilizing the robust Bellman equation (wang2023robust; wang2023model; wang2024robust; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample), its solvability is not proved, but rather implicitly assumed. We establish the well-posedness of the robust average-reward Bellman equation, study the structure of its solutions, and connect them with the optimal robust policies. This analysis provides structural insights and tools that are essential for solving multi-agent Markov games when long-term performance is of interest.
Existence and Algorithmic Identification of Robust Nash Equilibrium: We further introduce a formulation of induced distributionally robust MDPs to reduce the Nash Equilibrium identification in comprehensive multi-agent games to optimizing single-agent distributionally robust MDPs. Under the irreducibility setting, we utilize the solvability results we obtained to show a correspondence between the optimal robust policy and solutions of the Bellman equation, and prove the convexity and hemi-continuity of the best-response set, deriving the existence of a stationary Nash Equilibrium. We further study the more general weak communicating setting, where no such correspondence can be obtained. We hence instead study a proxy mapping which is a subset of the best-response set, and derive the NE based on it. These results address the fundamental open question, demonstrating that stable solutions exist in these complex, robust, and long-term horizon competitive environments.
Convergence Algorithm: We moreover propose a practical, iterative algorithm—a robust Nash iteration scheme—designed to compute the robust NE in the average-reward setting. We provide a rigorous proof of the algorithm’s convergence to the robust equilibrium policies, offering the first provably convergent method for solving average-reward DR-MGs.
Connection to Discounted Equilibria: We also aim to bridge the gap between DR-MGs under the average reward and the well-studied discounted-reward settings. We demonstrate the closeness between the robust average reward NE and robust discounted NEs as the discount factor approaches one. This connection not only provides deeper insight into the nature of the average reward equilibrium, but also suggests potential approximation methods based on the discounted framework.
1.2 Related Works
We discuss the most related works in this section.
Non-robust Markov Games. Markov Games (littman1994markov) provide the foundational mathematical framework for multi-agent sequential decision-making. The majority of early work focused on the discounted-reward setting, where the existence of a stationary Nash Equilibrium was established in (shapley1953stochastic). The average-reward criterion, while more suitable for systems with long operational horizons, presents greater analytical challenges. Unlike the discounted case, a stationary NE is not guaranteed to exist in a general average-reward MG. Its existence has only been proven under additional structural assumptions on the dynamics of the game (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; filar2012competitive).
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, distributionally robust MDPs are first developed in (iyengar2005robust; nilim2004robustness) for both finite and discounted settings. The studies under the average reward setting are recently developed in (wang2023robust; wang2023model; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample; grand2023beyond; grand2023reducing; chatterjee2023solving; wang2024robust). While the robust Bellman equation is central to these methods, its solvability was often implicitly assumed without rigorous proof.
Distributionally Robust Markov Games. Extending distributional robustness to the multi-agent setting is a recent and active area of research. However, the literature has overwhelmingly concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) and discounted-reward criteria (zhang2020robust; kardecs2011discounted). These settings are often more analytically tractable because the influence of future rewards diminishes, simplifying the analysis. However, none of these methods can be extended to the average-reward setting, as we shall discuss later.
2 Preliminaries and Problem Formulation
[…]
A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is a distribution over $\mathcal{A}$ for any given state $s$, meaning the agent takes action $a$ at state $s$ with probability $\pi(a|s)$. The goal of an MDP is to maximize the accumulative reward. Since different reward criteria give rise to markedly distinct problem settings, the resulting objective functions can vary greatly, modifying the key properties necessary to analyze the underlying system. The more extensively studied criterion is the discounted reward, whose objective function is the discounted value function: $V^{\pi}_{\mathsf{P}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right]$, where $\gamma\in[0,1)$ is some discount factor. When long-term performance is of interest, the average reward is generally adopted as the criterion of choice. By effectively setting $\gamma=1$, it does not discount the reward over time, but considers the behavior of the underlying Markov process under the steady-state distribution. Under this formulation, the average reward induced by following policy $\pi$ starting from state $s\in\mathcal{S}$ is defined as $g_{\mathsf{P}}^{\pi}(s)\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\big[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\big].$
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (littman1994markov), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\otimes_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}_{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the actions of all agents. At some state $s$, all agents act simultaneously by taking joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on stationary Nash Equilibrium (NE). More specifically, a stationary NE is a product policy $\pi^{\star}=(\pi^{\star}_{1},...,\pi^{\star}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{\star}$:
$V^{\pi^{\star}}_{\mathsf{P},i}(\rho)\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathsf{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}\in\Pi_{i},$ (1)
for some initial distribution $\rho$ and $V^{\pi}_{\mathsf{P},i}(\rho)\triangleq\mathbb{E}_{s\sim\rho}[V^{\pi}_{\mathsf{P},i}(s)]$, and $\Pi_{i}$ denotes the set of stationary policies for agent $i$, $\pi^{\star}_{-i}$ denotes the joint policy of all other agents except $i$ from $\pi^{\star}$, and $(\pi^{\star}_{-i},\pi_{i})$ is the joint policy. The existence of stationary NE is derived for MGs with discounted reward (shapley1953stochastic) and average-reward MGs with additional structural assumptions (filar2012competitive).
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\otimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ (iyengar2005robust; nilim2004robustness; zhang2020robust; blanchet2023double; shi2024sample), where the uncertainty set is independently defined over all $(s,a)$ pairs.
After all agents take their actions $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}(\rho)$, and robust average reward $g^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}(\rho)$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{\star}$ such that for some initial state distribution $\rho$,
$g^{\pi^{\star}}_{\mathcal{P},i}(\rho)\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}.$ (2)
3 Hardness in Average Reward Nash Equilibrium Identification
Extending existing methods for discounted or finite-horizon robust MGs to the average-reward setting is prevented by two major challenges:
Additional hidden environment player. The first major difference between a vanilla non-robust MG and a DR-MG is the introduction of an additional player, which selects the transition kernel or the underlying game model to diminish the gains of the other players (zhang2020robust). This additional player fundamentally changes the structure of non-robust MGs. For example, in the extensively studied special case of two-player zero-sum MGs, where $r_{1}=-r_{2}=r$, one fundamental structure of standard MGs is the Min-Max duality:
$\max_{\pi_{1}}\min_{\pi_{2}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)=\min_{\pi_{2}}\max_{\pi_{1}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho),$ (3)
which directly implies the existence of a Nash Equilibrium (NE) (shapley1953stochastic). Such results are further extended to non-robust average-reward MG studies through the limit method, which sets the discount factor $\gamma\to 1$ (filar2012competitive; sennott1994zero; rieder1995average).
However, when looking at the distributionally robust setting for MGs, this notion fails due to consideration of the worst-case scenario. Since agents have different rewards, their corresponding worst-case transition kernels are also different. Hence, the duality fails (note that from agent $2$’s perspective, the worst case is $\max_{\mathsf{P}}V^{\pi}_{\mathsf{P},r}$ as it aims to minimize w.r.t. $r$):
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho).$ (4)
Consequently, even through the uniform convergence of the discounted robust value functions enables the interchanging of the order of $\max,\min,\lim$ (wang2023robust), it does not imply the existence of NE. And this approach cannot be applied in multi-player general-sum games.
Complicated structure of the average reward. The second major challenge stems from the inherent complexity of the (robust) average reward, which heavily depends on structures of the underlying MDPs. Firstly, even in the non-robust setting, Markov games with average reward do not necessarily admit a stationary NE (filar2012competitive; wikecek2015stationary; lozovanu2016stationary; thuijsman1997survey). And with model uncertainty, the robust average reward has a more complicated structure, including a non-linear robust Bellman equation (wang2023robust; wang2024modelfree) with non-unique solutions. The studies and understandings of robust average reward is also significantly limited compared to non-robust and discounted settings (wang2023robust; wang2023model; wang2025bellman). Therefore, despite the existence of the NE being proved for both discounted robust MGs (kardecs2011discounted; zhang2020robust) and average-reward non-robust MGs (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; wikecek2015stationary), these results and methods cannot be applied for the NE of average-reward DR-MGs.
To better illustrate the hardness of solving average reward DR-MGs, we first show the following impossibility result. The constructed DR-MG is essentially a multi-chain (puterman1994markov), indicating the necessity of the additional structural assumptions to solve average reward DR-MGs.
Lemma 3.1.
There exists a two-player DR-MG without any stationary NE under average reward.
4 NE of Irreducible DR-MGs
As discussed above, the existence of a NE cannot be guaranteed without any structural assumptions. In this section, we first study a relatively easier setting, where the underlying DR-MG is assumed to be irreducible.
Assumption 1.
For any deterministic joint policy $\pi$ and any transition kernel $\mathsf{P}\in\mathcal{P}$, the induced Markov chain $(\mathcal{S},\mathcal{A},\mathsf{P}^{\pi})$ is irreducible, i.e., all states belong to a single recurrent class. Moreover, $\mathcal{P}$ is compact and convex.
Remark 1.
Irreducibility is a standard assumption (slightly stronger than the unichain assumption (wang2023robust; wang2023model; wang2024robust; roch2025finite; Roch2025reduction)) in robust average reward studies (xu2025finite; xu2025efficient). Under this assumption, the robust average reward $g^{\pi}_{\mathcal{P}}(\rho)=g^{\pi}_{\mathcal{P}}(s)$ is always a constant over all initial states, ensuring the tractability of robust average reward. Hence in this section, we omit the initial state and denote the robust average reward by $g^{\pi}_{\mathcal{P}}$.
As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs. In non-robust average reward MG studies, the NE can be proved by studying the induced stationary distribution (wikecek2015stationary). However, when considering robustness, the worst-case stationary distribution set becomes much more complicated. For example, it is non-convex (wang2022robust; zhang2024distributionally; ma2025rectified), whereas the convexity is necessary in non-robust proofs and hence the method cannot be adapted. We thus propose a direct approach by studying the best response policy.
Definition 1.
(Best Response Policy) In an average-reward DR-MG, fix agent $i$ and the joint policy $\pi_{-i}$ of all other agents. The set of best response policies of agent $i$ w.r.t. $\pi_{-i}$ is defined as
$\text{BR}_{i}(\pi_{-i})\triangleq\left\{\pi_{i}:g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}=\max_{\mu_{i}}g^{(\pi_{-i},\mu_{i})}_{\mathcal{P},i}\right\}$. (5)
That is, the best response policy is agent $i$’s optimal policy when other agents fix their policies. This notion and the following results are standard and extensively studied in game theory literature (osborne2004introduction; fudenberg1991game):
Lemma 4.1.
$\pi^{\star}$ is a Nash Equilibrium if $\pi^{\star}_{i}\in\text{BR}_{i}(\pi^{\star}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{\star}$ is a fixed point of the BR mapping: $\pi^{\star}\in\text{BR}(\pi^{\star})$, where $\text{BR}(\pi)=\otimes_{i}\text{BR}_{i}(\pi^{\star}_{-i})$.
Lemma 4.2 (Kakutani’s Fixed Point Theorem (kakutani1941generalization)).
Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) $f(x)$ is nonempty and convex $\forall x$; (2) $f$ is upper hemi-continuous (i.e., for any convergent sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$, it holds that $y\in f(x)$). Then there exists some $x^{\star}\in X$ such that $x^{\star}\in f(x^{\star})$.
These two results imply that it is sufficient to show that the best response policy mapping, BR, satisfies the conditions in Lemma 4.2. We further show that these conditions can be further reduced to single-agent cases.
Lemma 4.3.
If for any agent $i\in\mathcal{N}$ and any stationary joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then the BR also satisfies the conditions in Lemma 4.2.
It is hence sufficient to fix an agent $i$ and study the best response policy set $\text{BR}_{i}(\cdot)$, which is a single-agent problem.
4.1 Induced Distributionally Robust MDPs
We then develop a framework of induced distributionally robust MDPs to provide a framework for multi-agent studies.
Definition 2 (Induced Distributionally Robust MDP).
Given a distributionally robust Markov game $\mathcal{MG}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{P},r)$ and an agent $i\in\mathcal{N}$. The induced distributionally robust MDP of the $\mathcal{MG}$ by a joint policy $\pi_{-i}$ is defined as $\mathsf{M}_{i}(\pi_{-i})\triangleq(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, where
$(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}\triangleq\Big\{\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)\mathsf{P}\Big|\mathsf{P}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}\Big\}$,
$r^{\pi_{-i}}(s,a_{i})=\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)r_{i}(s,(a_{i},a_{-i})).$ (6)
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, due to being induced, irreducibility, convexity, and compactness are all preserved. Namely, if the distributionally robust Markov game $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 1, then the induced distributionally robust MDP $\mathsf{M}_{i}(\pi_{-i})$ satisfies Assumption 1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
However, even the single-agent average reward robust RL is not fully studied. Specifically, the most essential tool, its robust Bellman equation, is not well studied: its solvability is not yet justified (an incorrect proof is given in (wang2023robust)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Although a huge body of existing work develops convergent algorithms, their convergence guaranties either rely on the wrong result in (wang2023robust), e.g., (wang2023model; Roch2025reduction), or bypass its solvability (xu2025finite; xu2025efficient; chatterjee2023solving; meggendorfer2025solving). We hence derive the solvability of the robust Bellman equation under Assumption 1 as follows.
Theorem 4.1.
(1). (Robust Bellman Equation). Under Assumption 1, denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the gain/bias of $(\pi,\mathsf{P}_{w}):$ $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is the unique solution to the equation:
$V(s)=\sum_{a}\pi(a|s)(r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)),$ (7)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$; Moreover, for any solution $(g,V)$ to (7), it holds that $g=g^{\pi}_{\mathcal{P}},V=h^{\pi}_{\mathsf{P}_{w}}+c\textbf{e}$, for some constant $c$ and $\textbf{e}=(1,...,1).$
(2). (Optimal Robust Bellman Equation). The optimal robust Bellman equation also has a solution:
$V(s)=\max_{a}\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\}.$ (8)
Our results therefore prove the solvability of the robust Bellman equation, solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (7) can have non-unique solutions even up to constant vectors. Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward (puterman1990markov), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.
These results and connections are of fundamental importance in showing the existence of a Nash Equilibrium.
4.2 Existence of a Nash Equilibrium
With the results derived for single-agent robust Bellman equations, we can then show the following results.
Theorem 4.2.
For any distributionally robust MDP, its optimal robust policy set $\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$ is convex.
As $\text{BR}_{i}(\pi_{-i})$ is the set of optimal robust policies of the (induced) distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, this result hence directly implies the convexity of $\text{BR}_{i}(\pi_{-i})$. We highlight that this proof is not a direct extension from its counterpart under the discounted reward as we need to tackle the non-uniqueness of solutions, whereas the discounted Bellman equation admits a unique solution.
We further show the hemi-continuity of the optimal policy set of the induced distributionally robust MDP.
Theorem 4.3.
For convergent policy sequences $\{\pi^{n}_{-i}\}\to\pi_{-i}$ and $\{\mu^{n}_{i}\}\to\mu_{i}$, if $\mu^{n}_{i}$ is an optimal robust policy of the induced distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi^{n}_{-i}},r^{\pi^{n}_{-i}})$ for agent $i\in\mathcal{N}$, then $\mu_{i}$ is also optimal for $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$.
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, thus resulting in a distributionally robust MDP sequence with different dynamics. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
Remark 2.
The proofs of these results are based on the correspondence between the optimal policies and the solutions to the Bellman equations, heavily relying on Theorem˜4.1. In the next section when we study a weaker setting, such correspondence does not hold anymore and the proof derived cannot be applied there.
Combining all of the above results, we can finally show the existence of a Nash Equilibrium.
Theorem 4.4.
(Existence of a Nash Equilibrium) For an average-reward distributionally robust Markov game satisfying Assumption 1, a stationary Nash equilibrium exists.
Our result herein implies the solvability of DR-MGs in the average-reward setting. The existence of an equilibrium is the most fundamental property of a game, thus proving its existence is a necessary prerequisite for any further analysis or algorithmic design.
5 NE of Weakly Communicating DR-MGs
In this section, we further extend our studies to a more general setting of weakly communicating DR-MGs.
Assumption 2 (Weakly Communicating DR-MG).
$\mathcal{P}$ is convex and compact. For any transition kernel $\mathsf{P}\in\mathcal{P}$, and for any player $i$, the state space can be partitioned into two distinct sets $\mathcal{S}=\mathcal{S}_{0}^{i}\cup\mathcal{S}_{1}^{i}$, such that (1). Any state in $\mathcal{S}_{0}^{i}$ is transient under every stationary joint policy; And (2). For any two states $s,s^{\prime}\in\mathcal{S}_{1}^{i}$, and any policy $\pi_{-i}$, there exists some policy $\pi_{i}$ and $N$ such that $\mathbb{P}(S_{N}=s^{\prime}|S_{0}=s,(\pi_{i},\pi_{-i}),\mathsf{P})>0$.
This assumption is adapted from the non-robust settings (wikecek2015stationary; bertsekas2011dynamic). The weakly communicating assumption is strictly weaker than ˜1, as it only requires accessibility under some policy, instead of every policy. Moreover, under this assumption, only the optimal robust average reward $g^{\star}_{\mathcal{P}}(s)$ is a constant over initial states, whereas $g^{\pi}_{\mathcal{P}}(s)$ is not.
The solvability of the optimal robust Bellman equation is derived recently in (wang2025bellman), as follows.
Lemma 5.1.
(wang2025bellman) Under ˜2, (8) has a solution $(g^{\star},u^{\star})$, and $g^{\star}=g^{\star}_{\mathcal{P}}(\rho)$.
The results are strictly weaker than the ones we derived for irreducible DR-MGs. Firstly, Theorem˜4.1 ensures the solvability of the robust Bellman equation with any policy, whereas Lemma˜5.1 only consider the optimal equation. Hence, a correspondence between the optimal robust policy and solutions of (8) cannot be developed as in irreducible settings, and the previous proofs cannot be applied to show the convexity and continuity of the best-response mapping.
To address this issue, we instead study another mapping specified through the optimal robust Bellman equation: for $s\in\mathcal{S}$, $\nu\in\Delta(\mathcal{A}_{i})$, and a joint policy $\pi_{-i}$, we define
$\Psi_{s}(\nu)\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\nu(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+\sigma_{(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}}(u^{\star})\big),$
and set a map as the solution set of the Bellman equation:
$\mathcal{G}_{i}(\pi_{-i})\ \triangleq\ \bigl\{\nu:\Psi_{s}(\nu(\cdot|s))=\max_{\mu\in\Delta(\mathcal{A}_{i})}\Psi_{s}(\mu),\forall s\in\mathcal{S}\bigr\}.$
Based on the solvability result in Lemma˜5.1, we can show that $\mathcal{G}_{i}$ is nonempty. Different from irreducible cases, due to the existence of transient states, it is not guaranteed that an optimal policy can be obtained from a solution to the Bellman equations, hence $\mathcal{G}_{i}(\pi_{-i})$ may not contain all best-response policies. However, we can show that $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$, which is sufficient to derive the existence of the NE. Further, if the map $\mathcal{G}(\pi)=\times_{i}\mathcal{G}_{i}(\pi_{-i})$ satisfies the conditions in Lemma˜4.2, it then admits a fixed point which is also a fixed point under the BR mapping, and hence implies the existence of NE. Our results are presented in the following theorem.
Theorem 5.1.
Under ˜2, the following statements hold. (1). $\mathcal{G}_{i}$ is nonempty, compact, and convex; (2). $\mathcal{G}_{i}$ is upper hemi-continuous; (3). $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$; (4). A DR-MG always has a stationary NE.
The proofs are deferred to Appendix˜D, which rely on the strong duality of the robust average reward shown in (wang2025bellman), and are fundamentally different from the ones for Section˜4.
6 Algorithmic Solutions
We then develop an algorithmic solution for average-reward DR-MGs. Notably, finding stationary NE is significantly challenging in general-sum games, even under non-robust settings (hu2003nash; li2007reinforcement; li2003learning; zinkevich2005cyclic), and existing convergence results generally require additional structural assumptions on the underlying games. In this section, we design a value iteration based algorithm in Algorithm 1, named robust Nash-Iteration, to find a NE. The convergence guarantee of our algorithm, similarly to previous ones for general-sum games, is only obtained under some additional assumptions. Our algorithm is designed for irreducible DR-MGs, as learning under weakly communicating settings is generally challenging and remains open (sahabandu2024rl).
Algorithm 1 Robust Nash-Iteration
Initialize: $h_{i}=0,\forall i\in\mathcal{N}$
while TRUE do
  for all $i\in\mathcal{N}$ do
   $h^{0}_{i}\leftarrow h_{i}$
   for all $s,a\in\mathcal{S}\times\mathcal{A}$ do
    $Q_{i}(s,a,h_{i}^{0})\leftarrow r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i}^{0})$
    $\pi(s)\leftarrow\textbf{NE}(Q_{i}(s,a,h^{0}_{i}))$
    $h_{i}(s)\leftarrow\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i}^{0})]$
   end for
  end for
  if $\max_{i}\{\text{sp}(h_{i}-h^{0}_{i})\}=0$ then
   BREAK
  end if
end while
return $\pi$
Our algorithm maintains an estimation, $h_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). This assumption is not merely a technical convenience; it is the conceptual linchpin that stabilizes the learning dynamics of the system. By placing constraints on which Nash Equilibrium the oracle returns, we ensure that the sequence of generated policies and value functions behaves predictably, guiding the algorithm steadily toward a solution instead of allowing it to wander chaotically or diverge.
Theorem 6.1.
Under ˜1, and with a specific equilibrium selection oracle NE (see Assumption 3 in Appendix), the robust Nash-Iteration algorithm converges to a Nash Equilibrium.
This theorem provides the algorithmic contribution of our work, demonstrating that the proposed Robust Nash-Iteration algorithm is a principled method with guaranteed convergence. Given the known difficulties associated with multi-agent learning algorithms in general-sum games under the long-term performance criterion, our algorithm stands as a significant contribution. Taken together, our results provide a complete theoretical foundation for this problem. From establishing the existence of a robust equilibrium in Theorem 4.4 to now providing a concrete algorithm to find it, this work makes the class of average-reward DR-MGs amenable to computational analysis. It provides a blueprint for developing practical, reliable multi-agent systems that can find stable operating points even when faced with significant environmental uncertainty.
7 Connections with Discounted DR-NE
In this section, we study the connection between distributionally robust games with discounted and average rewards. Our motivation is to utilize the extensively studied and mathematical elegance of the discounted setting to solve the more challenging average-reward setting. Recall that a policy $\pi^{\star}_{\gamma}$ is a Nash Equilibrium of the $\gamma$-discounted DR-MG if
$V^{\pi^{\star}}_{\mathcal{P},i}\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (9)
where $V^{\pi^{\star}}_{\mathcal{P},i}$ is the $\gamma$-discounted robust value function. A policy $\pi^{\star}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{\star}}_{\mathcal{P},i}+\epsilon\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (10)
We then derive the following results for irreducible DR-MGs (˜1).
Theorem 7.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
This theorem establishes a crucial bridge between the discounted and average-reward criteria in DR-MGs. Specifically, we showed that the worst-case long-run average behavior of a multi-agent system can be approximated by its discounted behavior when the future is valued almost as much as the present (i.e., as $\gamma\to 1$).
The importance of this theorem lies in its computational implications. The analysis of $\gamma$-discounted games is often more tractable (kardecs2011discounted), benefiting from properties like the Banach fixed-point theorem which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as value iteration and policy iteration, are naturally formulated for the discounted setting. Our results hence provide a rigorous justification for using these well-understood tools to solve the more challenging average reward problem.
Remark 3.
Part (2) of Theorem 7.1 offers a more practical and constructive guarantee. In practice (like data-driven learning), it may be sufficient to find some approximation of the NE instead of the exact NE. This result therefore assures us that for any desired level of accuracy $\epsilon$, we can simply select a single discount factor $\gamma$ sufficiently close to 1. Any $\epsilon$-Nash Equilibrium found for this $\gamma$-discounted game will be an $\epsilon$-Nash Equilibrium for the average-reward game. This provides a clear and direct pathway for finding near-equilibrium policies for the average-reward criterion. Moreover, we can derive the following result to specify the reduction to discounted DR-MGs.
Lemma 7.1.
Define the robust diameter (auer2008near) of the distributionally robust Markov game by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
The result hence specifies a reduction factor whose Nash Equilibrium is an approximated equilibrium under the average reward criterion. Thus, we can apply existing algorithms for discounted DR-MGs (zhang2020robust) to approximately solve the average reward ones.
8 Numerical Experiments
In this section, we develop numerical experiments to validate the advantages of our formulation and methods. We mainly consider two aspects: the long-term performance and the robustness to environment uncertainty.
8.1 Long-Term Performance
To empirically validate our approach, we perform a series of experiments designed to compare the performance of our robust Nash-Iteration algorithm with average reward against the standard discounted-reward robust Nash-Iteration (zhang2020robust; shi2024sample) under DR-MGs. We aim to show that our algorithm will yield a joint policy that has better long-term performance than those derived from the discounted criterion.
We design a "Structured Random Environment" to create a complex strategic challenge with meaningful trade-offs between short-term gains and long-term consequences. The environment is a DR-MG with 20 states and 5 actions. The state space $S$ is partitioned into two clusters:
  • Prosperous States ($|S_{p}|=5$): A small set of states characterized by a higher mean reward (drawn from $\mathsf{N}(2,1)$).
  • Deprived States ($|S_{d}|=15$): A larger set of states with a lower mean reward (drawn from $\mathsf{N}(-2,1)$),
where $\mathsf{N}(\cdot,\cdot)$ is the Gaussian distribution. The nominal transition kernel $\mathsf{P}_{0}$ is also structured to create persistence within these clusters. Transitions originating from a state in a given cluster are five times more likely to terminate in a state within the same cluster. This design creates a strategic trap: an agent in a prosperous state might be tempted by an action with a high immediate reward that carries a significant risk of transitioning it to the deprived cluster, from which escape is difficult.
We further set the uncertainty set using a Kullback-Leibler (KL) divergence ball around the nominal kernel $\mathsf{P}_{0}$:
$\mathcal{P}^{a}_{s}=\{q\in\Delta(\mathcal{S})\ |\ D_{KL}(q||(\mathsf{P}_{0})^{a}_{s})\leq\theta\}.$
Based on the scale of our environment, the uncertainty budget $\theta$ is set to a modest value of 0.01.
We compare our average-reward Robust Nash-Iteration algorithm (Algorithm 1) and the Discounted-Reward Robust Nash-Iteration. We implement the two algorithms and evaluate their output policies under the robust average-reward criterion. The experimental procedure is as follows:
(1). The optimal average-reward policy, $\pi_{avg}$, is computed using the average-reward Robust Nash-Iteration algorithm. We evaluate the worst-case average reward of Player 1 under the policy, using the robust average reward single-agent RL method in (wang2023robust).
(2). For a range of discount factors $\gamma$ from 0.5 to 0.99, the corresponding optimal discounted policy, $\pi_{disc}(\gamma)$, is computed using the Discounted-Reward Robust Nash-Iteration algorithm. And for each $\pi_{disc}(\gamma)$, we similarly evaluate its worst-case long-term performance.
For both algorithms, we plot the performance of Player 1 as a performance metric. Results are presented in Figure 1 (under different random environments). Policies learned under the discounted criterion consistently underperform the dedicated average-reward policy. This gap is most pronounced for smaller values of $\gamma$, where the myopic nature of the discounted algorithm leads it to fall into the strategic trap of the environment by choosing actions that yield high immediate rewards but lead to the persistent, low-reward "deprived" cluster of states. As $\gamma$ approaches 1, the discounted policy becomes more far-sighted, and its performance converges toward the optimal baseline.
Our results hence provide strong evidence that, for DRMGs under the average reward criterion with long-term performance optimization goals, our robust Nash-Iteration algorithm achieves a better performance, compared to the common practice of approximating this objective with a discounted formulation using $\gamma\approx 1$.
Figure 1: Performance Comparisons of Average and Discounted Robust Nash-Iteration
Figure 2: Performance Comparisons of Robust/Non-Robust Nash-Iteration
8.2 Robustness
We then test the robustness of our method by comparing it with the non-robust average-reward Nash-Iteration algorithm (li2003learning), to test their ability to handle model uncertainty. Using the same environment, we evaluate both learners against a worst-case adversary throughout training to measure their inherent robustness. The results are presented in Figure 2. The policy learned by the robust algorithm consistently and monotonically improves its worst-case performance, rapidly converging to a high, stable value. This is because the algorithm directly optimizes for resilience against the adversary at each step. In contrast, the policy learned by the non-robust algorithm exhibits lower performance when evaluated against a worst-case adversary. Ultimately, it does not have the ability to reliably account for perturbations in the underlying environment, as opposed to its robust counterpart, which aligns with our theoretical results.
These numerical results therefore verify that our method is able to achieve superior long-term performance with increased stability when in the face of model uncertainty in multi-agent systems, validating our theoretical results.
9 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We established the well-posedness of the robust average-reward Bellman equation, proved the existence of a stationary Nash Equilibrium under both irreducible and weak communicating settings, and proposed the Robust Nash-Iteration algorithm as the first provably convergent method for this game. We also connected average-reward equilibria to their discounted counterparts, and our numerical experiments validated the superior long-term performance and stability of our approach. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
Appendix A Preliminaries
In this section we briefly review previous studies of single-agent distributionally robust MDPs.
A.1 Discounted Setting
Under the robust setting, the worst-case performance over the uncertainty set of MDPs is defined through the discounted reward. More specifically, the robust discounted value function of a policy $\pi$ for a discounted MDP is defined as
$V^{\pi}_{\mathcal{P},\gamma}(s)\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\mathbb{E}_{\pi,\kappa}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right],$ (11)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{t\geq 0}\mathcal{P}$.
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator (nilim2004robustness; iyengar2005robust; puterman1994markov):
$\mathbf{T}_{\pi}V(s)\triangleq\sum_{a\in\mathcal{A}}\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right),$ (12)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{p\in\mathcal{P}^{a}_{s}}p^{\top}V$ is the support function of $V$ on $\mathcal{P}^{a}_{s}$. Based on the contraction of $\mathbf{T}_{\pi}$, robust dynamic programming approaches, e.g., robust value iteration, can be designed (nilim2004robustness; iyengar2005robust).
The goal is to find the optimal robust policy:
$\pi^{\star}=\arg\max_{\pi}V^{\pi}_{\mathcal{P},\gamma}(s),$ (13)
and it is shown (iyengar2005robust) that the optimal robust value function $V^{\pi^{\star}}_{\mathcal{P},\gamma}(s)$ satisfies the optimal robust Bellman equation
$V(s)\triangleq\max_{a}\left\{\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right)\right\}.$ (14)
A.2 Average Reward Setting
For a fixed kernel $\mathsf{P}_{0}$, the average reward (or gain) and the bias are defined as
$g_{\mathsf{P}_{0}}^{\pi}(s)$ $\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right]$ (15)
and
$h_{\mathsf{P}_{0}}^{\pi}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\sum^{\infty}_{t=0}(r_{t}-g_{\mathsf{P}_{0}}^{\pi})|s_{0}=s\right].$ (16)
We can view equation (16) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by (puterman1994markov), thus it holds the following non-robust Bellman equation:
$h^{\pi}_{\mathsf{P}_{0}}(s)=\mathbb{E}_{\pi}\left[r(s,a)-g^{\pi}_{\mathsf{P}}(s)+\sum_{s^{\prime}\in\mathcal{S}}p^{a}_{s,s^{\prime}}h_{\mathsf{P}_{0}}^{\pi}(s^{\prime})\right].$ (17)
In the robust setting, we consider the robust average reward, defined as
$g_{\mathcal{P}}^{\pi}(s)$ $\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\liminf_{n\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right].$ (18)
It is shown that under the unichain assumption, the robust average reward is independent from the initial state: $g_{\mathcal{P}}^{\pi}(s_{1})=g_{\mathcal{P}}^{\pi}(s_{2})$, for any $s_{1},s_{2}$. The robust Bellman equation (for a policy $\pi$) and the robust optimal Bellman equation for the average reward setting are derived in (wang2023robust):
$V(s)+g$ $=\sum_{a}\pi(a|s)\big(r(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(V)\big),$ (19)
$V(s)$ $=\max_{a}\big\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big\},\quad\forall s\in\mathcal{S}.$ (20)
It is shown (wang2023model) that if the equation (19) ((20)) has a solution $(g,V)$, then the solution $g$ is the robust average reward $g^{\pi}_{\mathcal{P}}$ (the optimal robust average reward $g^{*}_{\mathcal{P}}=\max_{\pi}g^{\pi}_{\mathcal{P}}$). However, whether these two equations are solvable is not studied but directly assumed.
Appendix B Proof of Lemma˜3.1
We derive the result by constructing an example.
Fix $\varepsilon\in(0,1)$. Consider a two-player zero-sum distributionally robust Markov game
$(\mathcal{S},\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{P},r_{1},r_{2}),$
with state space $\mathcal{S}=\{s,\,G,\,B\}$, action sets $\mathcal{A}_{1}=\{1,2\}$ and $\mathcal{A}_{2}=\{1,2\}$, and rewards
$r_{1}(G,a_{1},a_{2})=1,\qquad r_{1}(B,a_{1},a_{2})=0,\qquad r_{1}(s,a_{1},a_{2})=0,\qquad r_{2}=-r_{1}.$
The ambiguity set consists of two kernels $p^{(1)}$ and $p^{(2)}$:
$p^{(1)}(\,\cdot\,|s,(1,1))$ $=\delta_{G},$ $p^{(1)}(\,\cdot\,|s,(2,2))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,1))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,2))$ $=\delta_{G},$
$p^{(1)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(1)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
and, for $x\in\{G,B\}$, $p^{(1)}(\,\cdot\,|x,(a_{1},a_{2}))=p^{(2)}(\,\cdot\,|x,(a_{1},a_{2}))=\delta_{x}$ (absorbing). Thus, under any $p\in\{p^{(1)},p^{(2)}\}$ the chain has two closed recurrent classes $\{G\}$ and $\{B\}$ (i.e., it is multi-chain).
Let the initial state be $S_{0}=s$. For stationary policies $\pi_{1}\in\Delta(\mathcal{A}_{1})$, $\pi_{2}\in\Delta(\mathcal{A}_{2})$, we specify them as
$\pi_{1}(1)=p\in[0,1],\qquad\pi_{2}(1)=q\in[0,1].$
Because $G,B$ are absorbing and $r_{1}\in\{0,1\}$ on $\{G,B\}$, the average reward of player 1 equals the probability (under the chosen kernel) of transitioning from $s$ to $G$ in one step. Thus we have that
$U(p,q)$ $\triangleq\min_{k\in\{1,2\}}\ \mathbb{P}_{p^{(k)}}\big(S_{1}=G\,\big|\,S_{0}=s,\,\pi_{1},\pi_{2}\big)$
$=\min\big\{\,p\,q+\varepsilon(1-p)(1-q),\ \varepsilon\,p\,q+(1-p)(1-q)\,\big\}.$
Then the robust average reward for player 1 is $U(p,q)$ and for player 2 is U(p,q)-U(p,q).
Lemma B.1.
For $U$ defined above and any $\varepsilon\in(0,1)$,
$\sup_{p\in[0,1]}\ \inf_{q\in[0,1]}U(p,q)\ =\ \frac{\varepsilon}{2},\qquad\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)\ =\ \varepsilon.$
Proof.
Fix $p\in[0,1]$. For $q\in[0,1]$, $U(p,q)$ is the pointwise minimum of two affine functions of $q$, hence concave in $q$. Therefore the infimum over $q$ is attained at an endpoint:
$\inf_{q\in[0,1]}U(p,q)=\min\{\,U(p,0),\,U(p,1)\,\}=\min\{\varepsilon(1-p),\,\varepsilon p\}=\varepsilon\min\{p,1-p\}.$
Maximizing over $p$ gives
$\sup_{p\in[0,1]}\inf_{q\in[0,1]}U(p,q)=\sup_{p\in[0,1]}\varepsilon\min\{p,1-p\}=\varepsilon\cdot\frac{1}{2}=\frac{\varepsilon}{2}.$
Fix $q\in[0,1]$. As a function of $p$, $U(p,q)$ is again the pointwise minimum of two affine functions, hence concave and piecewise linear with a single kink at the solution of equality:
$p\,q+\varepsilon(1-p)(1-q)=\varepsilon p\,q+(1-p)(1-q)\ \Longleftrightarrow\ p=1-q.$
Thus
$\sup_{p\in[0,1]}U(p,q)=\max\Big\{\,U(0,q),\ U(1,q),\ U(1-q,q)\,\Big\}=\max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
Hence
$\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)=\min_{q\in[0,1]}\ \max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
For $q\in\{0,1\}$ the maximum equals $\varepsilon$. For $q\in(0,1)$, $(1+\varepsilon)q(1-q)>\varepsilon\min\{q,1-q\}$ and thus the inner maximum is at least $\max\{\varepsilon q,\varepsilon(1-q)\}\geq\varepsilon/2$, while near $q=0$ or $1$ it is strictly larger than $\varepsilon/2$. The minimum over $q$ is attained at $q=0$ or $q=1$, and equals $\varepsilon$. ∎
Theorem B.1 (Nonexistence of a Nash equilibrium).
For the robust Markov game above, no (stationary mixed) Nash equilibrium exists.
Proof.
In a two-player zero-sum game, a stationary Nash equilibrium $(p^{\star},q^{\star})$ is a saddle point:
$U(p^{\star},q)\ \geq\ U(p^{\star},q^{\star})\ \geq\ U(p,q^{\star})\qquad\forall\,p,q\in[0,1],$
which implies
$\inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star}).$
Consequently,
$\sup_{p}\inf_{q}U(p,q)\ \geq\ \inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star})\ \geq\ \inf_{q}\sup_{p}U(p,q),$
so a saddle requires $\sup_{p}\inf_{q}U=\inf_{q}\sup_{p}U$. By Lemma B.1, $\sup_{p}\inf_{q}U=\varepsilon/2$ while $\inf_{q}\sup_{p}U=\varepsilon$, leading to a contradiction. Hence no saddle point exists and no Nash equilibrium exists. ∎
Appendix C Irreducible DR-MGs
Theorem C.1.
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the robust Bellman equation:
$V=r^{\pi}-g+\sigma^{\pi}(V)$ (21)
has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$.
Proof.
Note that for $\mathsf{P}_{w}$, it holds the corresponding non-robust Bellman equation:
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (22)
We further denote that
$\mathsf{P}_{V}=\arg\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}},$ (23)
then it also holds that
$V^{\pi}_{\mathsf{P}_{V}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{V}}.$ (24)
Combining (22) and (24) implies that
$(\mathsf{P}^{\pi}_{w}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}}.$ (25)
By the definition of $\mathsf{P}_{V}$, it holds that $\mathsf{P}_{V}^{\pi}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}_{w}^{\pi}V^{\pi}_{\mathsf{P}_{w}}$, thus
$(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{w}}\leq(\mathsf{P}^{\pi}_{V}-I)V^{\pi}_{\mathsf{P}_{V}},$ (26)
and it holds that
$\mathsf{P}^{\pi}_{V}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\leq(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}).$ (27)
Recursively applying it further implies that
$(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}})\geq(\mathsf{P}^{\pi}_{V})^{k}(V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}),\forall k.$ (28)
By denoting $V^{\pi}_{\mathsf{P}_{w}}-V^{\pi}_{\mathsf{P}_{V}}$ by $x$, the inequality above can be rewritten as
$x\geq(\mathsf{P}^{\pi}_{V})^{k}x,\forall k.$ (29)
We denote $s^{\star}\triangleq\arg\min_{s\in\mathcal{S}}x(s)$, then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s).$ (30)
By our assumption, for any $s^{\prime}\in\mathcal{S}$, there exists some $k$ such that $((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}>0$. Then it holds that
$x(s^{\star})\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq\sum_{s\in\mathcal{S}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s^{\star})=x(s^{\star}),$ (31)
which implies that
$x(s^{\star})=((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+\sum_{s\neq s^{\prime}}((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s}x(s)\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime})+(1-((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}})x(s^{\star}).$ (32)
We then have that
$((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\star})\geq((\mathsf{P}^{\pi}_{V})^{k})_{s^{\star},s^{\prime}}x(s^{\prime}),$ (33)
which implies $x(s^{\star})=x(s^{\prime}),\forall s^{\prime}\in\mathcal{S}$, and thus $x=me$ for some constant $m$, i.e., $V^{\pi}_{\mathsf{P}_{w}}=me+V^{\pi}_{\mathsf{P}_{V}}$.
Now we plug it in (22) and (24), and it holds that
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$ (34)
$V^{\pi}_{\mathsf{P}_{w}}$ $=r^{\pi}-g^{\pi}_{\mathsf{P}_{V}}+\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}},$ (35)
which implies that
$g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (36)
Note that $g^{\pi}_{\mathsf{P}_{w}}\leq g^{\pi}_{\mathsf{P}_{V}}$, due to $\mathsf{P}_{w}$ is the worst-case kernel w.r.t. $g^{\pi}_{\mathsf{P}}$, and $\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}\leq\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}$, due to the definition of $\mathsf{P}_{V}$, thus
$0\leq g^{\pi}_{\mathsf{P}_{V}}-g^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}-\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}\leq 0.$ (37)
It further implies that
$\mathsf{P}^{\pi}_{V}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}},$
i.e.,
$\sigma^{\pi}({V^{\pi}_{\mathsf{P}_{w}}})=\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}V^{\pi}_{\mathsf{P}_{w}}=\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}.$ (38)
Combine it with (22) and we have that
$V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\mathsf{P}^{\pi}_{w}V^{\pi}_{\mathsf{P}_{w}}=r^{\pi}-g^{\pi}_{\mathsf{P}_{w}}+\sigma^{\pi}(V^{\pi}_{\mathsf{P}_{w}}).$ (39)
Hence the Bellman equation (7) has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$, which completes the proof. ∎
Corollary C.1.
The optimal robust Bellman equation
$V(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}^{a}_{s}}(V)\}$ (40)
is solvable.
Proof.
Denote the optimal robust policy by $\pi^{\star}$, and the corresponding worst-case kernel by $\mathsf{P}^{\star}$. By Theorem 4.1, $(g^{\pi^{\star}}_{\mathcal{P}},h^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is the solution to the robust Bellman equation of $\pi^{\star}$:
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=r^{\pi^{\star}}(s)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{\pi^{\star}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right).$ (41)
We now set $\pi^{\prime}(s)=\arg\max_{a}\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\},\forall s$, then we have that
$r^{\pi^{\prime}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})=g^{\pi^{\prime}}_{\mathcal{P}}\leq g^{\pi^{\star}}_{\mathcal{P}}$ $=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right)$
$\leq r^{\pi^{\prime}}(s)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}),$ (42)
which further implies that
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)\leq\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})-\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})\leq\mathsf{P}^{\prime}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}).$ (43)
Now denote $x=h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}$, it holds that
$x\leq\mu x,$ (44)
for some positive distribution matrix $\mu$. Similarly, it implies $x(s)=x(s^{\prime})$, and thus $h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}=ce$. This further implies that all inequalities in (42) are equations, thus
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=\max_{a}\left\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right\},$ (45)
hence $(h^{\pi^{\star}}_{\mathsf{P}^{\star}},g^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is a solution to (8), completing the proof of the first part. The remaining proof follows in a similar way from (wang2023model). ∎
Lemma C.1.
If for any agent $i$ and any joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then BR also satisfies the conditions in Lemma 4.2.
Proof.
(1). If for any $i$ and $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, then $\text{BR}(\pi)=(\text{BR}_{1}(\pi_{-1}),...,\text{BR}_{N}(\pi_{-N}))$ is clearly nonempty.
(2). Consider two policies $\pi^{1},\pi^{2}\in\text{BR}(\pi)$. To show the convexity of $\text{BR}(\pi)$, we need to show $\pi^{\prime}=\alpha\pi^{1}+(1-\alpha)\pi^{2}=(\alpha\pi^{1}_{1}+(1-\alpha)\pi^{2}_{1},...,\alpha\pi^{1}_{N}+(1-\alpha)\pi^{2}_{N})\in\text{BR}(\pi)$. Note that for any $i$, $\pi^{1}_{i},\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, and due to the convexity of $\text{BR}_{i}(\pi_{-i})$, $\alpha\pi^{1}_{i}+(1-\alpha)\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, which completes the proof.
(3). Consider two convergent policy sequences $\{\pi^{n}\}\to\pi$ and $\{\mu^{n}\}\to\mu$, with $\mu^{n}\in\text{BR}(\pi^{n})$, we need to show that $\mu\in\text{BR}(\pi)$. Note that $\pi^{n}\to\pi$ implies that $\pi^{n}_{i}\to\pi_{i}$ for any $i$, and similarly $\mu^{n}\to\mu$. Since $\mu_{i}^{n}\in\text{BR}_{i}(\pi_{-i}^{n})$, by the hemi-continuity of $\text{BR}_{i}$, we have $\mu_{i}\in\text{BR}_{i}(\pi_{-i})$. Moreover note that
$\mu\in\text{BR}(\pi)\Leftrightarrow\mu_{i}\in\text{BR}_{i}(\pi_{-i}),\forall i,$ (46)
which hence completes the proof. ∎
C.1 Convexity
Lemma C.2.
Consider the optimal robust Bellman equation
$g+h(s)=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (47)
and denote any of its solutions by $(g,h)$. Additionally, consider an optimal robust policy $\pi$. Then $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\pi$:
$g+h(s)=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ (48)
Proof.
From Theorem 4.1, (48) is solvable, and we denote one of its solutions by $(g^{\pi},h^{\pi})$. Then we have that
$g+h(s)$ $=\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\},$ (49)
$g^{\pi}+h^{\pi}(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}.$ (50)
Now from Theorem 4.1, $g=g^{\pi}=g^{\star}_{\mathcal{P}}$, thus it holds that
$h^{\pi}(s)-h(s)$ $=\sum_{a}\pi(a|s)\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})\}-\max_{a}\{r(s,a)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq\sum_{a}\mu(a|s)\{\sigma_{\mathcal{P}_{s}^{a}}(h^{\pi})-\sigma_{\mathcal{P}_{s}^{a}}(h)\}$
$\leq q_{s}^{\mu}(h^{\pi}-h),$ (51)
for some policy $\mu$ and some kernel $q\in\mathcal{P}$. By the irreducibility of $q^{\mu}$, $h^{\pi}-h=ce$. Thus $(g,h)=(g^{\pi},h^{\pi}+ce)$ is also a solution to (48). ∎
Based on the equivalence between the best response policy set and the optimal policy set of the induced robust MDP, the convexity can be derived from the following result.
Lemma C.3.
Define $\Pi^{\star}\triangleq\{\pi\ |\ \text{optimal policies of }g^{\pi}_{\mathcal{P}}\}$. Then $\Pi^{\star}$ is convex.
Proof.
Let $\pi_{1},\pi_{2}\in\Pi^{\star}$ and define $\tilde{\pi}=\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1].$ Then by Theorem 4.1,
[…]
Alternatively, let $(g,h)$ be a solution to the optimal equation, i.e. $h(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}_{s}^{a}}(h)\}.$ From Lemma C.2, this $(g,h)$ also solves the equation w.r.t. $\pi_{1},\pi_{2}$, i.e.
$g+h=r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\quad\text{and}\quad g+h=r^{\pi_{2}}+\sigma^{\pi_{2}}(h).$ (53)
It then follows that,
$g+h$ $=\alpha\big(r^{\pi_{1}}+\sigma^{\pi_{1}}(h)\big)+(1-\alpha)\big(r^{\pi_{2}}+\sigma^{\pi_{2}}(h)\big)$ (54)
$=\sum_{a}\big(\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big)r(s,a)+\sum_{a}\big(\alpha\pi_{1}(a|s)+(1-\alpha)\pi_{2}(a|s)\big)\big(\sigma_{\mathcal{P}^{a}_{s}}(h)\big)$ (55)
$=r^{\tilde{\pi}}+\sigma^{\tilde{\pi}}(h).$ (56)
Consequently, $(g,h)$ is also a solution to the robust Bellman equation w.r.t. $\tilde{\pi}$. This implies that $g^{\tilde{\pi}}=g=g^{\star}$, and thus that $\tilde{\pi}$ is also optimal.
Hence the optimal policy set is convex, which completes the proof. ∎
C.2 Hemi-continuity
Lemma C.4.
Consider a policy sequence $\{\pi_{n}\}\rightarrow\pi$, and the sequence of induced transition matrices $\{\mathsf{P}_{\pi_{n}}\}$. It holds that $\mathsf{P}_{\pi_{n}}^{\star}\rightarrow\mathsf{P}_{\pi}^{\star}$ uniformly on $\mathcal{P}$.
Proof.
See Prop 8.4.6 in (puterman1994markov) for the proof. ∎
Lemma C.5.
Let $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ be a joint policy of all other agents except agent $i$. Define the operator $\mathcal{T}^{\pi}_{\mathcal{P}}:\mathbb{R}^{\mathcal{S}}\rightarrow\mathbb{R}^{\mathcal{S}\mathcal{A}_{i}}$ as $\mathcal{T}^{\pi}_{\mathcal{P}^{a}_{s}}(h)(s,a)=r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)$, where $(\mathcal{P}^{\pi})^{a}_{s}=\big\{q(\cdot)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)\mathsf{P}(\cdot|s,(a,b))\ :\ \mathsf{P}(\cdot|s,(a,b))\in\mathcal{P}^{(a,b)}_{s}\big\}$ and $r^{\pi}(s,a)=\sum_{b\in\mathcal{A}_{-i}}\pi(b)r(s,a,b).$ Consider the sequences $\{\pi_{n}\}\rightarrow\pi$ and $\{h_{n}\}\rightarrow h$, then $\mathcal{T}^{\pi_{n}}_{\mathcal{P}}(h_{n})\rightarrow\mathcal{T}^{\pi}_{\mathcal{P}}(h)$ uniformly on $\mathcal{S}\times\mathcal{A}_{i}$.
Proof.
Recalling the definition of $\mathcal{T}_{\mathcal{P}}$ we can obtain the difference,
[…]
$=\max_{s,a}\big|r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-r^{\pi}(s,a)\big|$
$\leq\max_{s,a}\big\{|r^{\pi_{n}}(s,a)-r^{\pi}(s,a)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|\big\}$
[…]
$B$ $=|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|$
$=|\min_{q\in(\mathcal{P}^{\pi_{n}})^{a}_{s}}qh-\min_{q\in(\mathcal{P}^{\pi})^{a}_{s}}qh|$
$=|\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi_{n}}h-\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi}h|$
$=\bigg|\min_{q\in\mathcal{P}^{a}_{s}}\Big(\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi_{n}(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big)-\min_{q\in\mathcal{P}^{a}_{s}}\Big(\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big)\bigg|$
$\leq\bigg|\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\Big(\pi_{n}(b|s)-\pi(b|s)\Big)q(s^{\prime}|s,a,b)h(s^{\prime})\bigg|,\quad\quad\text{for some }q\in\mathcal{P}^{a}_{s},$
$\leq||\pi_{n}(\cdot|s)-\pi(\cdot|s)||_{1}\cdot H,$
[…]
Note that $\pi_{n}(s)\rightarrow\pi(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},$ $\pi_{n}(s)-\pi(s)||_{1}\leq\epsilon$. Similarly with $h_{n}(s)\rightarrow h(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},\ |h_{n}(s)-h(s)|\leq\epsilon.$ Now by setting $N=\max_{s\in\mathcal{S}}\{N_{s}\}$, then $\forall n>N,$
[…]
Therefore, when $n>N,$ $\pi_{n}(s)-\pi(s)||_{1}+A+B\leq\epsilon+\epsilon+H\epsilon\lesssim\epsilon$. This implies that $\mathcal{T}^{n}_{\mathcal{P}}\rightarrow\mathcal{T}_{\mathcal{P}}$ uniformly on $\mathcal{S}\times\mathcal{A}$, which completes the proof. ∎
Theorem C.2 (Hemi-continuity).
Let $\pi_{n}\rightarrow\pi$ and $x_{n}\rightarrow x$, where $\pi_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ and $x_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$. If $x_{n}\in\text{BR}_{i}(\pi_{n}),$ then $x\in\text{BR}_{i}(\pi).$
Proof.
To show that $x\in\text{BR}_{i}(\pi),$ it suffices to show that $x$ is an optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi},r^{\pi}).$ Recalling the robust Bellman equation, it is also sufficient to show that $g^{(x,\pi)}_{\mathcal{P}}$ satisfies the following:
$g(s)=\max_{a\in\mathcal{A}_{i}}\{r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-h(s)\}=\max_{a\in\mathcal{A}_{i}}\{\mathcal{T}^{\pi}_{\mathcal{P}_{s}^{a}}(h)\}.$ (58)
See that since $x_{n}\in\text{BR}_{i}(\pi_{n})$, then $x_{n}$ is the optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{n}},r^{\pi_{n}}).$ Thus, there exists $h_{n}$, s.t. $(h_{n},g_{n}=g^{x_{n}}_{\mathcal{P}^{\pi_{n}}})$ is a solution to,
$g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g_{n}(s)=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}.$
See that $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=\min_{\mathsf{P}^{\pi_{n}}\in\mathcal{P}^{\pi_{n}}}g^{x_{n}}_{{}_{\mathsf{P}^{\pi_{n}}},r^{\pi_{n}}}=\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n},\pi_{n})}_{\mathsf{P},n}$, and so $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g^{\pi_{n},x_{n}}_{\mathcal{P}}$. Therefore, for any $(x_{n},\pi_{n})$, there exists some $h_{n}\in\mathbb{R}^{\mathcal{S}}$ such that for all $s\in\mathcal{S}$,
$g^{(x_{n},\pi_{n})}_{\mathcal{P}_{s}}(s)$ $=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}$
$=\max_{a}\{\mathcal{T}^{\pi_{n}}_{\mathcal{P}_{s}^{a}}(h_{n})\}.$ (59)
Moreover, since $h_{n}$ is some relative function (wang2023robust), $h_{n}||_{\infty}\leq H.$ Then, there exists some subsequence $\{n_{m}\},$ s.t. $h_{n_{m}}\rightarrow h$ for some $h$, and $g^{(x_{n},\pi_{n})}_{\mathcal{P}}\rightarrow g$ for some $g$. Then, we can take the limit as $m\rightarrow\infty$ in (C.2) to find:
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big\}$
$\overset{(i)}{=}\max_{a}\big\{\lim_{m\rightarrow\infty}\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big\}$
$=\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big\},$ (60)
where $(i)$ holds by the uniform convergence in Lemma C.5, and the final equality by $\pi_{n_{m}}\rightarrow\pi$, and $h_{n_{m}}\rightarrow h$. Looking at the left-hand side of equation (C.2) we find
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\cdot r^{(x_{n_{m}},\pi_{n_{m}})}.$
By Lemma C.4 we have $\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\rightarrow\mathsf{P}^{\star}_{(x,\pi)}$ uniformly. Thus, $\lim\min=\min\lim$ implies,
$\lim_{m\rightarrow\infty}g_{\mathcal{P}}^{(x_{n_{m}},\pi_{n_{m}})}$ $=\min_{\mathsf{P}\in\mathcal{P}}\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\min_{\mathsf{P}\in\mathcal{P}}g^{(x,\pi)}_{\mathsf{P}}$
$=g^{(x,\pi)}_{\mathcal{P}}$.
Therefore, $g^{(x,\pi)}_{\mathcal{P}}=\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big\}$ for some $h\in\mathbb{R}^{\mathcal{S}}$, which completes the proof. ∎
C.3 Technical Lemmas
In this section, we provide some supporting lemmas necessary for our main analysis.
Lemma C.6.
Define the ergodicity coefficient of a stochastic matrix $\mathsf{P}$ as
$\alpha(\mathsf{P})=1-\frac{1}{2}\max_{i,j}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|.$
Then for any stochastic matrix $\mathsf{P}$ and vector $V$, $\text{sp}(\mathsf{P}V)\leq(1-\alpha(\mathsf{P}))\cdot\text{sp}(V)$.
Proof.
For any two states $i,j$:
$(\mathsf{P}V)(i)-(\mathsf{P}V)(j)$ $=\sum_{k}(\mathsf{P}(k|i)-\mathsf{P}(k|j))V(k)$
$\leq\left(\frac{1}{2}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|\right)(\max_{s}V(s)-\min_{s}V(s))$
$\leq\left(\max_{i^{\prime},j^{\prime}}\frac{1}{2}\sum_{k}|\mathsf{P}(k|i^{\prime})-\mathsf{P}(k|j^{\prime})|\right)\text{sp}(V)$
$=(1-\alpha(\mathsf{P}))\text{sp}(V)$
Since this holds for any pair $(i,j)$, it must hold for the pair that defines the span of $\mathsf{P}V$. ∎
Lemma C.7 (Difference of Robust Operators).
For any $V_{1},V_{2}\in\mathbb{R}^{|\mathcal{S}|}$, there exists a valid transition kernel $\mathsf{P}^{\star}$ constructible from $\mathcal{P}$ such that $\sigma^{\pi}(V_{1})-\sigma^{\pi}(V_{2})=\mathsf{P}^{\star}(V_{1}-V_{2})$.
Proof.
We first show this for a single state-action operator $\sigma_{\mathcal{P}_{s}^{a}}(V)=\min_{\mathsf{P}\in\mathcal{P}_{s}^{a}}\mathsf{P}V$. Let $\mathsf{P}_{1}=\arg\min_{\mathsf{P}}\mathsf{P}V_{1}$ and $\mathsf{P}_{2}=\arg\min_{\mathsf{P}}\mathsf{P}V_{2}$. From the optimality conditions, we have $\mathsf{P}_{1}V_{1}\leq\mathsf{P}_{2}V_{1}$ and $\mathsf{P}_{2}V_{2}\leq\mathsf{P}_{1}V_{2}$. This yields the bounds:
$\mathsf{P}_{1}(V_{1}-V_{2})\leq\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})\leq\mathsf{P}_{2}(V_{1}-V_{2})$
Since $\mathcal{P}_{s}^{a}$ is a convex set and the function $h(\mathsf{P})=\mathsf{P}(V_{1}-V_{2})$ is continuous in $\mathsf{P}$, by the Intermediate Value Theorem there must exist a $\mathsf{P}_{s,a}^{\star}\in\mathcal{P}_{s}^{a}$ such that $\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})=\mathsf{P}_{s,a}^{\star}(V_{1}-V_{2})$. We then construct the global kernel $\mathsf{P}^{\star}$ row by row as $\mathsf{P}^{\star}(s^{\prime}|s)=\sum_{a}\pi(a|s)\mathsf{P}_{s,a}^{\star}(s^{\prime}|s,a)$. This leads to the desired vector equality. ∎
Appendix D Weakly Communicating DR-MGs
Robust average reward and best responses. For a stationary policy $\pi_{i}:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of agent $i$, define
$g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}\ \triangleq\ \min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\pi_{i}}_{P},\qquad\text{BR}_{i}(\pi_{-i})\ \triangleq\ \Bigl\{\pi_{i}:\ g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}=\sup_{\mu:\,\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\mu}_{P}\Bigr\}.$ (61)
Constant-gain robust Bellman equations. Under ˜2, there exists a constant-gain solution $(u^{\star},\alpha^{\star})$ to the statewise equations (Theorem 2 in (wang2025bellman)):
$u^{\star}(s)$ $=\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\ \sup_{\varphi\in\Delta(\mathcal{A}_{i})}\sum_{a_{i}}\varphi(a_{i})\Big(r_{i}^{\pi_{-i}}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Big),$ (62)
$u^{\star}(s)$ $=\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\ \inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sum_{a_{i}}\varphi(a_{i})\Big(r_{i}^{\pi_{-i}}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Big),\qquad s\in\mathcal{S},$ (63)
with attainment of both $\inf$ and $\sup$.
D.1 Convexity
For $s\in\mathcal{S}$ define
$L_{s}\!\left(\varphi,p_{s}\right)$ $\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\big),\quad\varphi\in\Delta(\mathcal{A}_{i}),\ p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s},$ (64)
$F_{s}(\varphi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}\!\left(\varphi,p_{s}\right),\qquad\Psi_{s}(p_{s})\triangleq\sup_{\varphi\in\Delta(\mathcal{A}_{i})}L_{s}\!\left(\varphi,p_{s}\right).$ (65)
Then $F_{s}$ is concave and upper hemi-continuous on the compact convex set $\Delta(\mathcal{A}_{i})$ (infimum of continuous affine maps). Using (62)–(63),
$u^{\star}(s)=\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi)=\min_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\Psi_{s}(p_{s}),\qquad s\in\mathcal{S}.$ (66)
Let
$\mathcal{M}_{s}\triangleq\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi).$ (67)
Then $\mathcal{M}_{s}$ is nonempty, convex, compact (boyd2004convex).
Lemma D.1 (Common supporting minimizer).
For each $s\in\mathcal{S}$ there exists $p_{s}^{\star}\in\arg\min_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\Psi_{s}(p_{s})$ such that
$u^{\star}(s)=\Psi_{s}(p_{s}^{\star})=L_{s}\!\left(\varphi,p_{s}^{\star}\right)=F_{s}(\varphi)\qquad\forall\,\varphi\in\mathcal{M}_{s}.$ (68)
Proof.
By Berge’s Maximum Theorem (berge1877topological), $\Psi_{s}$ is continuous on compact $(\mathcal{P}^{\pi_{-i}})_{s}$, so a minimizer $p_{s}^{\star}$ exists. Then (66) gives $\Psi_{s}(p_{s}^{\star})=u^{\star}(s)$. For any $\varphi\in\mathcal{M}_{s}$,
$u^{\star}(s)=F_{s}(\varphi)=\inf_{p_{s}}L_{s}\!\left(\varphi,p_{s}\right)\leq L_{s}\!\left(\varphi,p_{s}^{\star}\right)\leq\Psi_{s}(p_{s}^{\star})=u^{\star}(s).$ (69)
Equalities hold throughout. ∎
By $(s,a_{i})$-rectangularity, $p^{\star}\triangleq\{p_{s}^{\star}\}_{s\in\mathcal{S}}\in\mathcal{P}^{\pi_{-i}}$. And under ˜2, $p^{\star}$ admits a closed recurrent class $C_{p^{\star}}\subseteq\mathcal{S}$ and $\mathcal{S}\setminus C_{p^{\star}}$ is transient for all stationary policies.
Lemma D.2 (Slack inequalities).
For $\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ define $\delta_{\nu}(s)\triangleq u^{\star}(s)-F_{s}(\nu(\cdot\mid s))\geq 0$. Under $(\nu,p^{\star})$,
$\limsup_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\big(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\big)\right]$ $\leq 0,$ (70)
$\liminf_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\big(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\big)\right]$ $\geq-\,\limsup_{n\to\infty}\mathbb{E}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\right].$ (71)
If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for all $s\in C_{p^{\star}}$, then $g^{\nu}_{p^{\star}}=\alpha^{\star}$.
Proof.
From Lemma˜D.1, $u^{\star}(s)=\sup_{\varphi}L_{s}\!\left(\varphi,p_{s}^{\star}\right)$, hence
$u^{\star}(S_{t})\ \geq\ \mathbb{E}\!\big[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\big],$ (72)
and telescoping yields the first inequality. Also $F_{s}(\nu(\cdot\mid s))\leq L_{s}\!\left(\nu(\cdot\mid s),p_{s}^{\star}\right)$ implies
$u^{\star}(S_{t})-\delta_{\nu}(S_{t})\ \leq\ \mathbb{E}\!\big[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\big],$ (73)
and telescoping yields the second. If $\delta_{\nu}=0$ on $C_{p^{\star}}$ then its time-average vanishes because $\mathcal{S}\setminus C_{p^{\star}}$ is transient under $(\nu,p^{\star})$, so $g^{\nu}_{p^{\star}}=\alpha^{\star}$. ∎
Proposition D.1.
Let $\text{BR}_{i}(\pi_{-i})$ be the set of stationary best responses (optimal stationary policies of the induced DR-MDP). Then:
  • (a)
    (Necessary) If $\nu\in\text{BR}_{i}(\pi_{-i})$, then $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for every $s\in C_{p^{\star}}$.
  • (b)
    (Sufficient) If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for all $s\in\mathcal{S}$ (i.e., $\nu$ is statewise greedy for $(u^{\star},\alpha^{\star})$), then $\nu\in\text{BR}_{i}(\pi_{-i})$.
Proof.
(a) If $\nu(\cdot\mid s)\notin\mathcal{M}_{s}$ for some $s\in C_{p^{\star}}$, then $\delta_{\nu}(s)>0$. Since $s$ lies in the recurrent class of $p^{\star}$, the time-average of $\delta_{\nu}(S_{t})$ under $(\nu,p^{\star})$ is strictly positive; hence by Lemma D.2, $g^{\nu}_{p^{\star}}<\alpha^{\star}$, which contradicts optimality (because $\min_{p}g^{\nu}_{p}\leq g^{\nu}_{p^{\star}}<\alpha^{\star}$).
(b) If $\nu(\cdot\mid s)\in\mathcal{M}_{s}$ for every $s$, then by (66) we have $u^{\star}(s)=\inf_{p_{s}}L_{s}\!\left(\nu(\cdot\mid s),p_{s}\right)\leq L_{s}\!\left(\nu(\cdot\mid s),p_{s}\right)$ for all $p_{s}$, so summation and telescoping give $g^{\nu}_{p}\geq\alpha^{\star}$ for every $p\in\mathcal{P}^{\pi_{-i}}$. On the other hand, $\alpha^{\star}$ is the optimal robust value for the induced DR-MDP (Bellman optimality under weak communication), so $\min_{p}g^{\nu}_{p}\leq\alpha^{\star}$. Therefore $\min_{p}g^{\nu}_{p}=\alpha^{\star}$, i.e., $\nu\in\text{BR}_{i}(\pi_{-i})$. ∎
Theorem D.1 (Convexity).
Define
$\mathcal{G}_{i}\ \triangleq\ \bigl\{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})\ :\ \nu(\cdot\mid s)\in\mathcal{M}_{s}\ \ \forall s\in\mathcal{S}\bigr\}.$ (74)
Then $\mathcal{G}_{i}$ is nonempty, compact, and convex; and $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$.
Proof.
Each $\mathcal{M}_{s}$ is nonempty, compact, convex; the product set is nonempty, compact, convex. The inclusion follows from Proposition˜D.1. ∎
D.2 Hemi-continuity
For each $s\in\mathcal{S}$ and $(u,\alpha)\in\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$ define
$L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha)$ $\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\Big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha+p_{s,a_{i}}^{\top}u\Big),$ (75)
$F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha),$ (76)
$\mathcal{M}_{s}(\pi_{-i})$ $\triangleq\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u^{\star}_{\pi_{-i}},\alpha^{\star}_{\pi_{-i}}),$ (77)
$\mathcal{G}_{i}(\pi_{-i})$ $\triangleq\bigl\{\Delta:\mathcal{S}\to\Delta(\mathcal{A}_{i})\ :\ \Delta(\cdot\mid s)\in\mathcal{M}_{s}(\pi_{-i})\ \ \forall s\bigr\}.$ (78)
Here $(u^{\star}_{\pi_{-i}},\alpha^{\star}_{\pi_{-i}})$ denotes a constant-gain solution for the induced DR-MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i})$ (wang2025bellman).
Lemma D.3 (Hausdorff continuity of induced slices).
For fixed $(s,a_{i})$, define $M(w)\triangleq\sum_{a_{-i}}w(a_{-i})\,\mathcal{P}^{(a_{i},a_{-i})}_{s}$, $w\in\Delta(\mathcal{A}_{-i})$. If $w_{n}\to w$ in $\ell^{1}$, then $M(w_{n})\to M(w)$ in the Hausdorff metric; consequently $(\mathcal{P}^{\pi^{n}_{-i}})^{\,s}_{a_{i}}\to(\mathcal{P}^{\pi_{-i}})^{\,s}_{a_{i}}$ whenever $\pi^{n}_{-i}(\cdot\mid s)\to\pi_{-i}(\cdot\mid s)$.
Proof.
Our proof will utilize the Kuratowski convergence, which is equivalent to the convergence with respect to the Hausdorff metric on sets, if the space is a compact metric space (kuratowski2014topology).
By ˜2, each $\mathcal{P}^{(a_{i},a_{-i})}_{s}$ is nonempty and compact. Consider a sequence $x_{n}\in M(w_{n})$ with $x_{n}\to x$, which can be rewritten as $x_{n}=\sum_{a_{-i}}w_{n}(a_{-i})\,q^{n}_{a_{-i}}$ with $q^{n}_{a_{-i}}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}$. By compactness, for each $a_{-i}$, there exists a convergent subsequence $q^{n}_{a_{-i}}\to q_{a_{-i}}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}$. Through the diagonal argument, it holds that $x=\sum_{a_{-i}}w(a_{-i})\,q_{a_{-i}}\in M(w)$. On the other hand, fix $x=\sum_{a_{-i}}w(a_{-i})\,q_{a_{-i}}\in M(w)$ and set $x_{n}\triangleq\sum_{a_{-i}}w_{n}(a_{-i})\,q_{a_{-i}}\in M(w_{n})$; then $x_{n}\to x$. Hence the Kuratowski outer and inner limits coincide and, since the sets are compact, Hausdorff convergence follows (beer1993topologies). ∎
Lemma D.4 (Uniform continuity of min-envelope).
Let $(u_{n},\alpha_{n},\pi^{n}_{-i})\to(u,\alpha,\pi_{-i})$, with $\{u_{n}\}$ uniformly bounded. Then for each $s$,
$\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi_{-i}}_{s}(\varphi;u,\alpha)\big|\ \to\ 0.$ (79)
Proof.
For fixed $s$, the integrand $L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha)$ is continuous and affine in $(\varphi,p_{s},u,\alpha)$; $r^{\pi_{-i}}_{i}(s,a_{i})$ depends linearly on $\pi_{-i}(\cdot\mid s)$; $p_{s,a_{i}}^{\top}u$ is continuous. By Lemma D.3, the feasible sets $(\mathcal{P}^{\pi^{n}_{-i}})_{s}$ converge to $(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric. Berge’s Maximum Theorem (applied to the minimization of a continuous function over compact feasible sets that vary Hausdorff‑continuously) gives continuity of the value map $(\varphi,u,\alpha,\pi_{-i})\mapsto F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$. Compactness of $\Delta(\mathcal{A}_{i})$ yields uniform convergence in $\varphi$. ∎
Lemma D.5 (Argmax stability under uniform convergence (ross2013elementary)).
Let $K$ be compact and $f_{n},f:K\to\mathbb{R}$ with $f_{n}\to f$ uniformly. If $x_{n}\in\arg\max_{x\in K}f_{n}(x)$ and $x_{n}\to x$, then $x\in\arg\max_{x\in K}f(x)$.
Lemma D.6 (Stability of Bellman pairs).
Let $\{\pi^{n}_{-i}\}_{n\geq 1}\subset\prod_{s}\Delta(\mathcal{A}_{-i})$ satisfy $\pi^{n}_{-i}\to\pi_{-i}$ (product topology). For each $n$, let $(u_{n},\alpha_{n})$ solve the attained sup–inf constant-gain equation
$u_{n}(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\,F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n}),\qquad s\in\mathcal{S},$ (80)
and normalize $u_{n}(s_{0})=0$ at a fixed $s_{0}\in\mathcal{S}$. Then:
  • (i)
    $0\leq\alpha_{n}\leq 1$ for all $n$, and $\|u_{n}\|_{\infty}\leq 1$ for all $n$.
  • (ii)
    There exists a subsequence (not relabeled) such that $(u_{n},\alpha_{n})\to(u,\alpha)$ in $\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$.
  • (iii)
    The limit $(u,\alpha)$ solves the sup–inf constant-gain equation at $\pi_{-i}$:
    $u(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\,F^{\pi_{-i}}_{s}(\varphi;u,\alpha),\qquad s\in\mathcal{S}.$ (81)
Proof.
Since $r\in[0,1]$, every long-run average reward lies in $[0,1]$, hence $0\leq\alpha_{n}\leq 1$.
Let $m_{n}\triangleq\min_{s}u_{n}(s)$ and $M_{n}\triangleq\max_{s}u_{n}(s)$. For any $s$, any $\varphi$, and any feasible $p_{s}$, it holds that
$-1+m_{n}\;\leq\;-1+\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(p_{s,a_{i}}^{\top}u_{n}\big)\;\leq\;L_{s}\!\left(\varphi,p_{s};u_{n},\alpha_{n}\right)\;\leq\;1+\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(p_{s,a_{i}}^{\top}u_{n}\big)\;\leq\;1+M_{n}.$ (82)
Taking $\inf_{p_{s}}$ and then $\sup_{\varphi}$ yields, for every $s$,
[…]
Therefore $M_{n}-m_{n}\leq 1$. With the normalization $u_{n}(s_{0})=0$ we have $m_{n}\leq 0$ and hence $M_{n}\leq 1$, so $|u_{n}(s)|\leq 1$ for all $s$, i.e., $\|u_{n}\|_{\infty}\leq 1$.
As we have shown that $\{u_{n}\}$ is bounded in $\mathbb{R}^{\mathcal{S}}$ and $\{\alpha_{n}\}\subset[0,1]$, there exists a convergent subsequence (ross2013elementary) (not relabeled) such that $(u_{n},\alpha_{n})\to(u,\alpha)$ in $\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$.
We then consider a fixed $s\in\mathcal{S}$. For each $a_{i}$, the induced slice
$(\mathcal{P}^{\pi^{n}_{-i}})^{s}_{a_{i}}=\sum_{a_{-i}}\pi^{n}_{-i}(a_{-i}\!\mid s)\,\mathcal{P}^{(a_{i},a_{-i})}_{s}$ (84)
converges to $(\mathcal{P}^{\pi_{-i}})^{s}_{a_{i}}$ in the Hausdorff metric because the right-hand side is a Minkowski sum with convex weights that depend continuously on $\pi_{-i}(\cdot\mid s)$; hence $(\mathcal{P}^{\pi^{n}_{-i}})_{s}\to(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric (Lemma˜D.3).
Moreover, for $(u,\alpha)$ in a bounded set and $\varphi\in\Delta(\mathcal{A}_{i})$, the map
$(\varphi,p_{s},u,\alpha,\pi_{-i})\ \mapsto\ L^{\pi_{-i}}_{s}(\varphi,p_{s};u,\alpha)\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha+p_{s,a_{i}}^{\top}u\big)$ (85)
is continuous and affine in each argument; by Berge’s Maximum Theorem, the value function $(\varphi,u,\alpha,\pi_{-i})\mapsto F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$ is continuous when the feasible sets vary Hausdorff-continuously. Consequently,
$F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})\ \longrightarrow\ F^{\pi_{-i}}_{s}(\varphi;u,\alpha),\qquad\forall\,\varphi\in\Delta(\mathcal{A}_{i}).$ (86)
For any $n$, $s$, and $\varphi,\psi\in\Delta(\mathcal{A}_{i})$,
$\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi^{n}_{-i}}_{s}(\psi;u_{n},\alpha_{n})\big|\leq\sup_{p_{s}\in(\mathcal{P}^{\pi^{n}_{-i}})_{s}}\Big|\sum_{a_{i}\in\mathcal{A}_{i}}(\varphi-\psi)(a_{i})\big(r^{\pi^{n}_{-i}}_{i}(s,a_{i})-\alpha_{n}+p_{s,a_{i}}^{\top}u_{n}\big)\Big|\leq 3\,\|\varphi-\psi\|_{1},$ (87)
because $r^{\pi^{n}_{-i}}_{i}\in[0,1]$, $\alpha_{n}\in[0,1]$, and $|p_{s,a_{i}}^{\top}u_{n}|\leq\|u_{n}\|_{\infty}\leq 1$. Thus the family $\{F^{\pi^{n}_{-i}}_{s}(\cdot;u_{n},\alpha_{n})\}_{n}$ is equi-Lipschitz on the compact $\Delta(\mathcal{A}_{i})$. By (86) and a standard $\varepsilon$–net argument (or Arzelà–Ascoli’s criterion for equicontinuous families on compact domains), we obtain the uniform convergence
$\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\;\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi_{-i}}_{s}(\varphi;u,\alpha)\big|\ \longrightarrow\ 0.$ (88)
Finally, from (80) and (88),
$u_{n}(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})\ \longrightarrow\ \sup_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u,\alpha).$ (89)
Since $u_{n}(s)\to u(s)$, we conclude (81) for every $s\in\mathcal{S}$. This hence completes the proof. ∎
Theorem D.2 (Upper hemi-continuity of $\mathcal{G}_{i}(\cdot)$).
Let $\pi^{n}_{-i}\to\pi_{-i}$ in the product topology on stationary policies, and let $\Delta_{n}\in\mathcal{G}_{i}(\pi^{n}_{-i})$ for each $n$. If $\Delta_{n}\to\Delta$ pointwise (equivalently, in the product topology), then $\Delta\in\mathcal{G}_{i}(\pi_{-i})$.
Proof.
For each $n$, choose an attained constant‑gain solution $(u_{n},\alpha_{n})$ for $\mathsf{M}_{i}(\pi^{n}_{-i})$, normalized by $u_{n}(s_{0})=0$. By Lemma D.6, along a subsequence (not relabeled) $(u_{n},\alpha_{n})\to(u,\alpha)$ and $(u,\alpha)$ solves the limit equation for $\mathsf{M}_{i}(\pi_{-i})$.
Fix $s\in\mathcal{S}$. Because $\Delta_{n}\in\mathcal{G}_{i}(\pi^{n}_{-i})$, we have
$\Delta_{n}(\cdot\mid s)\in\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n}).$ (90)
By Lemma D.4, $F^{\pi^{n}_{-i}}_{s}(\cdot;u_{n},\alpha_{n})\to F^{\pi_{-i}}_{s}(\cdot;u,\alpha)$ uniformly on $\Delta(\mathcal{A}_{i})$. Since $\Delta_{n}(\cdot\mid s)\to\Delta(\cdot\mid s)$ and $\Delta(\mathcal{A}_{i})$ is compact, Lemma D.5 yields
$\Delta(\cdot\mid s)\in\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u,\alpha)=:\mathcal{M}_{s}(\pi_{-i}).$ (91)
As $s$ was arbitrary, $\Delta\in\mathcal{G}_{i}(\pi_{-i})$, proving the upper hemi-continuity. ∎
D.3 Existence of DR-NE
We can then prove the existence of DR-NE following the results shown above. Define a map $G(\pi)\triangleq\bigotimes_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$ on the compact convex policy space $X=\prod_{i,s}\Delta(\mathcal{A}_{i})\to 2^{X}$. Since each $\mathcal{G}_{i}(\cdot)$ is nonempty, compact, convex (Theorem˜D.1), and is upper hemi-continuous (Theorem˜D.2), thus $G$ satisfies Kakutani’s condition as in Lemma˜4.2. Hence there exists a fixed point $\pi^{\star}\in G(\pi^{\star})\subseteq\bigotimes_{i}\text{BR}_{i}(\pi^{\star}_{-i})$, which is a stationary DR-NE.
Appendix E Proof of Section 6
Assumption 3.
In Algorithm 1, at each step, the Nash Equilibrium $\pi(s)=\textbf{NE}(Q_{i}(s,a,h^{i}))$ satisfies one of the following conditions:
  • $\pi$ is global optimal: $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\geq\mathbb{E}_{\pi^{\prime}}[Q_{i}(s,a,h^{i})]$ for all joint policy $\pi^{\prime}$,$i,s,a$.
  • $\pi$ is a saddle point such that $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\leq\mathbb{E}_{(\pi^{\prime}_{-i},\pi_{i})}[Q_{i}(s,a,h^{i})]$, for all $i,s,a,$ and joint policy $\pi^{\prime}_{-i}$.
Such an assumption is necessary for convergence (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence).
Theorem E.1.
Additionally assume Assumption 3. The robust Nash-Iteration algorithm converge to an average reward Nash Equilibrium of the distributionally robust game.
Proof.
We first show that, if the algorithm converges to $\pi$, then $\pi$ is an average reward Nash Equilibrium. To show this, we note that the convergence of Alg 1 implies that there exists some constant $c_{1},...,c_{N}$, such that
$\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i})]-h_{i}(s)=c_{i},\forall i,s.$ (92)
This implies that
$h_{i}(s)+c_{i}=\sum_{a}\pi(a|s)(r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})),$ (93)
which is the robust Bellman equation w.r.t. $\pi$. Hence by Theorem 4.1, $(c_{i},h_{i})=(g^{\pi}_{\mathcal{P},i},h^{\pi}_{\mathcal{P},i})$ for some vector $h^{\pi}_{\mathcal{P},i}$, i.e.,
$h^{\pi}_{\mathcal{P},i}+g^{\pi}_{\mathcal{P},i}=r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i}).$ (94)
Now assume that $\pi$ is not an average reward Nash Equilibrium, then there exists some agent $i$ and some policy of it $\mu_{i}$, such that
$g^{(\pi_{-i},\mu)}_{\mathcal{P},i}>g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}$. (95)
Similarly, there exists some $h^{\mu}_{i}$ satisfying
$h^{\mu}_{i}+g^{(\pi_{-i},\mu)}_{\mathcal{P},i}=r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i}).$ (96)
Combining (94),(95),(96) implies that
$r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}=g^{(\pi_{-i},\mu)}_{\mathcal{P},i}>g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}=r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-h^{\pi}_{\mathcal{P},i}.$ (97)
Moreover, since $\pi$ is the Nash Equilibrium of $Q_{i}$, i.e.,
$r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i})\leq r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i}),$ (98)
hence
$r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-h^{\pi}_{\mathcal{P},i}<r^{(\pi_{-i},\mu)}_{i}+\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}\leq\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-h^{\mu}_{i}+r^{\pi}_{i}+\sigma^{\pi}(h^{\pi}_{\mathcal{P},i})-\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i}),$ (99)
i.e.,
$h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i}<\sigma^{(\pi_{-i},\mu)}(h^{\mu}_{i})-\sigma^{(\pi_{-i},\mu)}(h^{\pi}_{\mathcal{P},i})\leq P(h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i})$ (100)
for some $P\in\mathcal{P}$. Due to the irreducibility, it holds that
$h^{\mu}_{i}-h^{\pi}_{\mathcal{P},i}=c_{i}e$ (101)
for some constant $c_{i}$. Thus $g^{(\pi_{-i},\mu)}_{\mathcal{P},i}=g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}$, which is a contradiction to (95), and hence $\pi$ is an average reward Nash Equilibrium.
It then remains to show the convergence of Algorithm 1, and it is sufficient to show that for any $i$, the operator $\mathbf{T}=(\mathbf{T}_{1},...,\mathbf{T}_{N}):\mathbb{R}^{NS}\to\mathbb{R}^{NS}$ is a contraction under Span heminorm, where
$\mathbf{T}_{i}:\mathbb{R}^{NS}\to\mathbb{R}^{S}:\mathbf{T}_{i}(h=(h_{1},...,h_{N}))(s)=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})],$ (102)
here $\pi$ is the Nash Equilibrium specified by Algorithm 1. Consider $h,h^{\prime}$, and we consider the two cases from Assumption 3. If $\pi$ is global optimal, then
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)$ $=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$=\max_{\pi(s)}\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\max_{\pi^{\prime}(s)}\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\mathbb{E}_{\pi(s)}[\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\max_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i}));$ (103)
And if the second case from Assumption 3 holds, then
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)$ $=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\overset{(a)}{\leq}\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\overset{(b)}{\leq}\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i})]-\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})]$
$\leq\max_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})),$ (104)
where $(a)$ is due to $\pi^{\prime}$ is the Nash Equilibrium, and $(b)$ is due to the saddle point assumption. Similarly, we can show that
$\mathbf{T}_{i}(h)(s)-\mathbf{T}_{i}(h^{\prime})(s)\geq\min_{a}(\sigma_{\mathcal{P}^{a}_{s}}(h_{i})-\sigma_{\mathcal{P}^{a}_{s}}(h^{\prime}_{i})).$ (105)
Thus Lemma C.7 implies that
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq\text{sp}(\sigma_{\mathcal{P}}(h_{i})-\sigma_{\mathcal{P}}(h^{\prime}_{i}))=\text{sp}(\mathsf{P}(h_{i}-h^{\prime}_{i})),$ (106)
for some irreducible transition matrix $\mathsf{P}$. Applying Lemma C.6 implies
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq(1-\alpha(\mathsf{P}))\text{sp}(h\mathsf{P}_{0}i-h_{i}^{\prime}).$ (107)
Now since for irreducible transition matrix $\mathsf{P}$, $\alpha(\mathsf{P})<1$ (puterman1994markov), then
$\text{sp}(\mathbf{T}_{i}(h)-\mathbf{T}_{i}(h^{\prime}))\leq\gamma_{i}\text{sp}(h_{i}-h_{i}^{\prime})$ (108)
for some constant $\gamma_{i}<1$. Taking the maximum over $i\in\mathcal{N}$ hence implies the Span-contraction, which completes the proof. ∎
Appendix F Proof of Section 7
Theorem F.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
Proof.
(1). By the uniform convergence of the robust discounted value function derived in (wang2023robust), we have that
$\lim_{\gamma\to 1}(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}=g^{\pi}_{\mathcal{P},i},\text{ uniformly on }\Pi.$ (109)
Thus for any $\epsilon$, there exists some $\gamma_{e}$, such that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,$ (110)
for any $\pi\in\Pi$ and $\gamma>\gamma_{e}$. Moreover, there exists some constant $T$ such that for any $t>T$,
$\gamma_{t}>\gamma_{e},\|\pi_{t}-\pi\|\leq\epsilon.$ (111)
Since $\pi_{t}$ is the Nash Equilibrium of the discounted robust MG, it holds that
$V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}\geq V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i},\forall i\in\mathcal{N},\forall\mu^{i}.$ (112)
Thus we have that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$
$=-g^{\pi}_{\mathcal{P},i}+(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}$
$\quad-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}$. (113)
Consider any $t>T$, it first holds that due to (110):
$(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$ $\leq\epsilon,$ (114)
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}$ $\leq\epsilon.$ (115)
Moreover,
$(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}\leq(1-\gamma_{t})(V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-V^{\pi}_{\gamma_{t},\mathcal{P},i})\leq\epsilon,$ (116)
$(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq\epsilon,$ (117)
which is due to the $\frac{1}{1-\gamma}$-Lipschitz of the robust discounted value function in $\pi$, and the closeness of $\pi_{t}$ and $\pi$ in (111). Finally,
$-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq 0$ (118)
due to (112). Combining all results hence implies that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq\epsilon,$ (119)
for any $i$ and $\mu^{i}$ uniformly. Thus let $\epsilon\to 0$ we complete the proof.
(2). Set $\gamma$ to be any factor that is larger than $\gamma_{e}$. Then
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}=g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}.$ (120)
Note that for $\gamma>\gamma_{e}$, it holds that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq 2\epsilon+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}\leq 3\epsilon,$ (121)
due to $\pi$ is an $\epsilon$-NE for the discounted robust MG, which hence completes the proof. ∎
Lemma F.1.
The discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
Proof.
The proof is a direct corollary of Theorem 7.1 and results from (Roch2025reduction). Specifically, it is shown in Theorem 3.4 in (Roch2025reduction) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^2] The results in (Roch2025reduction) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ (Roch2025reduction)., it holds that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,\forall\pi,\forall i.$ (122)
Combining with Theorem 7.1 further completes the proof. ∎