### v3

# Distributionally Robust Markov Games with Average Reward

Abstract
We study distributionally robust Markov games (DR‑MGs) with the average‑reward criterion, a framework for multi-agent decision-making under uncertainty over extended horizons. In average reward DR-MGs, agents aim to maximize their worst-case infinite-horizon average reward, to ensure satisfactory performance under environment uncertainties and opponent actions. We first establish a connection between the best-response policies and the optimal policies for the induced single‑agent problems. Under a standard irreducible assumption, we derive a correspondence between the optimal policies and the solutions of the robust Bellman equation, and derive the existence of stationary Nash Equilibrium (NE) based on these results. We further study DR-MGs under the weakly communicating setting, where We construct a set-valued map and show its value is a subset of the best-response policies, convex and upper hemi-continuous, and derive the existence of NE. We then explore algorithmic solutions, by first proposing a Robust Nash‑Iteration algorithm and providing convergence guarantees under some additional assumptions and a NE computing oracle. We further develop a temporal-difference based algorithm for DR-MGs, and provide convergence guarantees without any additional oracle or assumptions. Finally, we connect average‑reward robust NE to discounted ones, showing that the average reward robust NE can be approximated by the discounted ones under a large discount factor. Our studies provide a comprehensive theoretical and algorithmic foundation for decision-making in complex, uncertain, and long-running multi-player environments.
1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments (shapley1953stochastic; filar2012competitive; littman1994markov). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the Sim-to-Real gap (mcmahan2024roping), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks (bukharin2023robust; ma2023decentralized). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
To address this vulnerability, the framework of distributionally robust Markov Games (DR-MGs) has been developed (zhang2020robust; kardecs2011discounted), extending the principles of distributionally robust Markov Decision Processes (MDPs) (bagnell2001solving; nilim2004robustness; iyengar2005robust) to the multi-agent setting. Instead of relying on a single, fixed MG model, the robust approach seeks to find an equilibrium that optimizes the worst-case performance over a predefined uncertainty set of plausible game models under potential model mismatches. The resulting robust equilibrium provides a performance guarantee across all models within this set, thereby ensuring each agent’s resilience against model mismatch.
The nature of a DR-MG problem is fundamentally shaped by the optimality criterion used. Much of the prior work has concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) or discounted-reward setting (zhang2020robust; kardecs2011discounted), where the objective is to maximize the expected reward over a finite number of steps or a discounted sum of rewards.
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics (krnjaic2024scalable), communication networks (qu2020scalable), autonomous vehicle coordination (peake2020multi), financial markets (vadori2024towards), or even peer-to-peer energy trading (qiu2021multi). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate (hernandez2000zero). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium (filar2012competitive); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions (puterman2014markov; atia2021steady). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
Moreover, none of the methods for DR-MGs under the discounted reward or finite horizon can be extended to the average-reward setting. The existence of (non-stationary) Nash Equilibrium under finite horizon is derived through backward induction (blanchet2023double), which does not apply to infinite-horizon settings. The existence of stationary Nash Equilibrium under the discounted reward criterion is derived in (kardecs2011discounted), which however heavily relies on the uniqueness of the solution to the discounted robust Bellman equation (iyengar2005robust) and hence cannot be applied to the average reward setting. Another natural attempt is to use the discounted reward as an intermediate step to study the robust Nash Equilibrium under the average reward by setting the discount factor approaching to $1$. Although this limit method is effective in the single-agent setting (wang2023robust), it cannot solve the DR-MGs, due to the complicated structure (also see our detailed discussion in Section 3). To date, the literature on average-reward DR-MGs remains sparse, leaving fundamental questions about the structure of robust equilibria and the design of convergent solution algorithms largely unanswered.
1.1 Contributions
In this paper, we develop a comprehensive study of DR-MGs under the average-reward formulation. Our contributions are summarized as follows.
Existence of Robust Nash Equilibrium: We first study the solvability of average reward DR-MGs, i.e., the existence of robust NE. We first note that, finding the best response policy of an agent is equivelent to find the optimal policy of the single-agent induced robust MDP. We thus revisit the robust Bellman equation under average reward, and show that its solvability and the corresponding between its solutions and the optimal policies, under the irreducible assumption. Based on these results, we further prove the convexity and hemi-continuity of the best-response mapping, deriving the existence of a stationary Nash Equilibrium. We further study the more general weak communicating DR-MGs, where no such correspondence can be obtained due to their more complex structures. We instead study a proxy mapping which is a subset of the best-response mapping, and show the existence of NE based on it. These results address the fundamental open question, demonstrating that solutions exist in these complex, robust, and long-term horizon competitive environments.
Convergence Algorithm: We then look for concrete algorithms to find robust NE in average reward DR-MGs. Insipired by the standard value iteration approach, we first propose a robust Nash iteration algorithm. Under some additional assumptions on the game structure and an NE computing oracle, we prove the convergence to a robust NE of our algorithm. Moreover, to bypass these assumptions, we further develop a temporal-difference based method, which applies proximal descent on the temporal difference. Our algorithm does not require any NE oracle, and can be executed efficiently. We further develop its provably convergence guarantee without any additional assumption. Our algorithm designs and results hence provide the first comprehensive and algorithmic understanding on solving average-reward DR-MGs.
Connection to Discounted Equilibria: We further study the connection between the DR-MGs under average and discounted reward. We show that, the robust NE under the average reward can be effectively approximated by the robust NE under the discounted reward with some discount factor large enough. Such a connection enables us to solve the complicated and challenging average reward DR-MGs through the discounted ones, utilizing their element mathematical properties and extensively developed algorithms.
1.2 Related Works
We discuss the most related works in this section.
Non-robust Markov Games. Markov Games (littman1994markov) provide the foundational mathematical framework for multi-agent sequential decision-making. The majority of early work focused on the discounted-reward setting, where the existence of a stationary Nash Equilibrium was established in (shapley1953stochastic). The average-reward criterion, while more suitable for systems with long operational horizons, presents greater analytical challenges. Unlike the discounted case, a stationary NE is not guaranteed to exist in a general average-reward MG. Its existence has only been proven under additional structural assumptions on the dynamics of the game (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; filar2012competitive).
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, distributionally robust MDPs are first developed in (iyengar2005robust; nilim2004robustness) for both finite and discounted settings. The studies under the average reward setting are recently developed in (wang2023robust; wang2023model; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample; grand2023beyond; grand2023reducing; chatterjee2023solving; wang2024robust). However, these works are all developed for single-agent settings, and do not address the challenges we faced in multi-agent DR-MGs.
Distributionally Robust Markov Games. Extending distributional robustness to the multi-agent setting is a recent and active area of research. However, the literature has overwhelmingly concentrated on the finite-horizon (ma2023decentralized; shi2024breaking; shi2024sample; jiao_minimax-optimal_2024; li2025sample; blanchet2023double) and discounted-reward criteria (zhang2020robust; kardecs2011discounted). These settings are often more analytically tractable because the influence of future rewards diminishes, simplifying the analysis. However, none of these methods can be extended to the average-reward setting, as we shall discuss later.
2 Preliminaries and Problem Formulation
[…]
A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ is a distribution over $\mathcal{A}$ for any given state $s$, meaning the agent takes action $a$ at state $s$ with probability $\pi(a|s)$. The goal of an MDP is to maximize the accumulative reward. Since different reward criteria give rise to markedly distinct problem settings, the resulting objective functions can vary greatly, modifying the key properties necessary to analyze the underlying system. The more extensively studied criterion is the discounted reward, whose objective function is the discounted value function: $V^{\pi}_{\mathsf{P}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right]$, where $\gamma\in[0,1)$ is some discount factor. When long-term performance is of interest, the average reward is generally adopted as the criterion of choice. By effectively setting $\gamma=1$, it does not discount the reward over time, but considers the behavior of the underlying Markov process under the steady-state distribution. Under this formulation, the average reward induced by following policy $\pi$ starting from state $s\in\mathcal{S}$ is defined as $g_{\mathsf{P}}^{\pi}(s)\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}}\big[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\big].$
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (littman1994markov), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\times_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}_{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the joint actions of all agents. At some state $s$, all agents act simultaneously by taking a joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint (product) policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on stationary Nash Equilibrium (NE). More specifically, a stationary NE is a product policy $\pi^{\star}=(\pi^{\star}_{1},...,\pi^{\star}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{\star}$:
$V^{\pi^{\star}}_{\mathsf{P},i}(\rho)\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathsf{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}\in\Pi_{i},$ (1)
for some initial distribution $\rho$ and $V^{\pi}_{\mathsf{P},i}(\rho)\triangleq\mathbb{E}_{s\sim\rho}[V^{\pi}_{\mathsf{P},i}(s)]$, and $\Pi_{i}$ denotes the set of stationary policies for agent $i$, $\pi^{\star}_{-i}$ denotes the joint policy of all other agents except $i$ from $\pi^{\star}$, and $(\pi^{\star}_{-i},\pi_{i})$ is the joint policy. The existence of stationary NE is derived for MGs with discounted reward (shapley1953stochastic) and average-reward MGs with additional structural assumptions (filar2012competitive).
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\times_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ (iyengar2005robust; nilim2004robustness; zhang2020robust; blanchet2023double; shi2024sample), where the uncertainty set is independently defined over all $(s,a)$ pairs.
After all agents take their joint action $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}(\rho)$, and robust average reward $g^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}(\rho)$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{\star}$ such that for some initial state distribution $\rho$,
$g^{\pi^{\star}}_{\mathcal{P},i}(\rho)\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i}.$ (2)
3 Hardness in Average Reward Nash Equilibrium Identification
Extending existing studies for discounted or finite-horizon robust MGs to the average-reward setting is prevented by two major challenges:
Additional hidden environment player. The first major difference between a vanilla non-robust MG and a DR-MG is the introduction of an additional player, which selects the transition kernel or the underlying game model to diminish the gains of the other players (zhang2020robust). This additional player fundamentally changes the structure of non-robust MGs. For example, in the extensively studied special case of two-player zero-sum MGs, where $r_{1}=-r_{2}=r$, one fundamental structure of standard MGs is the Min-Max duality:
$\max_{\pi_{1}}\min_{\pi_{2}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)=\min_{\pi_{2}}\max_{\pi_{1}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho),$ (3)
which directly implies the existence of a Nash Equilibrium (NE) (shapley1953stochastic). Such results are further extended to non-robust average-reward MG studies through the limit method, which sets the discount factor $\gamma\to 1$ (filar2012competitive; sennott1994zero; rieder1995average).
However, when looking at the distributionally robust setting for MGs, this notion fails due to consideration of the worst-case scenario. Since agents have different rewards, their corresponding worst-case transition kernels are also different. Hence, the duality fails (note that from agent $2$’s perspective, the worst case is $\max_{\mathsf{P}}V^{\pi}_{\mathsf{P},r}$ as it aims to minimize w.r.t. $r$):
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho).$
Consequently, even through the uniform convergence of the discounted robust value functions enables the interchanging of the order of $\max,\min,\lim$ (wang2023robust), it does not imply the existence of NE. And this approach cannot be applied in multi-player general-sum games.
Complicated structure of the average reward. The second major challenge stems from the inherent complexity of the (robust) average reward, which heavily depends on structures of the underlying MDPs. Firstly, even in the non-robust setting, Markov games with average reward do not necessarily admit a stationary NE (filar2012competitive; wikecek2015stationary; lozovanu2016stationary; thuijsman1997survey). And with model uncertainty, the robust average reward has a more complicated structure, including a non-linear robust Bellman equation (wang2023robust; wang2024modelfree) with non-unique solutions. The studies and understandings of robust average reward is also significantly limited compared to non-robust and discounted settings (wang2023robust; wang2023model; wang2025bellman). Therefore, despite the existence of the NE being proved for both discounted robust MGs (kardecs2011discounted; zhang2020robust) and average-reward non-robust MGs (guo2008new; hernandez2000zero; guo2003zero; zheng2024zero; nowak1999optimal; iwase1976markov; kuenle2003optimality; tanaka1977continuous; jaskiewicz2001optimality; wikecek2015stationary), these results and methods cannot be applied for the NE of average-reward DR-MGs.
To better illustrate the hardness of solving average reward DR-MGs, we first show the following impossibility result. The constructed DR-MG is essentially a multi-chain (puterman2014markov), indicating the necessity of the additional structural assumptions to solve average reward DR-MGs.
Lemma 3.1.
There exists a DR-MG without any stationary robust NE under average reward.
4 NE of Irreducible DR-MGs
As discussed above, the existence of a NE cannot be guaranteed without any structural assumptions. In this section, we first study a relatively easier setting, where the underlying DR-MG is assumed to be irreducible.
Assumption 1.
For any deterministic joint policy $\pi$ and any transition kernel $\mathsf{P}\in\mathcal{P}$, the induced Markov chain $(\mathcal{S},\mathcal{A},\mathsf{P}^{\pi})$ is irreducible, i.e., all states belong to a single recurrent class. Moreover, $\mathcal{P}$ is compact and convex.
Remark 1.
Irreducibility is a standard assumption (slightly stronger than the unichain assumption (wang2023robust; wang2023model; wang2024robust; roch2025finite; Roch2025reduction)) in robust average reward studies (xu2025finite; xu2025efficient). Under this assumption, the robust average reward $g^{\pi}_{\mathcal{P}}(\rho)=g^{\pi}_{\mathcal{P}}(s)$ is always a constant over all initial states, ensuring the tractability of robust average reward. Hence in this section, we omit the initial state and denote the robust average reward by $g^{\pi}_{\mathcal{P}}$.
As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs. In non-robust average reward MG studies, the NE can be proved by studying the induced stationary/occupancy distribution (wikecek2015stationary). However, when considering robustness, the worst-case stationary distribution set becomes much more complicated. For example, it is non-convex (wang2022robust; zhang2024distributionally; ma2025rectified), whereas the convexity is necessary in non-robust proofs and hence the method cannot be adapted. We thus propose a direct approach by studying the best response mapping.
Definition 1.
(Best Response mapping) In an average-reward DR-MG, fix agent $i$ and the joint policy $\pi_{-i}$ of all other agents. The best response mapping of agent $i$ w.r.t. $\pi_{-i}$ is a set-valued mapping defined as
$\text{BR}_{i}(\pi_{-i})\triangleq\left\{\pi_{i}:g^{(\pi_{-i},\pi_{i})}_{\mathcal{P},i}=\max_{\mu_{i}}g^{(\pi_{-i},\mu_{i})}_{\mathcal{P},i}\right\}$. (4)
That is, the best response policy is agent $i$’s optimal policy when other agents fix their policies. This notion and the following results are standard and extensively studied in game theory literature (osborne2004introduction; fudenberg1991game):
Lemma 4.1.
$\pi^{\star}$ is a Nash Equilibrium if and only if $\pi^{\star}_{i}\in\text{BR}_{i}(\pi^{\star}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{\star}$ is a fixed point of the BR mapping: $\pi^{\star}\in\text{BR}(\pi^{\star})$, where $\text{BR}(\pi)=\times_{i}\text{BR}_{i}(\pi^{\star}_{-i})$.
Lemma 4.2 (Kakutani’s Fixed Point Theorem (kakutani1941generalization)).
Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) $f(x)$ is nonempty and convex, $\forall x$; (2) $f$ is upper hemi-continuous (i.e., for any convergent sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$, it holds that $y\in f(x)$). Then there exists some $x^{\star}\in X$ such that $x^{\star}\in f(x^{\star})$.
These two results imply that it is sufficient to show that the best response policy mapping, BR, satisfies the conditions in Lemma 4.2. We further show that these conditions can be further reduced to single-agent cases.
Lemma 4.3.
If for any agent $i\in\mathcal{N}$ and any stationary joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then the BR also satisfies the conditions in Lemma 4.2.
4.1 Induced Distributionally Robust MDPs
We then develop a framework of induced distributionally robust MDPs to provide a framework for multi-agent studies.
Definition 2 (Induced Distributionally Robust MDP).
Given a distributionally robust Markov game $\mathcal{MG}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{P},r)$ and an agent $i\in\mathcal{N}$. The induced distributionally robust MDP of the $\mathcal{MG}$ by a joint policy $\pi_{-i}$ is defined as $\mathsf{M}_{i}(\pi_{-i})\triangleq(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, where
$(\mathcal{P}^{\pi_{-i}})^{a_{i}}_{s}\triangleq\Big\{\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)\mathsf{P}\Big|\mathsf{P}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}\Big\}$,
$r^{\pi_{-i}}(s,a_{i})=\sum_{a_{-i}\in\mathcal{A}_{-i}}\pi_{-i}(a_{-i}|s)r_{i}(s,(a_{i},a_{-i})).$ (5)
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, due to being induced, irreducibility, convexity, and compactness are all preserved. Namely, if the DR-MG $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 1, then the induced distributionally robust MDP $\mathsf{M}_{i}(\pi_{-i})$ satisfies Assumption 1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
However, even the single-agent average reward robust RL is not fully studied. Specifically, the most essential tool, its robust Bellman equation, is not well studied: its solvability or characterizations of solutions are not yet justified (an incorrect proof is given in (wang2023robust)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Although a huge body of existing work develops convergent algorithms, their convergence guaranties either rely on the wrong result in (wang2023robust), e.g., (wang2023model; Roch2025reduction), or bypass its solvability (xu2025finite; xu2025efficient; chatterjee2023solving; meggendorfer2025solving). We hence derive the solvability of the robust Bellman equation under Assumption 1 as follows.
Theorem 4.1.
(1). (Robust Bellman Equation). Under Assumption 1, denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the gain/bias of $(\pi,\mathsf{P}_{w}):$ $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is the unique solution to the equation:
$V(s)=\sum_{a}\pi(a|s)(r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)),$ (6)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$; The uniqueness is in the sense that, for any solution $(g,V)$ to (6), it holds that $g=g^{\pi}_{\mathcal{P}},V=h^{\pi}_{\mathsf{P}_{w}}+c\textbf{e}$, for some constant $c$ and $\textbf{e}=(1,...,1).$
(2). (Optimal Robust Bellman Equation). The optimal robust Bellman equation also has a solution:
$V(s)=\max_{a}\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\}.$ (7)
Our results therefore prove the solvability of the robust Bellman equation, especially (6), solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (6) may have non-unique solutions even up to constant vectors (see discussion in wang2023model). Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward (puterman2014markov), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.
These results and connections are of fundamental importance in showing the existence of a Nash Equilibrium.
4.2 Existence of a Nash Equilibrium
With the results derived for single-agent robust Bellman equations, we can then show the following results.
Theorem 4.2.
For any distributionally robust MDP under ˜1, the optimal policy set $\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$ is convex.
As $\text{BR}_{i}(\pi_{-i})$ is the set of optimal robust policies of the (induced) distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$, this result hence directly implies the convexity of $\text{BR}_{i}(\pi_{-i})$. We highlight that this proof is not a direct extension from its counterpart under the discounted reward as we need to tackle the non-contraction of the Bellman equation, whereas the discounted Bellman equation is a contraction and admits a unique solution. Also, compared to non-robust studies, we need to additionally tackle the non-linearity and uncertainty transition kernel introduced by the worst-case consideration. We also highlight that our proof heavily relies on the solvability and uniqueness (up to constant vectors) of (6), which enables us to fully characterize the optimal policies through the optimal robust Bellman equation.
We further show the hemi-continuity of the optimal policy set of the induced distributionally robust MDP.
Theorem 4.3.
Under ˜1. For convergent policy sequences $\{\pi^{n}_{-i}\}\to\pi_{-i}$ and $\{\mu^{n}_{i}\}\to\mu_{i}$, if $\mu^{n}_{i}$ is an optimal robust policy of the induced distributionally robust MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi^{n}_{-i}},r^{\pi^{n}_{-i}})$ for agent $i\in\mathcal{N}$, then $\mu_{i}$ is also optimal for $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$.
Note that in this result, the inducing policy $\pi^{n}_{-i}$ varies, resulting in a distributionally robust MDP sequence with different dynamics. Hence, we need to reveal the inherent continuity of induced distributionally robust MDPs.
Remark 2.
The proofs of these results rely on Theorem˜4.1, i.e., the correspondence between the optimal policies and the solutions to the Bellman equations, and the fact that the Bellman equation for any policy $\pi$ has a unique solution. In the next section where we study a weaker setting, such results do not hold anymore, and the proof derived cannot be applied there.
Combining all of the above results, we can finally show the existence of a Nash Equilibrium.
Theorem 4.4.
(Existence of a Nash Equilibrium) For an average-reward DR-MG satisfying Assumption 1, a stationary robust Nash equilibrium exists.
Our result herein implies the solvability of DR-MGs in the average-reward setting. The existence of an equilibrium is the most fundamental property of a game, thus proving its existence is a necessary prerequisite for any further analysis or algorithmic design.
5 NE of Weakly Communicating DR-MGs
In this section, we extend our existence result from irreducible DR-MGs to a more general class of weakly communicating DR-MGs. The weakly communicating assumption is known to be the weakest standard structural assumption under which efficient algorithms and sharp structural results are available for average-reward control; see, e.g., (wikecek2015stationary; bertsekas2011dynamic; wan2021learning).
Assumption 2 (Weakly communicating DR-MG).
The uncertainty set $\mathcal{P}$ is convex and compact. Moreover, for any transition kernel $P\in\mathcal{P}$ and any player $i$, the state space can be decomposed as $\mathcal{S}=\mathcal{S}^{i}_{0}\cup\mathcal{S}^{i}_{1},$ with $\mathcal{S}^{i}_{0}\cap\mathcal{S}^{i}_{1}=\emptyset,$ such that: (1). Every state in $\mathcal{S}^{i}_{0}$ is transient under every stationary joint policy; (2). For any two states $s,s^{\prime}\in\mathcal{S}^{i}_{1}$ and any stationary policy $\pi_{-i}$ of the other players, there exist a stationary policy $\pi_{i}$ for player $i$ and an integer $N$ (possibly depending on $s,s^{\prime},\pi_{-i},P$) such that $\mathbb{P}\big(S_{N}=s^{\prime}\mid S_{0}=s,(\pi_{i},\pi_{-i}),P\big)>0.$
As in the non-robust case, weakly communicating is strictly weaker than irreducibility: it only requires accessibility under some policy, rather than under all joint policies. Under Assumption 2, only the optimal robust average reward $g^{\star}_{\mathcal{P}}(s)$ is constant across initial states, whereas the robust average reward $g^{\pi}_{\mathcal{P}}(s)$ of a general stationary policy $\pi$ may depend on $s$; see, e.g., (puterman2014markov; atia2021steady) for the non-robust case and wang2025bellman for the robust extension.
Moreover, the solvability guarantees for weakly communicating robust Bellman equations are weaker. The solvability of the optimal robust Bellman equation under the weakly communicating assumption was recently established by wang2025bellman.
Lemma 5.1 (Optimal robust Bellman equation (wang2025bellman)).
Let an average-reward DR-MDP satisfy Assumption 2. Then the optimal robust Bellman equation
$V(s)\;=\;\max_{a\in\mathcal{A}}\Big\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\Big\},\qquad s\in\mathcal{S},$ (8)
admits a solution $(g^{\star},u^{\star})$, with $g^{\star}=g^{\star}_{\mathcal{P}}(\rho)$ equal to the optimal robust average reward for any initial distribution $\rho$. Moreover, $g^{\star}$ is the optimal min–max value $\sup_{\pi}\inf_{P\in\mathcal{P}}g^{\pi}_{P}$, and strong min–max duality holds.
Compared with the irreducible case (Theorem 4.1), these results are strictly weaker. Theorem 4.1 gives solvability of the robust Bellman equation for every stationary policy $\pi$, as well as a tight correspondence between optimal policies and solutions of the optimal Bellman equation. Under Assumption 2, Lemma 5.1 guarantees solvability only for the optimal equation (8); the equation associated with a fixed policy $\pi$ may fail to have a solution, and one can no longer characterize optimal policies solely via the Bellman equations. As a consequence, the proof strategy of irreducible DR-MGs cannot be directly reused.
To address this issue, we construct a proxy set-valued map, derived from solutions of the optimal robust Bellman equation, and show that this map is contained in the true best-response map. Fix a player $i$ and a stationary joint policy $\pi_{-i}$ of all other players, and consider the induced single-agent DR-MDP $M_{i}(\pi_{-i})=(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i})$. Applying Lemma 5.1 to $M_{i}(\pi_{-i})$ yields a solution $(g^{\star}_{\pi_{-i}},u^{\star}_{\pi_{-i}})$ of the optimal Bellman equation for this induced DR-MDP, with $g^{\star}_{\pi_{-i}}$ equal to the optimal robust average reward $g^{\star}_{\pi_{-i}}\;=\;\sup_{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\nu}_{P}.$
For each state $s\in\mathcal{S}$ and each joint policy $\nu\in\Delta(\mathcal{A}_{i})$ of player $i$, define the one-step Bellman functional as
$F^{\pi_{-i}}_{s}(\nu)\;:=\;\sum_{a_{i}\in\mathcal{A}_{i}}\nu(a_{i})\Big(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}_{\pi_{-i}}+\sigma_{(\mathcal{P}^{\pi_{-i}})_{s,a_{i}}}(u^{\star}_{\pi_{-i}})\Big),$ (9)
where $r^{\pi_{-i}}_{i}(s,a_{i})=\sum_{a_{-i}}\pi_{-i}(a_{-i}\mid s)\,r_{i}(s,(a_{i},a_{-i}))$ and $(\mathcal{P}^{\pi_{-i}})_{s,a_{i}}$ is the induced uncertainty slice at $(s,a_{i})$. Intuitively, $F^{\pi_{-i}}_{s}(\nu)$ is the robust one-step Bellman return at state $s$ when player $i$ plays $\nu(\cdot\mid s)$ and the others follow $\pi_{-i}$. We then define the Bellman-greedy response set of player $i$ w.r.t. $\pi_{-i}$ by
$\mathcal{G}_{i}(\pi_{-i})~\triangleq~\bigl\{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i}):\;F^{\pi_{-i}}_{s}(\nu(\cdot\mid s))=\max_{\mu\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\mu),\ \forall s\in\mathcal{S}\bigr\}.$ (10)
Thus $\mathcal{G}_{i}(\pi_{-i})$ consists of stationary policies that are pointwise maximizers of the optimal Bellman right-hand side for the induced DR-MDP $M_{i}(\pi_{-i})$.
By Lemma 5.1 and strong duality for weakly communicating DR-MDPs (wang2025bellman, Thm. 2), the maxima in (9) are attained for every $s$, so $\mathcal{G}_{i}(\pi_{-i})$ is nonempty. Unlike in the irreducible case, however, the presence of transient states means that a policy greedy w.r.t. a Bellman solution need not be globally optimal; in particular, not every robust best response arises as a maximizer of (9). Nonetheless, we will show that
$\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i}),$
so $\mathcal{G}_{i}(\pi_{-i})$ is a subset of the best-response set that is still rich enough to establish existence of Nash equilibria.
Finally, define the joint map
$\mathcal{G}(\pi)\;\triangleq\;\times_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$
on the compact convex policy space $\mathcal{X}=\times_{i\in\mathcal{N}}\times_{s\in\mathcal{S}}\Delta(\mathcal{A}_{i})$. If each $\mathcal{G}_{i}$ satisfies the conditions of Kakutani’s fixed-point theorem (Lemma 4.2), then $G$ admits a fixed point $\pi^{\star}\in G(\pi^{\star})$, and the inclusion $\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i})$ guarantees that this $\pi^{\star}$ is also a fixed point of the best-response map BR and hence a robust NE.
We summarize the structural properties of the maps $\mathcal{G}_{i}$ in the next theorem; the proof is deferred to Appendix D.
Theorem 5.1 (NE of weakly communicating DR-MGs).
Suppose the average-reward DR-MG satisfies Assumption 2. Then, for every player $i\in\mathcal{N}$, the following hold:
  1. 1.
    For every stationary joint policy $\pi_{-i}$ of the other players, the set $\mathcal{G}_{i}(\pi_{-i})$ is nonempty, compact, and convex.
  2. 2.
    The set-valued map $\mathcal{G}_{i}$ is upper hemi-continuous in the product topology.
  3. 3.
    For every stationary joint policy $\pi_{-i}$, any Bellman-greedy response is a robust best response: $\mathcal{G}_{i}(\pi_{-i})\subseteq\text{BR}_{i}(\pi_{-i}).$
Consequently, the joint map $\mathcal{G}(\pi)=\times_{i}\mathcal{G}_{i}(\pi_{-i})$ satisfies Kakutani’s conditions (Lemma 4.2), admits a fixed point $\pi^{\star}\in G(\pi^{\star})$, and any such $\pi^{\star}$ is a stationary robust Nash equilibrium of the DR-MG.
Our results hence imply that DR-MGs with average reward is also solvable, i.e., have stationary robust NE, even under the weakly communicating settings. Due to the complicated structures, our proof techniques are fundamentally different from the irreducible case, and are deferred to Appendix D.
6 Algorithmic Solutions
In this section, we develop algorithmic solutions for irreducible average-reward DR-MGs under ˜1. Notably, finding stationary NE is significantly challenging in general-sum games, even under non-robust settings (hu2003nash; li2007reinforcement; li2003learning; zinkevich2005cyclic).
6.1 Nash Iteration
Inspired by the widely used value iteration algorithm, we first develop our robust Nash-Iteration in Algorithm 1.
Algorithm 1 Robust Nash-Iteration
Initialization: $h_{i}=0,\forall i\in\mathcal{N}$
while TRUE do
  for all $i\in\mathcal{N}$ do
   $h^{0}_{i}\leftarrow h_{i}$
   for all $s,a\in\mathcal{S}\times\mathcal{A}$ do
    $Q_{i}(s,a,h_{i}^{0})\leftarrow r_{i}(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(h_{i}^{0})$
    $\pi(s)\leftarrow\textbf{NE}(Q_{i}(s,a,h^{0}_{i}))$
    $h_{i}(s)\leftarrow\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i}^{0})]$
   end for
  end for
  if $\max_{i}\{\text{sp}(h_{i}-h^{0}_{i})\}=0$ then
   BREAK
  end if
end while
Output: $\pi$
Value-iteration has been extensively studied in algorithmic Markov game studies (hu2003nash; li2007reinforcement; li2003learning; liu2021sharp; shi2024sample; farhat2025online). Our algorithm maintains an estimation, $Q_{i}$, for all agents, and in each step, the algorithm finds a Nash Equilibrium for the matrix-form game $\{Q_{i}(s,a,h^{0}_{i})\},i\in\mathcal{N}$ through an oracle NE. Notably, due to the non-uniqueness of Nash Equilibria and the complicated structure of general-sum MGs, the algorithm is not guaranteed to converge if an arbitrary oracle NE is used. Therefore, additional assumptions on it are required, even in non-robust MG studies (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). We thus adopt these assumptions and derive the convergence guarantee as follows.
Theorem 6.1.
Under ˜1 and some additional assumptions (see Assumption 3 in Appendix), Algorithm˜1 converges to a robust NE.
6.2 Proximal Descent Algorithm
Despite the convergence guarantee we obtained in Theorem˜6.1, the algorithm suffers from two disadvantages. Firstly, the convergence is only guaranteed with additional assumptions, which are standard but may not hold generally (filar2012competitive); Moreover, in each step, Algorithm˜1 solves for a NE of a matrix game, which can be PPAD-complete in the worst case (daskalakis2009complexity; chen2009settling). Although assuming such an oracle is also standard in Markov game studies (liu2021sharp; hu2003nash; shi2024sample), yet we are motivated to find other algorithms with better computational complexity.
In this section, we propose another algorithm to address the two disadvantages. Our algorithm is based on an equivalent characterization of the robust NE, in terms of the temporal difference (TD) errors. Specifically, given a joint policy $\pi$ and agent $k$’s robust average reward and robust bias $(g_{k}^{\pi},V_{k}^{\pi})$ (unique up to constant vector, per Theorem˜4.1), define a TD error as
$\Omega_{R}^{k}(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi})\triangleq\ g_{k}^{\pi}+V_{k}^{\pi}(s)-\sum_{a_{-k}}\pi_{-k}(a_{-k}\mid s)(r_{k}(s,a_{-k},a_{k})+\sigma_{\mathcal{P}_{s}^{a_{-k},a_{k}}}(V_{k}^{\pi})).$
Note that from (6), it holds that $\sum_{a_{k}}\pi_{k}(a_{k}|s)\Omega_{R}^{k}(s,a_{k};\pi_{-k}|V_{k}^{\pi},g_{k}^{\pi})=0,\forall k,s.$ Moreover, we define
$G_{k,s}(\pi)\triangleq-\min_{a_{k}}\Omega_{R}^{k}(s,a_{k};\cdot),\quad G(\pi)\triangleq\sum_{k=1}^{K}\sum_{s\in S}G_{k,s}(\pi).$
We then develop a characterization of robust NE as follows.
Lemma 6.1.
For any policy $\pi$, $G(\pi)\geq 0$. Moreover, $G(\pi)=0$ if and only if $\pi$ is a robust NE.
This result directly implies that finding a robust NE is equivalent to solve the equation $G(\pi)=0$. Moreover, since $G(\pi)\geq 0$, and $G$ is entry-wise convex, identifying a robust NE can be reduced to a minimization problem: $\min_{\pi}G(\pi)$. Such a characterization and transform is inspired by TD-based reinforcement learning approaches (sutton2018reinforcement) and their applications in standard Markov games (sahabandu2024rl).
However, solving such a multi-agent optimization problem may still be challenging, especially due to the inherent instability of multi-agent algorithm (albrecht2024multi). To address this issue and enable stable convergence, we design a block-wise proximal descent algorithm. Specifically, we index blocks by player-state pairs $(k,s)\in\mathcal{B}\triangleq\mathcal{N}\times\mathcal{S}$, where each block variable is the probability vector $\pi_{k}(\cdot\mid s)\in\Delta(\mathcal{A}_{k})$. We fix a total order $\prec$ on $\mathcal{B}$ (e.g., lexicographic by states then players). For a block $b\equiv(k,s)$, we define the sets of earlier and later blocks:
$\mathcal{B}_{<k,s>}\triangleq\{(i,t)\in\mathcal{B}:\ (i,t)\prec(k,s)\},$ (11)
$\mathcal{B}_{>k,s>}\triangleq\{(i,t)\in\mathcal{B}:\ (i,t)\succ(k,s)\}.$ (12)
Our algorithm is then updated block-wisely. Namely, when update for a block, we will set the earlier block variables to the updated ones, and apply a one-step proximal descent (i.e., mirror descent) to the optimistic estimation of $G$, $\widehat{G}_{n}\big(\pi^{n+1}_{<k,s>},\mu,\pi^{n}_{>k,s>}\big)$ (see its definition in (114)). We note that such a design is similar to the optimistic iteration studied in two-player zero-sum games (wei2021last; cen2022faster), and is shown to be more stable than simultaneously updating (daskalakis2020independent; nekoei2023dealing). We then present our algorithm in Algorithm˜2. Our algorithm is inspired by the actor-critic structure in reinforcement learning (sutton2018reinforcement).
Algorithm 2 Proximal Descent for DR-MGs
1:Initialization: $\pi=\pi^{0}$, stepsizes $\{\eta_{n}\}_{n\geq 0}$.
2:for $n=0,1,2,\dots$ do
3:  For each $k$, solve (6) at $\pi^{n}$ to obtain $(g_{k}^{n},V_{k}^{n})$.
4:  for $(k,s)\in\mathcal{B}$ do
5:   $\pi^{n+1}_{k}(s)\ \in\ \arg\min_{\mu\in\Delta(\mathcal{A}_{k})}\big\{\frac{1}{2\eta_{n}}\,\big\|\mu-\pi^{n}_{k}(s)\big\|_{2}^{2}+\widehat{G}_{n}\big(\pi^{n+1}_{<k,s>},\mu,\pi^{n}_{>k,s>}\big)\big\}$
6:  end for
7:end for
Notably, each step of our Algorithm˜2 can be executed in a polynomial complexity. Specifically, solving (6) can be done with polynomial complexity, e.g., (xu2025finite); The proximal update step is a convex optimization per block, which is also polynomial-time. Compared to Algorithm˜1 which requires a NP-hard subroutine, our algorithm is more practical and feasible.
We then derive our convergence guarantee merely under ˜1, as follows.
Theorem 6.2.
Set $\sum_{n}^{\infty}\eta_{n}<\infty$. Then $G(\pi^{n})$ converges: $\lim_{n\to\infty}G(\pi^{n})=0,$ and $\pi^{n}$ weakly converges to a robust NE, i.e., every cluster point $\pi^{\star}$ of $\{\pi^{n}\}$ is a robust NE. Moreover, if the DR-MG only has finitely many robust NE, then $\{\pi^{n}\}$ converges to some NE.
We show that, with a suitable choice of stepsizes, the TD error converges to 0. We note that this convergence does not imply convergence of $\pi^{n}$, as $\pi^{n}$ can be cycling among distinct equilibria, while $G(\pi^{n})\to 0$. However, we show that any convergent subsequence of $\{\pi^{n}\}$ converges to a robust NE, and constructing such a subsequence is straightforward in fixed-point theory and monotone operator theory, by adopting techniques like Halpern Iteration (halpern_fixed_1967) or Tikhonov Regularization.
Remark 3.
Although our Algorithm˜2 enjoys a slightly weaker convergence guarantee than Algorithm˜1, it does not require any additional assumptions on the game structure or NE-computing oracle, and can be executed in a polynomial complexity. Our algorithm hence stands for the first practical algorithm for average reward DR-MGs with provably convergence guarantees.
7 Connections with Discounted DR-NE
In this section, we study the connection between distributionally robust games with discounted and average rewards. Our motivation is to utilize the extensively studied and mathematical elegance of the discounted setting to solve the more challenging average-reward setting. Recall that a policy $\pi^{\star}_{\gamma}$ is a Nash Equilibrium of the $\gamma$-discounted DR-MG if
$V^{\pi^{\star}}_{\mathcal{P},i}(\rho)\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i},$ (13)
where $V^{\pi^{\star}}_{\mathcal{P},i}(\rho)$ is the $\gamma$-discounted robust value function. A policy $\pi^{\star}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{\star}}_{\mathcal{P},i}(\rho)+\epsilon\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i},$ (14)
We then derive the following results for irreducible DR-MGs (˜1).
Theorem 7.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted DR-MG by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward DR-MG.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted DR-MG is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
(3). Moreover, define the robust diameter (auer2008near) of the DR-MG by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discounted robust NE for $\gamma=1-\frac{\epsilon}{D}$ is an $\epsilon$-NE under average reward.
This theorem establishes a crucial bridge between the discounted and average-reward criteria in DR-MGs. Specifically, we showed that the worst-case long-run average behavior of a multi-agent system can be approximated by its discounted behavior when the future is valued almost as much as the present (i.e., as $\gamma\to 1$).
Our results provide significant computational implications. The analysis of $\gamma$-discounted games is often more tractable (kardecs2011discounted), benefiting from properties like the Banach fixed-point theorem, which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as Q-learning and policy gradient, can be applied for the discounted setting (zhang2020robust). This result therefore assures us that for any desired level of accuracy $\epsilon$, we can simply select a single discount factor $\gamma$ sufficiently close to 1 and solve for this $\gamma$-discounted game.
8 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We proved the existence of a stationary Nash Equilibrium under both irreducible and weak communicating settings, and proposed two practical algorithms as the first provably convergent method for average reward DR-MGs. We also connected average-reward equilibria to their discounted counterparts. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
Appendix A Preliminaries
In this section we briefly review previous studies of single-agent distributionally robust MDPs.
A.1 Discounted Setting
Under the robust setting, the worst-case performance over the uncertainty set of MDPs is defined through the discounted reward. More specifically, the robust discounted value function of a policy $\pi$ for a discounted MDP is defined as
$V^{\pi}_{\mathcal{P},\gamma}(s)\triangleq\min_{\kappa\in\times_{t\geq 0}\mathcal{P}}\mathbb{E}_{\pi,\kappa}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right],$ (15)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\times_{t\geq 0}\mathcal{P}$.
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator (nilim2004robustness; iyengar2005robust; puterman2014markov):
$\mathbf{T}_{\pi}V(s)\triangleq\sum_{a\in\mathcal{A}}\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right),$ (16)
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{p\in\mathcal{P}^{a}_{s}}p^{\top}V$ is the support function of $V$ on $\mathcal{P}^{a}_{s}$. Based on the contraction of $\mathbf{T}_{\pi}$, robust dynamic programming approaches, e.g., robust value iteration, can be designed (nilim2004robustness; iyengar2005robust).
The goal is to find the optimal robust policy:
$\pi^{\star}=\arg\max_{\pi}V^{\pi}_{\mathcal{P},\gamma}(s),$ (17)
and it is shown (iyengar2005robust) that the optimal robust value function $V^{\pi^{\star}}_{\mathcal{P},\gamma}(s)$ satisfies the optimal robust Bellman equation
$V(s)\triangleq\max_{a}\left\{\pi(a|s)\left(r(s,a)+\gamma\sigma_{\mathcal{P}^{a}_{s}}(V)\right)\right\}.$ (18)
A.2 Average Reward Setting
For a fixed kernel $\mathsf{P}_{0}$, the average reward (or gain) and the bias are defined as
$g_{\mathsf{P}_{0}}^{\pi}(s)$ $\triangleq\liminf_{n\to\infty}\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right]$ (19)
and
$h_{\mathsf{P}_{0}}^{\pi}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}_{0}}\left[\sum^{\infty}_{t=0}(r_{t}-g_{\mathsf{P}_{0}}^{\pi})|s_{0}=s\right].$ (20)
We can view equation (20) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by (puterman2014markov), thus it holds the following non-robust Bellman equation:
$h^{\pi}_{\mathsf{P}_{0}}(s)=\mathbb{E}_{\pi}\left[r(s,a)-g^{\pi}_{\mathsf{P}}(s)+\sum_{s^{\prime}\in\mathcal{S}}p^{a}_{s,s^{\prime}}h_{\mathsf{P}_{0}}^{\pi}(s^{\prime})\right].$ (21)
In the robust setting, we consider the robust average reward, defined as
$g_{\mathcal{P}}^{\pi}(s)$ $\triangleq\min_{\kappa\in\times_{t\geq 0}\mathcal{P}}\liminf_{n\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right].$ (22)
It is shown that under the unichain assumption, the robust average reward is independent from the initial state: $g_{\mathcal{P}}^{\pi}(s_{1})=g_{\mathcal{P}}^{\pi}(s_{2})$, for any $s_{1},s_{2}$. The robust Bellman equation (for a policy $\pi$) and the robust optimal Bellman equation for the average reward setting are derived in (wang2023robust):
$V(s)+g$ $=\sum_{a}\pi(a|s)\big(r(s,a)+\sigma_{\mathcal{P}^{a}_{s}}(V)\big),$ (23)
$V(s)$ $=\max_{a}\big\{r(s,a)-g+\sigma_{\mathcal{P}^{a}_{s}}(V)\big\},\quad\forall s\in\mathcal{S}.$ (24)
It is shown (wang2023model) that if the equation (23) ((24)) has a solution $(g,V)$, then the solution $g$ is the robust average reward $g^{\pi}_{\mathcal{P}}$ (the optimal robust average reward $g^{*}_{\mathcal{P}}=\max_{\pi}g^{\pi}_{\mathcal{P}}$). However, whether these two equations are solvable is not studied but directly assumed.
Appendix B Proof of Lemma˜3.1
We derive the result by constructing an example.
Fix $\varepsilon\in(0,1)$. Consider a two-player zero-sum distributionally robust Markov game
$(\mathcal{S},\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{P},r_{1},r_{2}),$
with state space $\mathcal{S}=\{s,\,G,\,B\}$, action sets $\mathcal{A}_{1}=\{1,2\}$ and $\mathcal{A}_{2}=\{1,2\}$, and rewards
$r_{1}(G,a_{1},a_{2})=1,\qquad r_{1}(B,a_{1},a_{2})=0,\qquad r_{1}(s,a_{1},a_{2})=0,\qquad r_{2}=-r_{1}.$
The ambiguity set consists of two kernels $p^{(1)}$ and $p^{(2)}$:
$p^{(1)}(\,\cdot\,|s,(1,1))$ $=\delta_{G},$ $p^{(1)}(\,\cdot\,|s,(2,2))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,1))$ $=\varepsilon\,\delta_{G}+(1-\varepsilon)\,\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,2))$ $=\delta_{G},$
$p^{(1)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(1)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
$p^{(2)}(\,\cdot\,|s,(1,2))$ $=\delta_{B},$ $p^{(2)}(\,\cdot\,|s,(2,1))$ $=\delta_{B},$
and, for $x\in\{G,B\}$, $p^{(1)}(\,\cdot\,|x,(a_{1},a_{2}))=p^{(2)}(\,\cdot\,|x,(a_{1},a_{2}))=\delta_{x}$ (absorbing). Thus, under any $p\in\{p^{(1)},p^{(2)}\}$ the chain has two closed recurrent classes $\{G\}$ and $\{B\}$ (i.e., it is multi-chain).
Let the initial state be $S_{0}=s$. For stationary policies $\pi_{1}\in\Delta(\mathcal{A}_{1})$, $\pi_{2}\in\Delta(\mathcal{A}_{2})$, we specify them as
$\pi_{1}(1)=p\in[0,1],\qquad\pi_{2}(1)=q\in[0,1].$
Because $G,B$ are absorbing and $r_{1}\in\{0,1\}$ on $\{G,B\}$, the average reward of player 1 equals the probability (under the chosen kernel) of transitioning from $s$ to $G$ in one step. Thus we have that
$U(p,q)$ $\triangleq\min_{k\in\{1,2\}}\ \mathbb{P}_{p^{(k)}}\big(S_{1}=G\,\big|\,S_{0}=s,\,\pi_{1},\pi_{2}\big)$
$=\min\big\{\,p\,q+\varepsilon(1-p)(1-q),\ \varepsilon\,p\,q+(1-p)(1-q)\,\big\}.$
Then the robust average reward for player 1 is $U(p,q)$ and for player 2 is U(p,q)-U(p,q).
Lemma B.1.
For $U$ defined above and any $\varepsilon\in(0,1)$,
$\sup_{p\in[0,1]}\ \inf_{q\in[0,1]}U(p,q)\ =\ \frac{\varepsilon}{2},\qquad\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)\ =\ \varepsilon.$
Proof.
Fix $p\in[0,1]$. For $q\in[0,1]$, $U(p,q)$ is the pointwise minimum of two affine functions of $q$, hence concave in $q$. Therefore the infimum over $q$ is attained at an endpoint:
$\inf_{q\in[0,1]}U(p,q)=\min\{\,U(p,0),\,U(p,1)\,\}=\min\{\varepsilon(1-p),\,\varepsilon p\}=\varepsilon\min\{p,1-p\}.$
Maximizing over $p$ gives
$\sup_{p\in[0,1]}\inf_{q\in[0,1]}U(p,q)=\sup_{p\in[0,1]}\varepsilon\min\{p,1-p\}=\varepsilon\cdot\frac{1}{2}=\frac{\varepsilon}{2}.$
Fix $q\in[0,1]$. As a function of $p$, $U(p,q)$ is again the pointwise minimum of two affine functions, hence concave and piecewise linear with a single kink at the solution of equality:
$p\,q+\varepsilon(1-p)(1-q)=\varepsilon p\,q+(1-p)(1-q)\ \Longleftrightarrow\ p=1-q.$
Thus
$\sup_{p\in[0,1]}U(p,q)=\max\Big\{\,U(0,q),\ U(1,q),\ U(1-q,q)\,\Big\}=\max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
Hence
$\inf_{q\in[0,1]}\ \sup_{p\in[0,1]}U(p,q)=\min_{q\in[0,1]}\ \max\Big\{\,\varepsilon(1-q),\ \varepsilon q,\ (1+\varepsilon)q(1-q)\,\Big\}.$
For $q\in\{0,1\}$ the maximum equals $\varepsilon$. For $q\in(0,1)$, $(1+\varepsilon)q(1-q)>\varepsilon\min\{q,1-q\}$ and thus the inner maximum is at least $\max\{\varepsilon q,\varepsilon(1-q)\}\geq\varepsilon/2$, while near $q=0$ or $1$ it is strictly larger than $\varepsilon/2$. The minimum over $q$ is attained at $q=0$ or $q=1$, and equals $\varepsilon$. ∎
Theorem B.1 (Nonexistence of a Nash equilibrium).
For the robust Markov game above, no (stationary stochastic) Nash equilibrium exists.
Proof.
In a two-player zero-sum game, a stationary Nash equilibrium $(p^{\star},q^{\star})$ is a saddle point:
$U(p^{\star},q)\ \geq\ U(p^{\star},q^{\star})\ \geq\ U(p,q^{\star})\qquad\forall\,p,q\in[0,1],$
which implies
$\inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star}).$
Consequently,
$\sup_{p}\inf_{q}U(p,q)\ \geq\ \inf_{q}U(p^{\star},q)\ =\ U(p^{\star},q^{\star})\ =\ \sup_{p}U(p,q^{\star})\ \geq\ \inf_{q}\sup_{p}U(p,q),$
so a saddle requires $\sup_{p}\inf_{q}U=\inf_{q}\sup_{p}U$. By Lemma B.1, $\sup_{p}\inf_{q}U=\varepsilon/2$ while $\inf_{q}\sup_{p}U=\varepsilon$, leading to a contradiction. Hence no saddle point exists and no Nash equilibrium exists. ∎
Appendix C Irreducible DR-MGs
In this appendix we work in the single-agent setting and fix a stationary policy $\pi$. The state space $\mathcal{S}$ and action space $\mathcal{A}$ are finite. For each $(s,a)$, the local ambiguity set $\mathcal{P}_{s}^{a}\subset\Delta(\mathcal{S})$ is nonempty, compact and convex, and the overall ambiguity set is the rectangular product
$\mathcal{P}=\times_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}_{s}^{a}.$
For each $P\in\mathcal{P}$, we denote by $P^{\pi}$ the induced Markov kernel on $\mathcal{S}$:
$P^{\pi}(s^{\prime}\mid s)\;=\;\sum_{a\in\mathcal{A}}\pi(a\mid s)P(s^{\prime}\mid s,a).$
Under Assumption 1, $P^{\pi}$ is irreducible for every stationary $\pi$ and for every $P\in\mathcal{P}$.
For fixed $P$ and $\pi$, the average reward (gain) and bias are defined as
$g^{\pi}_{P}(s)=\liminf_{n\to\infty}\mathbb{E}^{\pi,P}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}r_{t}\,\middle|\,S_{0}=s\right],\qquad h^{\pi}_{P}(s)=\mathbb{E}^{\pi,P}\!\left[\sum_{t=0}^{\infty}\bigl(r_{t}-g^{\pi}_{P}\bigr)\,\middle|\,S_{0}=s\right].$
By irreducibility, $g^{\pi}_{P}(s)$ does not depend on $s$, so we write simply $g^{\pi}_{P}$; moreover $(g^{\pi}_{P},h^{\pi}_{P})$ is a solution to the (non-robust) Poisson/Bellman equation (puterman2014markov):
$g^{\pi}_{P}+h^{\pi}_{P}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}\in\mathcal{S}}P^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P}(s^{\prime}),\quad s\in\mathcal{S},$ (25)
unique up to an additive constant on the bias.
We define the robust average reward of $\pi$ as
$g^{\pi}_{\mathcal{P}}\;\triangleq\;\min_{P\in\mathcal{P}}g^{\pi}_{P}.$
Compactness of $\mathcal{P}$ and continuity of $P\mapsto g^{\pi}_{P}$ (see, e.g., Puterman, 2014, Prop. 8.4.6) imply that the minimum is attained: there exists at least one worst-case kernel
$P^{w}\in\arg\min_{P\in\mathcal{P}}g^{\pi}_{P}.$
For each state $s$ and action $a$, define the local robust one-step operator
$\sigma_{s}^{a}(v)\;\triangleq\;\min_{p\in\mathcal{P}_{s}^{a}}p^{\top}v,\qquad v\in\mathbb{R}^{\mathcal{S}},$
and the $\pi$-averaged robust Bellman operator
$\sigma^{\pi}_{s}(v)\;\triangleq\;\min_{P\in\mathcal{P}}\sum_{s^{\prime}\in\mathcal{S}}P^{\pi}(s^{\prime}\mid s)v(s^{\prime})\;=\;\min_{p\in\mathcal{P}_{s}^{\pi}}p^{\top}v,$
where $\mathcal{P}_{s}^{\pi}\triangleq\bigl\{\sum_{a}\pi(a\mid s)\,p^{a}:p^{a}\in\mathcal{P}_{s}^{a}\bigr\}$.
Theorem C.1.
Consider an uncertainty set $\mathcal{P}$ satisfying Assumption 1 and fix a stationary policy $\pi$. Let $P^{w}\in\arg\min_{P\in\mathcal{P}}g^{\pi}_{P}$ be any worst-case kernel, and let $(g^{\pi}_{P^{w}},h^{\pi}_{P^{w}})$ be the corresponding (non-robust) gain and bias, i.e. a solution to (25) with $P=P^{w}$. Then the pair $(g^{\pi}_{\mathcal{P}},h^{\pi}_{P^{w}})$ solves the robust Bellman equation
$g^{\pi}_{\mathcal{P}}+h(s)\;=\;r^{\pi}(s)+\sigma^{\pi}_{s}(h),\qquad s\in\mathcal{S},$ (26)
and is unique up to addition of a constant to $h$. Equivalently, writing $V=h$ and rearranging, $V$ satisfies
$V(s)\;=\;\sum_{a\in\mathcal{A}}\pi(a\mid s)\bigl(r(s,a)-g^{\pi}_{\mathcal{P}}+\sigma_{s}^{a}(V)\bigr),\quad s\in\mathcal{S},$
which is exactly equation (6) in the main text.
Proof.
Fix $\pi$ and let $P^{w}$ be any worst-case kernel for the robust average reward, so $g^{\pi}_{\mathcal{P}}=g^{\pi}_{P^{w}}$. By the irreducible average-reward MDP theory (puterman2014markov), there exists a bias $h^{\pi}_{P^{w}}$ such that
$g^{\pi}_{P^{w}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{w}{}^{\pi}(s^{\prime}\mid s)h^{\pi}_{P^{w}}(s^{\prime}),\qquad s\in\mathcal{S}.$ (27)
Now fix any vector $v\in\mathbb{R}^{\mathcal{S}}$ and define a kernel $P^{v}\in\mathcal{P}$ by taking, independently for each state $s$,
$P^{v,\pi}(\cdot\mid s)\in\arg\min_{p\in\mathcal{P}_{s}^{\pi}}p^{\top}v,$
which exists by compactness of $\mathcal{P}_{s}^{\pi}$. In particular, for $v=h^{\pi}_{P^{w}}$ we define $P^{V}\triangleq P^{h^{\pi}_{P^{w}}}$ so that
$\sigma^{\pi}_{s}(h^{\pi}_{P^{w}})\;=\;\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{w}}(s^{\prime}),\qquad s\in\mathcal{S}.$
By Assumption 1, $P^{V}{}^{\pi}$ is irreducible.
For the kernel $P^{V}$, denote by $(g^{\pi}_{P^{V}},h^{\pi}_{P^{V}})$ the corresponding non-robust gain/bias pair. It satisfies
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{V}}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{V}}(s^{\prime}),\qquad s\in\mathcal{S}.$ (28)
Rewriting (27)–(28) in vector form yields
$(\mathrm{I}-P^{w}{}^{\pi})h^{\pi}_{P^{w}}=r^{\pi}-g^{\pi}_{P^{w}}\mathbf{1},\qquad(\mathrm{I}-P^{V}{}^{\pi})h^{\pi}_{P^{V}}=r^{\pi}-g^{\pi}_{P^{V}}\mathbf{1}.$
Since $P^{w}$ is worst-case for $\pi$, we have $g^{\pi}_{P^{w}}\leq g^{\pi}_{P^{V}}$, and hence
$(\mathrm{I}-P^{w}{}^{\pi})h^{\pi}_{P^{w}}\;=\;r^{\pi}-g^{\pi}_{P^{w}}\mathbf{1}\;\geq\;r^{\pi}-g^{\pi}_{P^{V}}\mathbf{1}\;=\;(\mathrm{I}-P^{V}{}^{\pi})h^{\pi}_{P^{V}}.$ (29)
Equivalently,
$(P^{w}{}^{\pi}-\mathrm{I})h^{\pi}_{P^{w}}\;\leq\;(P^{V}{}^{\pi}-\mathrm{I})h^{\pi}_{P^{V}}.$
By the definition of $P^{V}$ as a minimizer for $h^{\pi}_{P^{w}}$ we also have, componentwise,
$P^{V}{}^{\pi}h^{\pi}_{P^{w}}\;\leq\;P^{w}{}^{\pi}h^{\pi}_{P^{w}}.$ (30)
Let $x\triangleq h^{\pi}_{P^{w}}-h^{\pi}_{P^{V}}$. Using (29) and adding/subtracting $P^{V}{}^{\pi}h^{\pi}_{P^{w}}$ we obtain
$x$ $=h^{\pi}_{P^{w}}-h^{\pi}_{P^{V}}\;\geq\;P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{V}}$
$=P^{V}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{V}}+\bigl(P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{w}}\bigr)$
$=P^{V}{}^{\pi}x+\bigl(P^{w}{}^{\pi}h^{\pi}_{P^{w}}-P^{V}{}^{\pi}h^{\pi}_{P^{w}}\bigr).$
By (30) the last term is componentwise nonnegative, so
$x\;\geq\;P^{V}{}^{\pi}x.$
Since $P^{V}{}^{\pi}$ is irreducible and stochastic, the inequality $x\geq P^{V}{}^{\pi}x$ forces $x$ to be constant, as we prove as follows.
let $s^{\star}$ be a state with minimal coordinate $x(s^{\star})=\min_{s}x(s)$. Then for any $k\geq 1$,
$x(s^{\star})\;\geq\;\sum_{s}(P^{V}{}^{\pi})^{k}(s^{\star},s)\,x(s)\;\geq\;x(s^{\star})\sum_{s}(P^{V}{}^{\pi})^{k}(s^{\star},s)\;=\;x(s^{\star}),$
so all inequalities must be equalities. Thus whenever $(P^{V}{}^{\pi})^{k}(s^{\star},s^{\prime})>0$ we must have $x(s^{\prime})=x(s^{\star})$. By irreducibility, every state is reachable from $s^{\star}$ under some power of $P^{V}{}^{\pi}$, so $x$ is constant: $x=c\mathbf{1}$ for some $c\in\mathbb{R}$, i.e.,
$h^{\pi}_{P^{w}}=h^{\pi}_{P^{V}}+c\,\mathbf{1}.$ (31)
Using (31) in the Poisson equation (28) for $P^{V}$ gives
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s)$ $=g^{\pi}_{P^{V}}+h^{\pi}_{P^{V}}(s)+c\;=\;r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{V}}(s^{\prime})+c$
$=r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P^{w}}(s^{\prime}).$
Equivalently,
$g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s).$ (32)
By construction of $P^{V}$ we have
$(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s)=\sigma^{\pi}_{s}(h^{\pi}_{P^{w}}).$
On the other hand, from the Poisson equation (27) for $P^{w}$,
$g^{\pi}_{P^{w}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+(P^{w}{}^{\pi}h^{\pi}_{P^{w}})(s)\;\geq\;r^{\pi}(s)+(P^{V}{}^{\pi}h^{\pi}_{P^{w}})(s)\;=\;g^{\pi}_{P^{V}}+h^{\pi}_{P^{w}}(s),$
where the inequality uses the minimizing property of $P^{V}$ for $h^{\pi}_{P^{w}}$. Thus $g^{\pi}_{P^{w}}\geq g^{\pi}_{P^{V}}$. Since $P^{w}$ is worst-case, $g^{\pi}_{P^{w}}\leq g^{\pi}_{P^{V}}$, and we conclude that $g^{\pi}_{P^{w}}=g^{\pi}_{P^{V}}$.
Combining this equality with (32), and recalling that $g^{\pi}_{P^{w}}=g^{\pi}_{\mathcal{P}}$, we obtain
$g^{\pi}_{\mathcal{P}}+h^{\pi}_{P^{w}}(s)\;=\;r^{\pi}(s)+\sigma^{\pi}_{s}(h^{\pi}_{P^{w}}),$
which is exactly (26) with $h=h^{\pi}_{P^{w}}$.
We then the uniqueness.
Suppose $(g,h)$ is any solution of (26). By the same construction as above, there exists some kernel $P^{V}\in\mathcal{P}$ such that
$\sigma^{\pi}_{s}(h)=(P^{V}{}^{\pi}h)(s),\quad s\in\mathcal{S}.$
Thus $(g,h)$ solves the non-robust Poisson equation for $(\pi,P^{V})$:
$g+h(s)=r^{\pi}(s)+\sum_{s^{\prime}}P^{V}{}^{\pi}(s^{\prime}\mid s)h(s^{\prime}).$
By uniqueness of the Poisson solution (up to constants) for irreducible chains, $g=g^{\pi}_{P^{V}}$ and $h=h^{\pi}_{P^{V}}+c\mathbf{1}$ for some $c$. Since any such $P^{V}$ is also worst-case for the average reward (by the same argument as above), we have $g=g^{\pi}_{\mathcal{P}}$ and $h=h^{\pi}_{P^{w}}+c\mathbf{1}$ for some worst-case kernel $P^{w}$. This proves the claimed uniqueness up to constants. ∎
Corollary C.1.
The optimal robust Bellman equation
$V(s)=\max_{a}\{r(s,a)-g(s)+\sigma_{\mathcal{P}^{a}_{s}}(V)\}$ (33)
is solvable.
Proof.
Denote the optimal robust policy by $\pi^{\star}$, and the corresponding worst-case kernel by $\mathsf{P}^{\star}$. By Theorem 4.1, $(g^{\pi^{\star}}_{\mathcal{P}},h^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is the solution to the robust Bellman equation of $\pi^{\star}$:
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=r^{\pi^{\star}}(s)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{\pi^{\star}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right).$ (34)
We now set $\pi^{\prime}(s)=\arg\max_{a}\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\},\forall s$, then we have that
$r^{\pi^{\prime}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})=g^{\pi^{\prime}}_{\mathcal{P}}\leq g^{\pi^{\star}}_{\mathcal{P}}$ $=\sum_{a}\pi^{\star}(a|s)\left(r(s,a)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right)$
$\leq r^{\pi^{\prime}}(s)-h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)+\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}),$ (35)
which further implies that
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}(s)\leq\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})-\sigma^{\pi^{\prime}}_{s}(h^{\pi^{\prime}}_{\mathsf{P}^{\prime}})\leq\mathsf{P}^{\prime}(h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}).$ (36)
Now denote $x=h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}$, it holds that
$x\leq\mu x,$ (37)
for some positive distribution matrix $\mu$. Similarly, it implies $x(s)=x(s^{\prime})$, and thus $h^{\pi^{\star}}_{\mathsf{P}^{\star}}-h^{\pi^{\prime}}_{\mathsf{P}^{\prime}}=ce$. This further implies that all inequalities in (35) are equations, thus
$h^{\pi^{\star}}_{\mathsf{P}^{\star}}(s)=\max_{a}\left\{r(s,a)-g^{\pi^{\star}}_{\mathcal{P}}+\sigma^{a}_{s}(h^{\pi^{\star}}_{\mathsf{P}^{\star}})\right\},$ (38)
hence $(h^{\pi^{\star}}_{\mathsf{P}^{\star}},g^{\pi^{\star}}_{\mathsf{P}^{\star}})$ is a solution to (7), completing the proof of the first part. The remaining proof follows in a similar way from (wang2023model). ∎
Lemma C.1.
If for any agent $i$ and any joint policy $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, convex, and $\text{BR}_{i}(\cdot)$ is upper hemi-continuous, then BR also satisfies the conditions in Lemma 4.2.
Proof.
(1). If for any $i$ and $\pi_{-i}$, $\text{BR}_{i}(\pi_{-i})$ is nonempty, then $\text{BR}(\pi)=(\text{BR}_{1}(\pi_{-1}),...,\text{BR}_{N}(\pi_{-N}))$ is clearly nonempty.
(2). Consider two policies $\pi^{1},\pi^{2}\in\text{BR}(\pi)$. To show the convexity of $\text{BR}(\pi)$, we need to show $\pi^{\prime}=\alpha\pi^{1}+(1-\alpha)\pi^{2}=(\alpha\pi^{1}_{1}+(1-\alpha)\pi^{2}_{1},...,\alpha\pi^{1}_{N}+(1-\alpha)\pi^{2}_{N})\in\text{BR}(\pi)$. Note that for any $i$, $\pi^{1}_{i},\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, and due to the convexity of $\text{BR}_{i}(\pi_{-i})$, $\alpha\pi^{1}_{i}+(1-\alpha)\pi^{2}_{i}\in\text{BR}_{i}(\pi_{-i})$, which completes the proof.
(3). Consider two convergent policy sequences $\{\pi^{n}\}\to\pi$ and $\{\mu^{n}\}\to\mu$, with $\mu^{n}\in\text{BR}(\pi^{n})$, we need to show that $\mu\in\text{BR}(\pi)$. Note that $\pi^{n}\to\pi$ implies that $\pi^{n}_{i}\to\pi_{i}$ for any $i$, and similarly $\mu^{n}\to\mu$. Since $\mu_{i}^{n}\in\text{BR}_{i}(\pi_{-i}^{n})$, by the hemi-continuity of $\text{BR}_{i}$, we have $\mu_{i}\in\text{BR}_{i}(\pi_{-i})$. Moreover note that
$\mu\in\text{BR}(\pi)\Leftrightarrow\mu_{i}\in\text{BR}_{i}(\pi_{-i}),\forall i,$ (39)
which hence completes the proof. ∎
C.1 Convexity
Lemma C.2.
Consider the optimal robust Bellman equation
$g+h(s)=\max_{a\in\mathcal{A}}\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\},\qquad s\in\mathcal{S},$ (40)
and let $(g,h)$ be any of its solutions. Let $\pi^{\star}$ be an optimal robust policy, i.e. $\pi^{\star}\in\arg\max_{\pi}g^{\pi}_{\mathcal{P}}$. Then $(g,h)$ also solves the robust Bellman equation associated with $\pi^{\star}$:
$g+h(s)=\sum_{a\in\mathcal{A}}\pi^{\star}(a\mid s)\bigl(r(s,a)+\sigma_{s}^{a}(h)\bigr),\qquad s\in\mathcal{S}.$ (41)
Proof.
By Theorem C.1, any optimal policy $\pi^{\star}$ admits a worst-case kernel $P^{\star}\in\mathcal{P}$ such that $(g^{\pi^{\star}}_{P^{\star}},h^{\pi^{\star}}_{P^{\star}})$ solves the fixed-policy equation (26) with $\pi=\pi^{\star}$. Moreover $g^{\pi^{\star}}_{P^{\star}}=g^{\star}_{\mathcal{P}}$ (the optimal robust average reward), and the optimal robust Bellman equation (40) also has a solution $(g,h)$ with $g=g^{\star}_{\mathcal{P}}$. We must show that $h$ and $h^{\pi^{\star}}_{P^{\star}}$ differ only by a constant.
Equation (40) can be written as
$g+h(s)=\max_{a}\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\}.$ (42)
Optimality of $\pi^{\star}$ implies that at each state $s$, the support of $\pi^{\star}(\cdot\mid s)$ lies in the set of maximizers of the right-hand side:
$\pi^{\star}(a\mid s)>0\;\Longrightarrow\;a\in\arg\max_{a^{\prime}}\{r(s,a^{\prime})+\sigma_{s}^{a^{\prime}}(h)\}.$
Therefore,
$g+h(s)=\sum_{a}\pi^{\star}(a\mid s)\bigl\{r(s,a)+\sigma_{s}^{a}(h)\bigr\},$ (43)
i.e., $(g,h)$ is a solution of the fixed-policy robust Bellman equation (41) with $\pi=\pi^{\star}$.
By Theorem C.1, the solution of the fixed-policy robust Bellman equation for $\pi^{\star}$ is unique up to constants and has gain $g=g^{\star}_{\mathcal{P}}=g^{\pi^{\star}}_{P^{\star}}$. Thus $h$ and $h^{\pi^{\star}}_{P^{\star}}$ differ only by an additive constant, and in particular $(g,h)$ is of the form
$g=g^{\star}_{\mathcal{P}},\qquad h=h^{\pi^{\star}}_{P^{\star}}+c\mathbf{1}.$
This hence completes the proof. ∎
Lemma C.3 (Convexity of optimal policies).
Define the optimal policy set
$\Pi^{\star}\triangleq\bigl\{\pi\;\big|\;g^{\pi}_{\mathcal{P}}=g^{\star}_{\mathcal{P}}\bigr\}.$
Then $\Pi^{\star}$ is convex.
Proof.
Let $\pi_{1},\pi_{2}\in\Pi^{\star}$ and let $\tilde{\pi}\triangleq\alpha\pi_{1}+(1-\alpha)\pi_{2}$ for some $\alpha\in[0,1]$, i.e.
$\tilde{\pi}(a\mid s)=\alpha\pi_{1}(a\mid s)+(1-\alpha)\pi_{2}(a\mid s),\quad s\in\mathcal{S},a\in\mathcal{A}.$
Let $(g,h)$ be any solution of the optimal robust Bellman equation (40). By Lemma C.2, $(g,h)$ also solves the fixed-policy robust Bellman equations for $\pi_{1}$ and $\pi_{2}$:
$g+h(s)=r^{\pi_{1}}(s)+\sigma^{\pi_{1}}_{s}(h),\qquad g+h(s)=r^{\pi_{2}}(s)+\sigma^{\pi_{2}}_{s}(h).$
Taking the convex combination with weights $(\alpha,1-\alpha)$ yields
$g+h(s)$ $=\alpha\bigl(r^{\pi_{1}}(s)+\sigma^{\pi_{1}}_{s}(h)\bigr)+(1-\alpha)\bigl(r^{\pi_{2}}(s)+\sigma^{\pi_{2}}_{s}(h)\bigr)$
$=\sum_{a}\tilde{\pi}(a\mid s)r(s,a)+\sum_{a}\tilde{\pi}(a\mid s)\sigma_{s}^{a}(h)=r^{\tilde{\pi}}(s)+\sigma^{\tilde{\pi}}_{s}(h).$
Thus $(g,h)$ also solves the fixed-policy robust Bellman equation for $\tilde{\pi}$. By Theorem C.1, the gain of any such solution must be the robust average reward of that policy; hence
$g^{\tilde{\pi}}_{\mathcal{P}}=g.$
Since $g=g^{\star}_{\mathcal{P}}$ is the maximal robust average reward (because $\pi_{1},\pi_{2}$ are optimal), we conclude that $\tilde{\pi}$ is also optimal, i.e. $\tilde{\pi}\in\Pi^{\star}$. Therefore $\Pi^{\star}$ is convex. ∎
The convexity of the best-response sets in the multi-agent case then follows by applying Lemma C.3 to the induced single-agent DR-MDPs $M_{i}(\pi_{-i})$, as discussed.
C.2 Hemi-continuity
Lemma C.4.
Let $\{\pi_{n}\}$ be a sequence of stationary policies converging to $\pi$ in the product topology. Let $P^{\pi_{n}}$ and $P^{\pi}$ denote the induced kernels for a fixed $P\in\mathcal{P}$, and let $(P^{\pi_{n}})^{\star}$ and $(P^{\pi})^{\star}$ be their stationary distributions. Then
$(P^{\pi_{n}})^{\star}\;\to\;(P^{\pi})^{\star}$
uniformly in $P\in\mathcal{P}$.
Proof.
This is a direct consequence of the continuity properties of irreducible Markov chains; see (puterman2014markov) (Prop. 8.4.6). Uniformity in $P$ follows from the compactness of $\mathcal{P}$. ∎
Lemma C.5.
Fix an agent $i$ and let $\pi:\mathcal{S}\to\Delta(\mathcal{A}_{-i})$ be a joint policy of all other agents. Define the operator $T^{\pi}_{\mathcal{P}}:\mathbb{R}^{\mathcal{S}}\to\mathbb{R}^{\mathcal{S}\times\mathcal{A}_{i}}$ by
$T^{\pi}_{\mathcal{P}}h(s,a_{i})\;\triangleq\;r^{\pi}(s,a_{i})+\sigma^{\pi}_{s,a_{i}}(h),$
where
$r^{\pi}(s,a_{i})\triangleq\sum_{a_{-i}}\pi(a_{-i}\mid s)\,r_{i}\bigl(s,(a_{i},a_{-i})\bigr),$
and
$\sigma^{\pi}_{s,a_{i}}(h)\triangleq\min_{q\in\mathcal{P}^{\pi}_{s,a_{i}}}q^{\top}h,\qquad\mathcal{P}^{\pi}_{s,a_{i}}\triangleq\left\{\sum_{a_{-i}}\pi(a_{-i}\mid s)P(\cdot\mid s,(a_{i},a_{-i})):P(\cdot\mid s,(a_{i},a_{-i}))\in\mathcal{P}_{s}^{(a_{i},a_{-i})}\right\}.$
Let $\{\pi_{n}\}$ be a sequence of opponent policies with $\pi_{n}\to\pi$, and $\{h_{n}\}$ a sequence of functions $h_{n}\to h$ uniformly on $\mathcal{S}$. Then
$T^{\pi_{n}}_{\mathcal{P}}(h_{n})\;\to\;T^{\pi}_{\mathcal{P}}(h)$
uniformly on $\mathcal{S}\times\mathcal{A}_{i}$.
Proof.
Recalling the definition of $\mathcal{T}_{\mathcal{P}}$ we can obtain the difference,
[…]
$=\max_{s,a}\big|r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-r^{\pi}(s,a)\big|$
$\leq\max_{s,a}\big\{|r^{\pi_{n}}(s,a)-r^{\pi}(s,a)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)|+|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|\big\}$
[…]
$B$ $=|\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h)-\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)|$
$=|\min_{q\in(\mathcal{P}^{\pi_{n}})^{a}_{s}}qh-\min_{q\in(\mathcal{P}^{\pi})^{a}_{s}}qh|$
$=|\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi_{n}}h-\min_{q\in\mathcal{P}^{a}_{s}}q^{\pi}h|$
$=\bigg|\min_{q\in\mathcal{P}^{a}_{s}}\Big(\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi_{n}(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big)-\min_{q\in\mathcal{P}^{a}_{s}}\Big(\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\pi(b|s)q(s^{\prime}|s,a,b)h(s^{\prime})\Big)\bigg|$
$\leq\bigg|\sum_{s^{\prime}\in\mathcal{S}}\sum_{b\in\mathcal{A}_{-i}}\Big(\pi_{n}(b|s)-\pi(b|s)\Big)q(s^{\prime}|s,a,b)h(s^{\prime})\bigg|,\quad\quad\text{for some }q\in\mathcal{P}^{a}_{s},$
$\leq||\pi_{n}(\cdot|s)-\pi(\cdot|s)||_{1}\cdot H,$
[…]
Note that $\pi_{n}(s)\rightarrow\pi(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},$ $\pi_{n}(s)-\pi(s)||_{1}\leq\epsilon$. Similarly with $h_{n}(s)\rightarrow h(s),\ \forall s\in\mathcal{S}$, then $\forall\epsilon,\ \exists N_{s}$ s.t. $\forall n>N_{s},\ |h_{n}(s)-h(s)|\leq\epsilon.$ Now by setting $N=\max_{s\in\mathcal{S}}\{N_{s}\}$, then $\forall n>N,$
[…]
Therefore, when $n>N,$ $\pi_{n}(s)-\pi(s)||_{1}+A+B\leq\epsilon+\epsilon+H\epsilon\lesssim\epsilon$. This implies that $\mathcal{T}^{n}_{\mathcal{P}}\rightarrow\mathcal{T}_{\mathcal{P}}$ uniformly on $\mathcal{S}\times\mathcal{A}$, which completes the proof. ∎
Theorem C.2 (Hemi-continuity).
Let $\pi_{n}\rightarrow\pi$ and $x_{n}\rightarrow x$, where $\pi_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{-i})$ and $x_{n}:\mathcal{S}\rightarrow\Delta(\mathcal{A}_{i})$. If $x_{n}\in\text{BR}_{i}(\pi_{n}),$ then $x\in\text{BR}_{i}(\pi).$
Proof.
To show that $x\in\text{BR}_{i}(\pi),$ it suffices to show that $x$ is an optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi},r^{\pi}).$ Recalling the robust Bellman equation, it is also sufficient to show that $g^{(x,\pi)}_{\mathcal{P}}$ satisfies the following:
$g(s)=\max_{a\in\mathcal{A}_{i}}\{r^{\pi}(s,a)+\sigma_{(\mathcal{P}^{\pi})^{a}_{s}}(h)-h(s)\}=\max_{a\in\mathcal{A}_{i}}\{\mathcal{T}^{\pi}_{\mathcal{P}_{s}^{a}}(h)\}.$ (45)
See that since $x_{n}\in\text{BR}_{i}(\pi_{n})$, then $x_{n}$ is the optimal policy of the induced AMDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{n}},r^{\pi_{n}}).$ Thus, there exists $h_{n}$, s.t. $(h_{n},g_{n}=g^{x_{n}}_{\mathcal{P}^{\pi_{n}}})$ is a solution to,
$g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g_{n}(s)=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}.$
See that $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=\min_{\mathsf{P}^{\pi_{n}}\in\mathcal{P}^{\pi_{n}}}g^{x_{n}}_{{}_{\mathsf{P}^{\pi_{n}}},r^{\pi_{n}}}=\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n},\pi_{n})}_{\mathsf{P},n}$, and so $g^{x_{n}}_{\mathcal{P}^{\pi_{n}}}=g^{\pi_{n},x_{n}}_{\mathcal{P}}$. Therefore, for any $(x_{n},\pi_{n})$, there exists some $h_{n}\in\mathbb{R}^{\mathcal{S}}$ such that for all $s\in\mathcal{S}$,
$g^{(x_{n},\pi_{n})}_{\mathcal{P}_{s}}(s)$ $=\max_{a}\{r^{\pi_{n}}(s,a)+\sigma_{(\mathcal{P}^{\pi_{n}})^{a}_{s}}(h_{n})-h_{n}(s)\}$
$=\max_{a}\{\mathcal{T}^{\pi_{n}}_{\mathcal{P}_{s}^{a}}(h_{n})\}.$ (46)
Moreover, since $h_{n}$ is some relative function (wang2023robust), $h_{n}||_{\infty}\leq H.$ Then, there exists some subsequence $\{n_{m}\},$ s.t. $h_{n_{m}}\rightarrow h$ for some $h$, and $g^{(x_{n},\pi_{n})}_{\mathcal{P}}\rightarrow g$ for some $g$. Then, we can take the limit as $m\rightarrow\infty$ in (C.2) to find:
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big\}$
$\overset{(i)}{=}\max_{a}\big\{\lim_{m\rightarrow\infty}\mathcal{T}_{\mathcal{P}}^{\pi_{n_{m}}}(h_{n_{m}})\big\}$
$=\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big\},$ (47)
where $(i)$ holds by the uniform convergence in Lemma C.5, and the final equality by $\pi_{n_{m}}\rightarrow\pi$, and $h_{n_{m}}\rightarrow h$. Looking at the left-hand side of equation (C.2) we find
$\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathcal{P}}$ $=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\lim_{m\rightarrow\infty}\min_{\mathsf{P}\in\mathcal{P}}\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\cdot r^{(x_{n_{m}},\pi_{n_{m}})}.$
By Lemma C.3 we have $\mathsf{P}^{\star}_{(x_{n_{m}},\pi_{n_{m}})}\rightarrow\mathsf{P}^{\star}_{(x,\pi)}$ uniformly. Thus, $\lim\min=\min\lim$ implies,
$\lim_{m\rightarrow\infty}g_{\mathcal{P}}^{(x_{n_{m}},\pi_{n_{m}})}$ $=\min_{\mathsf{P}\in\mathcal{P}}\lim_{m\rightarrow\infty}g^{(x_{n_{m}},\pi_{n_{m}})}_{\mathsf{P}}$
$=\min_{\mathsf{P}\in\mathcal{P}}g^{(x,\pi)}_{\mathsf{P}}$
$=g^{(x,\pi)}_{\mathcal{P}}$.
Therefore, $g^{(x,\pi)}_{\mathcal{P}}=\max_{a}\big\{\mathcal{T}_{\mathcal{P}}^{\pi}(h)\big\}$ for some $h\in\mathbb{R}^{\mathcal{S}}$, which completes the proof. ∎
Appendix D Weakly Communicating DR-MGs
Throughout this section we fix a player $i$ and a stationary joint policy $\pi_{-i}$ of the other players, and work with the induced single-agent DR-MDP
$M_{i}(\pi_{-i})\;=\;(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i}).$
All quantities defined below (transition sets, Bellman solutions, etc.) depend on $(i,\pi_{-i})$, but we often suppress this dependence to lighten notation.
Robust average reward and best responses.
For a stationary policy $\pi_{i}:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of player $i$, denote by $g^{\pi_{i}}_{P}$ its average reward in $M_{i}(\pi_{-i})$ under a kernel $P\in\mathcal{P}^{\pi_{-i}}$. We define the robust worst-case average reward and the associated best-response set as
$g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}$ $\triangleq\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\pi_{i}}_{P},$ (48)
$\text{BR}_{i}(\pi_{-i})$ $\triangleq\Bigl\{\pi_{i}:g^{\pi_{i}}_{\mathcal{P}^{\pi_{-i}}}=\sup_{\mu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\mu}_{P}\Bigr\}.$ (49)
Constant-gain robust Bellman equations.
Under Assumption 2 the induced DR-MDP $M_{i}(\pi_{-i})$ is weakly communicating, so we may invoke the main result of (wang2025bellman).
Lemma D.1 (Constant-gain robust Bellman equations, strong duality).
For the induced DR-MDP $M_{i}(\pi_{-i})$ there exists a constant-gain solution $(u^{\star},g^{\star})$ to the state-wise min–max Bellman equations
$u^{\star}(s)$ $=\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),$ (50)
$u^{\star}(s)$ $=\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),\qquad s\in\mathcal{S},$ (51)
with attainment of both the $\inf$ and the $\sup$ for every state $s$. Moreover $g^{\star}$ equals the optimal robust average reward of $M_{i}(\pi_{-i})$,
$g^{\star}=\sup_{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})}\min_{P\in\mathcal{P}^{\pi_{-i}}}g^{\nu}_{P},$ (52)
and the sup in (52) is attained.
Lemma D.1 is a direct specialization of wang2025bellman to the induced DR-MDP $M_{i}(\pi_{-i})$. In particular, $(u^{\star},g^{\star})$ is a solution of the optimal robust Bellman equation for $M_{i}(\pi_{-i})$, and $g^{\star}$ is the optimal min–max value.
D.1 Convexity
For each state $s\in\mathcal{S}$ define the local saddle Lagrangian
$L_{s}(\varphi,p_{s})~\triangleq~\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\bigl(r^{\pi_{-i}}_{i}(s,a_{i})-g^{\star}+p_{s,a_{i}}^{\top}u^{\star}\bigr),\quad\varphi\in\Delta(\mathcal{A}_{i}),\;p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}.$ (53)
We introduce the value functions
$F_{s}(\varphi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}(\varphi,p_{s}),$ (54)
$\Psi_{s}(p_{s})$ $\triangleq\sup_{\varphi\in\Delta(\mathcal{A}_{i})}L_{s}(\varphi,p_{s}).$ (55)
The function $F_{s}$ is concave and upper semicontinuous on the compact convex set $\Delta(\mathcal{A}_{i})$ (being the infimum of continuous affine maps), while $\Psi_{s}$ is convex and lower semicontinuous on $(\mathcal{P}^{\pi_{-i}})_{s}$.
By the duality relations (50)–(51),
$u^{\star}(s)=\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi)=\min_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\Psi_{s}(p_{s}),\qquad s\in\mathcal{S}.$ (56)
Let
$M_{s}\;\triangleq\;\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F_{s}(\varphi),$ (57)
which is nonempty, convex, and compact by standard convex analysis (boyd2004convex).
Lemma D.2 (Common supporting minimizer).
For each $s\in\mathcal{S}$ there exists $p_{s}^{\star}\in\arg\min_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}\Psi_{s}(p_{s})$ such that
$u^{\star}(s)=\Psi_{s}(p_{s}^{\star})=L_{s}\!\left(\varphi,p_{s}^{\star}\right)=F_{s}(\varphi)\qquad\forall\,\varphi\in\mathcal{M}_{s}.$ (58)
Proof.
By Berge’s Maximum Theorem (berge1877topological), $\Psi_{s}$ is continuous on compact $(\mathcal{P}^{\pi_{-i}})_{s}$, so a minimizer $p_{s}^{\star}$ exists. Then it holds that $\Psi_{s}(p_{s}^{\star})=u^{\star}(s)$. For any $\varphi\in\mathcal{M}_{s}$,
$u^{\star}(s)=F_{s}(\varphi)=\inf_{p_{s}}L_{s}\!\left(\varphi,p_{s}\right)\leq L_{s}\!\left(\varphi,p_{s}^{\star}\right)\leq\Psi_{s}(p_{s}^{\star})=u^{\star}(s).$ (59)
Equalities hold throughout. ∎
By $(s,a_{i})$-rectangularity, $p^{\star}\triangleq\{p_{s}^{\star}\}_{s\in\mathcal{S}}\in\mathcal{P}^{\pi_{-i}}$. And under ˜2, $p^{\star}$ admits a closed recurrent class $C_{p^{\star}}\subseteq\mathcal{S}$ and $\mathcal{S}\setminus C_{p^{\star}}$ is transient for all stationary policies.
Slack inequalities and characterization of ${G}_{i}$.
For a stationary policy $\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ of player $i$ define the Bellman slack
$\delta_{\nu}(s)\;\triangleq\;u^{\star}(s)-F_{s}\bigl(\nu(\cdot\mid s)\bigr)\;\geq\;0,\qquad s\in\mathcal{S},$ (60)
where the inequality follows from (56).
Recall the definitions (for fixed $i$ and $\pi_{-i}$)
$L_{s}(\phi,p_{s})$ $\triangleq\sum_{a_{i}\in A_{i}}\phi(a_{i})\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Bigr),\quad\phi\in\Delta(A_{i}),\ p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s},$ (61)
$F_{s}(\phi)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}L_{s}(\phi,p_{s}),\qquad\Psi_{s}(p_{s})\triangleq\sup_{\phi\in\Delta(A_{i})}L_{s}(\phi,p_{s}).$ (62)
By the duality relations (65)–(66), we have for every $s\in\mathcal{S}$
$u^{\star}(s)\;=\;\max_{\phi\in\Delta(A_{i})}F_{s}(\phi)\;=\;\min_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}\Psi_{s}(p_{s}).$ (63)
Let
$M_{s}\;\triangleq\;\operatorname{argmax}_{\phi\in\Delta(A_{i})}F_{s}(\phi),$
which is nonempty, convex, and compact.
Lemma D.1 already shows that there exists a common supporting minimizer $p_{s}^{\star}\in(\mathcal{P}_{\pi_{-i}})_{s}$ such that
$u^{\star}(s)=\Psi_{s}(p_{s}^{\star})=L_{s}(\phi,p_{s}^{\star})=F_{s}(\phi),\quad\forall\,\phi\in M_{s}.$ (64)
By $(s,a_{i})$-rectangularity, $p^{\star}\triangleq\{p_{s}^{\star}\}_{s\in\mathcal{S}}\in\mathcal{P}_{\pi_{-i}}$. Under Assumption 2, this kernel admits at least one closed recurrent class $C_{p^{\star}}\subseteq\mathcal{S}$, and every state outside $C_{p^{\star}}$ is transient under any stationary policy. [^2] This is the standard decomposition of a finite weakly communicating Markov decision process into transient states and recurrent communicating classes; see, e.g., puterman2014markov.
Lemma D.3 (Slack inequalities).
For a stationary policy $\nu:S\to\Delta(A_{i})$ define
$\delta_{\nu}(s)\;\triangleq\;u^{\star}(s)-F_{s}\bigl(\nu(\cdot\mid s)\bigr)\ \geq 0,\quad s\in\mathcal{S}.$
Let $(S_{t},A_{t})$ be the trajectory generated under $(\nu,p^{\star})$ starting from any initial distribution. Then
$\limsup_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]$  $\;\leq\;0,$ (65)
$\liminf_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]$ $\;\geq\;\liminf_{n\to\infty}\;\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr].$ (66)
Moreover, if $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$, then
$g^{\nu}_{p^{\star}}\;\triangleq\;\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})\Bigr]=\alpha^{\star}.$
Proof.
Step 1: upper bound. Fix any state $s\in\mathcal{S}$. By definition of $F_{s}$ and $u^{\star}(s)$ we have, for any action distribution $\phi\in\Delta(A_{i})$ and any $p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}$,
$L_{s}(\phi,p_{s})\;\geq\;F_{s}(\phi)\;\leq\;u^{\star}(s).$
In particular, plugging $\phi=\nu(\cdot\mid s)$ and the special kernel $p^{\star}_{s}$ from Lemma D.1 yields
$L_{s}\bigl(\nu(\cdot\mid s),p^{\star}_{s}\bigr)\;\leq\;\Psi_{s}(p^{\star}_{s})=u^{\star}(s).$
Writing this out,
$u^{\star}(s)\;\geq\;\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p^{\star\top}_{s,a_{i}}u^{\star}\Bigr).$
Under $(\nu,p^{\star})$, $A_{t}\sim\nu(\cdot\mid S_{t})$ and the next state satisfies $\mathbb{E}[u^{\star}(S_{t+1})\mid S_{t}=s,A_{t}=a_{i}]=p^{\star\top}_{s,a_{i}}u^{\star}$. Therefore
$u^{\star}(S_{t})\;\geq\;\mathbb{E}\bigl[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\bigr]\quad\text{a.s.}$ (67)
Taking unconditional expectation and summing from $t=0$ to n1n-1 gives
$\mathbb{E}[u^{\star}(S_{0})]\;\geq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\bigr)\Bigr].$
Rearrange the telescoping sum in $u^{\star}$:
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})-n\alpha^{\star}\Bigr]\;\leq\;\mathbb{E}[u^{\star}(S_{0})]-\mathbb{E}[u^{\star}(S_{n})].$
Divide by $n$ and use boundedness of $u^{\star}$ to get
$\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}\bigr)\Bigr]\;\leq\;\frac{\|u^{\star}\|_{\infty}+\|u^{\star}\|_{\infty}}{n}\to 0,$
which yields (73) after taking the $\limsup$.
Step 2: lower bound with slack. By definition of $\delta_{\nu}(s)$ and $F_{s}$,
$u^{\star}(s)-\delta_{\nu}(s)=F_{s}\bigl(\nu(\cdot\mid s)\bigr)=\inf_{p_{s}\in(\mathcal{P}_{\pi_{-i}})_{s}}\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p_{s,a_{i}}^{\top}u^{\star}\Bigr).$
In particular, evaluating the infimum at $p^{\star}_{s}$,
$u^{\star}(s)-\delta_{\nu}(s)\;\leq\;\sum_{a_{i}}\nu(a_{i}\mid s)\Bigl(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha^{\star}+p^{\star\top}_{s,a_{i}}u^{\star}\Bigr).$
As above, this translates to
$u^{\star}(S_{t})-\delta_{\nu}(S_{t})\;\leq\;\mathbb{E}\bigl[r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\,\big|\,S_{t}\bigr]\quad\text{a.s.}$ (68)
Taking expectations and summing from $t=0$ to n1n-1,
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(u^{\star}(S_{t})-\delta_{\nu}(S_{t})\bigr)\Bigr]\;\leq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\bigl(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star}+u^{\star}(S_{t+1})\bigr)\Bigr].$
Rearranging,
$\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]\;\geq\;\mathbb{E}\!\Bigl[\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]+\mathbb{E}[u^{\star}(S_{0})]-\mathbb{E}[u^{\star}(S_{n})].$
Divide by $n$ and pass to the $\liminf$ to obtain (74).
Step 3: equality under greedy actions on $C_{p^{\star}}$. Assume now that $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$. Then, by (61)–(62) and Lemma D.1,
$F_{s}\bigl(\nu(\cdot\mid s)\bigr)=u^{\star}(s),\quad s\in C_{p^{\star}},$
hence $\delta_{\nu}(s)=0$ on $C_{p^{\star}}$.
Under $(\nu,p^{\star})$, states in $S\setminus C_{p^{\star}}$ are transient, and $C_{p^{\star}}$ is a closed recurrent class (by Assumption 2 and the standard structure of finite Markov chains). Thus the fraction of time spent in $S\setminus C_{p^{\star}}$ converges to zero almost surely, while any invariant distribution is supported on $C_{p^{\star}}$ only. It follows that
$\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]=0.$
Combining this with the upper and lower bounds (73)–(74) gives
$\lim_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}r^{\pi_{-i}}_{i}(S_{t},A_{t})\Bigr]=\alpha^{\star},$
i.e., $g^{\nu}_{p^{\star}}=\alpha^{\star}$. ∎
Proposition D.1 (Characterization of G{G}).
Define
${G}_{i}(\pi_{-i})\;\triangleq\;\Bigl\{\nu:\mathcal{S}\to\Delta(A_{i})\ :\ \nu(\cdot\mid s)\in M_{s}\ \text{for all }s\in C_{p^{\star}}\Bigr\},$
where $C_{p^{\star}}$ is any closed recurrent class of the kernel $p^{\star}$ constructed in Lemma D.1. Then:
  1. 1.
    If $\nu\in\mathrm{BR}_{i}(\pi_{-i})$, every closed recurrent class $C_{p^{\star}}$ of $p^{\star}$ satisfies $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$.
  2. 2.
    If $\nu\in\mathcal{G}_{i}(\pi_{-i})$, then $\nu\in\mathrm{BR}_{i}(\pi_{-i})$.
Proof.
(1) Suppose $\nu\in\mathrm{BR}_{i}(\pi_{-i})$ and let $C_{p^{\star}}$ be any closed recurrent class of $p^{\star}$. If there existed $s_{0}\in C_{p^{\star}}$ with $\nu(\cdot\mid s_{0})\notin M_{s_{0}}$, then $F_{s_{0}}(\nu(\cdot\mid s_{0}))<u^{\star}(s_{0})$, so $\delta_{\nu}(s_{0})>0$. Since $s_{0}$ lies in a recurrent class of $p^{\star}$, the chain $(S_{t})$ under $(\nu,p^{\star})$ visits $s_{0}$ with positive asymptotic frequency, implying
$\liminf_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}\delta_{\nu}(S_{t})\Bigr]>0.$
Lemma D.3 then yields
$\limsup_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]<0,$
i.e., $g^{\nu}_{p^{\star}}<\alpha^{\star}$. On the other hand,
$\min_{p\in\mathcal{P}_{\pi_{-i}}}g^{\nu}_{p}\;\leq\;g^{\nu}_{p^{\star}}<\alpha^{\star},$
which contradicts the fact that $\alpha^{\star}$ is the optimal robust value $\sup_{\tilde{\nu}}\min_{p}g^{\tilde{\nu}}_{p}$ attained by $\nu\in\mathrm{BR}_{i}$. Hence no such $s_{0}$ can exist, and $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$.
(2) Conversely, assume $\nu\in\mathcal{G}_{i}(\pi_{-i})$, so $\nu(\cdot\mid s)\in M_{s}$ for all $s\in C_{p^{\star}}$. By Lemma D.3, $g^{\nu}_{p^{\star}}=\alpha^{\star}$. For any $p\in\mathcal{P}_{\pi_{-i}}$, the same one-step inequality as in the proof of Lemma D.3 (but now with $p$ in place of $p^{\star}$ and again using $F_{s}(\nu(\cdot\mid s))\leq L_{s}(\nu(\cdot\mid s),p_{s})$) implies
$\liminf_{n\to\infty}\mathbb{E}\Bigl[\frac{1}{n}\sum_{t=0}^{n-1}(r^{\pi_{-i}}_{i}(S_{t},A_{t})-\alpha^{\star})\Bigr]\;\geq\;0,$
so $g^{\nu}_{p}\geq\alpha^{\star}$ for all $p\in\mathcal{P}_{\pi_{-i}}$. By definition of $\alpha^{\star}$ as the optimal min–max value,
$\min_{p\in\mathcal{P}_{\pi_{-i}}}g^{\nu}_{p}\;\leq\;\sup_{\tilde{\nu}}\min_{p}g_{p}^{\tilde{\nu}}=\alpha^{\star},$
and therefore $\min_{p}g^{\nu}_{p}=\alpha^{\star}$, i.e., $\nu$ is a robust best response to $\pi_{-i}$. ∎
Theorem D.1 (Convexity).
Define
$\mathcal{G}_{i}\ \triangleq\ \bigl\{\nu:\mathcal{S}\to\Delta(\mathcal{A}_{i})\ :\ \nu(\cdot\mid s)\in\mathcal{M}_{s}\ \ \forall s\in\mathcal{S}\bigr\}.$ (69)
Then $\mathcal{G}_{i}$ is nonempty, compact, and convex; and $\mathcal{G}_{i}\subseteq\text{BR}_{i}(\pi_{-i})$.
Proof.
Each $\mathcal{M}_{s}$ is nonempty, compact, convex; the product set is nonempty, compact, convex. The inclusion follows from Proposition˜D.1. ∎
D.2 Hemi-continuity
For each $s\in\mathcal{S}$ and $(u,\alpha)\in\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$ define
$L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha)$ $\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\Big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha+p_{s,a_{i}}^{\top}u\Big),$ (70)
$F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$ $\triangleq\inf_{p_{s}\in(\mathcal{P}^{\pi_{-i}})_{s}}L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha),$ (71)
$\mathcal{M}_{s}(\pi_{-i})$ $\triangleq\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u^{\star}_{\pi_{-i}},\alpha^{\star}_{\pi_{-i}}),$ (72)
$\mathcal{G}_{i}(\pi_{-i})$ $\triangleq\bigl\{\Delta:\mathcal{S}\to\Delta(\mathcal{A}_{i})\ :\ \Delta(\cdot\mid s)\in\mathcal{M}_{s}(\pi_{-i})\ \ \forall s\bigr\}.$ (73)
Here $(u^{\star}_{\pi_{-i}},\alpha^{\star}_{\pi_{-i}})$ denotes a constant-gain solution for the induced DR-MDP $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}}_{i})$ (wang2025bellman).
Lemma D.4 (Hausdorff continuity of induced slices).
For fixed $(s,a_{i})$, define $M(w)\triangleq\sum_{a_{-i}}w(a_{-i})\,\mathcal{P}^{(a_{i},a_{-i})}_{s}$, $w\in\Delta(\mathcal{A}_{-i})$. If $w_{n}\to w$ in $\ell^{1}$, then $M(w_{n})\to M(w)$ in the Hausdorff metric; consequently $(\mathcal{P}^{\pi^{n}_{-i}})^{\,s}_{a_{i}}\to(\mathcal{P}^{\pi_{-i}})^{\,s}_{a_{i}}$ whenever $\pi^{n}_{-i}(\cdot\mid s)\to\pi_{-i}(\cdot\mid s)$.
Proof.
Our proof will utilize the Kuratowski convergence, which is equivalent to the convergence with respect to the Hausdorff metric on sets, if the space is a compact metric space (kuratowski2014topology).
By ˜2, each $\mathcal{P}^{(a_{i},a_{-i})}_{s}$ is nonempty and compact. Consider a sequence $x_{n}\in M(w_{n})$ with $x_{n}\to x$, which can be rewritten as $x_{n}=\sum_{a_{-i}}w_{n}(a_{-i})\,q^{n}_{a_{-i}}$ with $q^{n}_{a_{-i}}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}$. By compactness, for each $a_{-i}$, there exists a convergent subsequence $q^{n}_{a_{-i}}\to q_{a_{-i}}\in\mathcal{P}^{(a_{i},a_{-i})}_{s}$. Through the diagonal argument, it holds that $x=\sum_{a_{-i}}w(a_{-i})\,q_{a_{-i}}\in M(w)$. On the other hand, fix $x=\sum_{a_{-i}}w(a_{-i})\,q_{a_{-i}}\in M(w)$ and set $x_{n}\triangleq\sum_{a_{-i}}w_{n}(a_{-i})\,q_{a_{-i}}\in M(w_{n})$; then $x_{n}\to x$. Hence the Kuratowski outer and inner limits coincide and, since the sets are compact, Hausdorff convergence follows (beer1993topologies). ∎
Lemma D.5 (Uniform continuity of min-envelope).
Let $(u_{n},\alpha_{n},\pi^{n}_{-i})\to(u,\alpha,\pi_{-i})$, with $\{u_{n}\}$ uniformly bounded. Then for each $s$,
$\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi_{-i}}_{s}(\varphi;u,\alpha)\big|\ \to\ 0.$ (74)
Proof.
For fixed $s$, the integrand $L_{s}^{\pi_{-i}}(\varphi,p_{s};u,\alpha)$ is continuous and affine in $(\varphi,p_{s},u,\alpha)$; $r^{\pi_{-i}}_{i}(s,a_{i})$ depends linearly on $\pi_{-i}(\cdot\mid s)$; $p_{s,a_{i}}^{\top}u$ is continuous. By Lemma D.4, the feasible sets $(\mathcal{P}^{\pi^{n}_{-i}})_{s}$ converge to $(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric. Berge’s Maximum Theorem (applied to the minimization of a continuous function over compact feasible sets that vary Hausdorff‑continuously) gives continuity of the value map $(\varphi,u,\alpha,\pi_{-i})\mapsto F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$. Compactness of $\Delta(\mathcal{A}_{i})$ yields uniform convergence in $\varphi$. ∎
Lemma D.6 (Argmax stability under uniform convergence (ross2013elementary)).
Let $K$ be compact and $f_{n},f:K\to\mathbb{R}$ with $f_{n}\to f$ uniformly. If $x_{n}\in\arg\max_{x\in K}f_{n}(x)$ and $x_{n}\to x$, then $x\in\arg\max_{x\in K}f(x)$.
Lemma D.7 (Stability of Bellman pairs).
Let $\{\pi^{n}_{-i}\}_{n\geq 1}\subset\times_{s}\Delta(\mathcal{A}_{-i})$ satisfy $\pi^{n}_{-i}\to\pi_{-i}$ (product topology). For each $n$, let $(u_{n},\alpha_{n})$ solve the attained sup–inf constant-gain equation
$u_{n}(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\,F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n}),\qquad s\in\mathcal{S},$ (75)
and normalize $u_{n}(s_{0})=0$ at a fixed $s_{0}\in\mathcal{S}$. Then:
  • (i)
    $0\leq\alpha_{n}\leq 1$ for all $n$, and $\|u_{n}\|_{\infty}\leq 1$ for all $n$.
  • (ii)
    There exists a subsequence (not relabeled) such that $(u_{n},\alpha_{n})\to(u,\alpha)$ in $\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$.
  • (iii)
    The limit $(u,\alpha)$ solves the sup–inf constant-gain equation at $\pi_{-i}$:
    $u(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\,F^{\pi_{-i}}_{s}(\varphi;u,\alpha),\qquad s\in\mathcal{S}.$ (76)
Proof.
Since $r\in[0,1]$, every long-run average reward lies in $[0,1]$, hence $0\leq\alpha_{n}\leq 1$.
Let $m_{n}\triangleq\min_{s}u_{n}(s)$ and $M_{n}\triangleq\max_{s}u_{n}(s)$. For any $s$, any $\varphi$, and any feasible $p_{s}$, it holds that
$-1+m_{n}\;\leq\;-1+\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(p_{s,a_{i}}^{\top}u_{n}\big)\;\leq\;L_{s}\!\left(\varphi,p_{s};u_{n},\alpha_{n}\right)\;\leq\;1+\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(p_{s,a_{i}}^{\top}u_{n}\big)\;\leq\;1+M_{n}.$ (77)
Taking $\inf_{p_{s}}$ and then $\sup_{\varphi}$ yields, for every $s$,
[…]
Therefore $M_{n}-m_{n}\leq 1$. With the normalization $u_{n}(s_{0})=0$ we have $m_{n}\leq 0$ and hence $M_{n}\leq 1$, so $|u_{n}(s)|\leq 1$ for all $s$, i.e., $\|u_{n}\|_{\infty}\leq 1$.
As we have shown that $\{u_{n}\}$ is bounded in $\mathbb{R}^{\mathcal{S}}$ and $\{\alpha_{n}\}\subset[0,1]$, there exists a convergent subsequence (ross2013elementary) (not relabeled) such that $(u_{n},\alpha_{n})\to(u,\alpha)$ in $\mathbb{R}^{\mathcal{S}}\times\mathbb{R}$.
We then consider a fixed $s\in\mathcal{S}$. For each $a_{i}$, the induced slice
$(\mathcal{P}^{\pi^{n}_{-i}})^{s}_{a_{i}}=\sum_{a_{-i}}\pi^{n}_{-i}(a_{-i}\!\mid s)\,\mathcal{P}^{(a_{i},a_{-i})}_{s}$ (79)
converges to $(\mathcal{P}^{\pi_{-i}})^{s}_{a_{i}}$ in the Hausdorff metric because the right-hand side is a Minkowski sum with convex weights that depend continuously on $\pi_{-i}(\cdot\mid s)$; hence $(\mathcal{P}^{\pi^{n}_{-i}})_{s}\to(\mathcal{P}^{\pi_{-i}})_{s}$ in the Hausdorff metric (Lemma˜D.4).
Moreover, for $(u,\alpha)$ in a bounded set and $\varphi\in\Delta(\mathcal{A}_{i})$, the map
$(\varphi,p_{s},u,\alpha,\pi_{-i})\ \mapsto\ L^{\pi_{-i}}_{s}(\varphi,p_{s};u,\alpha)\triangleq\sum_{a_{i}\in\mathcal{A}_{i}}\varphi(a_{i})\big(r^{\pi_{-i}}_{i}(s,a_{i})-\alpha+p_{s,a_{i}}^{\top}u\big)$ (80)
is continuous and affine in each argument; by Berge’s Maximum Theorem, the value function $(\varphi,u,\alpha,\pi_{-i})\mapsto F^{\pi_{-i}}_{s}(\varphi;u,\alpha)$ is continuous when the feasible sets vary Hausdorff-continuously. Consequently,
$F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})\ \longrightarrow\ F^{\pi_{-i}}_{s}(\varphi;u,\alpha),\qquad\forall\,\varphi\in\Delta(\mathcal{A}_{i}).$ (81)
For any $n$, $s$, and $\varphi,\psi\in\Delta(\mathcal{A}_{i})$,
$\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi^{n}_{-i}}_{s}(\psi;u_{n},\alpha_{n})\big|\leq\sup_{p_{s}\in(\mathcal{P}^{\pi^{n}_{-i}})_{s}}\Big|\sum_{a_{i}\in\mathcal{A}_{i}}(\varphi-\psi)(a_{i})\big(r^{\pi^{n}_{-i}}_{i}(s,a_{i})-\alpha_{n}+p_{s,a_{i}}^{\top}u_{n}\big)\Big|\leq 3\,\|\varphi-\psi\|_{1},$ (82)
because $r^{\pi^{n}_{-i}}_{i}\in[0,1]$, $\alpha_{n}\in[0,1]$, and $|p_{s,a_{i}}^{\top}u_{n}|\leq\|u_{n}\|_{\infty}\leq 1$. Thus the family $\{F^{\pi^{n}_{-i}}_{s}(\cdot;u_{n},\alpha_{n})\}_{n}$ is equi-Lipschitz on the compact $\Delta(\mathcal{A}_{i})$. By (81) and a standard $\varepsilon$–net argument (or Arzelà–Ascoli’s criterion for equicontinuous families on compact domains), we obtain the uniform convergence
$\sup_{\varphi\in\Delta(\mathcal{A}_{i})}\;\big|F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})-F^{\pi_{-i}}_{s}(\varphi;u,\alpha)\big|\ \longrightarrow\ 0.$ (83)
Finally, from (75) and (83),
$u_{n}(s)\;=\;\sup_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n})\ \longrightarrow\ \sup_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u,\alpha).$ (84)
Since $u_{n}(s)\to u(s)$, we conclude (76) for every $s\in\mathcal{S}$. This hence completes the proof. ∎
Theorem D.2 (Upper hemi-continuity of $\mathcal{G}_{i}(\cdot)$).
Let $\pi^{n}_{-i}\to\pi_{-i}$ in the product topology on stationary policies, and let $\Delta_{n}\in\mathcal{G}_{i}(\pi^{n}_{-i})$ for each $n$. If $\Delta_{n}\to\Delta$ pointwise (equivalently, in the product topology), then $\Delta\in\mathcal{G}_{i}(\pi_{-i})$.
Proof.
For each $n$, choose an attained constant‑gain solution $(u_{n},\alpha_{n})$ for $\mathsf{M}_{i}(\pi^{n}_{-i})$, normalized by $u_{n}(s_{0})=0$. By Lemma D.7, along a subsequence (not relabeled) $(u_{n},\alpha_{n})\to(u,\alpha)$ and $(u,\alpha)$ solves the limit equation for $\mathsf{M}_{i}(\pi_{-i})$.
Fix $s\in\mathcal{S}$. Because $\Delta_{n}\in\mathcal{G}_{i}(\pi^{n}_{-i})$, we have
$\Delta_{n}(\cdot\mid s)\in\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi^{n}_{-i}}_{s}(\varphi;u_{n},\alpha_{n}).$ (85)
By Lemma D.5, $F^{\pi^{n}_{-i}}_{s}(\cdot;u_{n},\alpha_{n})\to F^{\pi_{-i}}_{s}(\cdot;u,\alpha)$ uniformly on $\Delta(\mathcal{A}_{i})$. Since $\Delta_{n}(\cdot\mid s)\to\Delta(\cdot\mid s)$ and $\Delta(\mathcal{A}_{i})$ is compact, Lemma D.6 yields
$\Delta(\cdot\mid s)\in\arg\max_{\varphi\in\Delta(\mathcal{A}_{i})}F^{\pi_{-i}}_{s}(\varphi;u,\alpha)=:\mathcal{M}_{s}(\pi_{-i}).$ (86)
As $s$ was arbitrary, $\Delta\in\mathcal{G}_{i}(\pi_{-i})$, proving the upper hemi-continuity. ∎
D.3 Existence of DR-NE
We can then prove the existence of DR-NE following the results shown above. Define a map $\mathcal{G}(\pi)\triangleq\times_{i\in\mathcal{N}}\mathcal{G}_{i}(\pi_{-i})$ on the compact convex policy space $X=\times_{i,s}\Delta(\mathcal{A}_{i})\to 2^{X}$. Since each $\mathcal{G}_{i}(\cdot)$ is nonempty, compact, convex (Theorem˜D.1), and is upper hemi-continuous (Theorem˜D.2), thus $G$ satisfies Kakutani’s condition as in Lemma˜4.2. Hence there exists a fixed point $\pi^{\star}\in\mathcal{G}(\pi^{\star})\subseteq\times_{i}\text{BR}_{i}(\pi^{\star}_{-i})$, which is a stationary DR-NE.
Appendix E Proof of Section 6
E.1 Nash Iteration
Assumption 3.
In Algorithm 1, at each step, the Nash Equilibrium $\pi(s)=\textbf{NE}(Q_{i}(s,a,h^{i}))$ satisfies one of the following conditions:
  • $\pi$ is global optimal: $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\geq\mathbb{E}_{\pi^{\prime}}[Q_{i}(s,a,h^{i})]$ for all joint policy $\pi^{\prime}$,$i,s,a$.
  • $\pi$ is a saddle point such that $\mathbb{E}_{\pi}[Q_{i}(s,a,h^{i})]\leq\mathbb{E}_{(\pi^{\prime}_{-i},\pi_{i})}[Q_{i}(s,a,h^{i})]$, for all $i,s,a,$ and joint policy $\pi^{\prime}_{-i}$.
Moreover, we assume that $P^{a}_{s}>0$ for all $s,a$ and $P\in\mathcal{P}^{a}_{s}$.
The first part of our assumption is necessary for convergence under average reward Markov games (li2003learning; hu2003nash; littman2001friend; Bowling2001; bowling2000convergence). The second assumption is a technical assumption to ensure convergence, and can be removed by modifying the update rule by a multi-step operator, see (wang2023model; xu2025finite).
Theorem E.1.
Additionally assume Assumption 3. The robust Nash-Iteration algorithm converge to an average reward Nash Equilibrium of the distributionally robust game.
Proof.
We split the proof into two parts.
(i) Any limit point is an average-reward NE.
Suppose Algorithm 1 converges to a stationary joint policy $\pi$, and let $h_{i}$ be the limit of the bias estimates for player $i$. By the update rule in Algorithm 1, at iteration $k$,
$h_{i}^{k+1}(s)=\mathbb{E}_{\pi^{k}(s)}\bigl[Q_{i}(s,a,h_{i}^{k})\bigr],$
where $\pi^{k}(s)$ is the selected NE of the stage game with payoff matrix $Q_{i}(s,a,h_{i}^{k})$. At convergence we have $h_{i}^{k+1}\to h_{i}^{*}$ and $\pi^{k}\to\pi$, and the stopping condition $\max_{i}\operatorname{sp}(h_{i}^{k+1}-h_{i}^{k})=0$ implies that the difference $h_{i}^{k+1}-h_{i}^{k}$ converges to a constant vector $c_{i}\mathbf{1}$ for each $i$ (here $\operatorname{sp}(\cdot)$ denotes the span seminorm). Thus there exist constants $c_{1},\dots,c_{N}$ such that for all $s$ and $i$,
$\mathbb{E}_{\pi(s)}[Q_{i}(s,a,h_{i}^{*})]-h_{i}^{*}(s)=c_{i}.$ (87)
Unfolding the definition of $Q_{i}$,
$Q_{i}(s,a,h_{i}^{*})=r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i}^{*}),$
and taking the expectation in (95) gives
$h_{i}^{*}(s)+c_{i}=\sum_{a}\pi(a\mid s)\Bigl(r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i}^{*})\Bigr),\quad s\in\mathcal{S}.$ (88)
Equation (88) is the robust average-reward Bellman equation associated with the fixed policy $\pi$ and bias $h_{i}^{*}$. By the solvability result for the robust Bellman equation under Assumption 1 (Theorem 4.1 in the main text, based on the single-agent theory in wang2025bellman), it follows that $(c_{i},h_{i}^{*})=(g^{\pi}_{P,i},h^{\pi}_{P,i})$ for some robust bias $h^{\pi}_{P,i}$, i.e.,
$h^{\pi}_{P,i}+g^{\pi}_{P,i}\mathbf{1}=r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i}).$ (89)
Now assume, for contradiction, that $\pi$ is not an average-reward Nash equilibrium. Then there exists a player $i$ and an alternative policy $\mu_{i}$ such that
$g^{(\pi_{-i},\mu_{i})}_{P,i}>g^{(\pi_{-i},\pi_{i})}_{P,i}=g^{\pi}_{P,i}.$ (90)
By the same Bellman solvability result, there is a bias $h_{i}^{\mu}$ such that
$h_{i}^{\mu}+g^{(\pi_{-i},\mu_{i})}_{P,i}\mathbf{1}=r^{(\pi_{-i},\mu_{i})}_{i}+\sigma_{(\pi_{-i},\mu_{i})}(h_{i}^{\mu}).$ (91)
Combining (89)–(91) yields
$r^{(\pi_{-i},\mu_{i})}_{i}+\sigma_{(\pi_{-i},\mu_{i})}(h_{i}^{\mu})-h_{i}^{\mu}=g^{(\pi_{-i},\mu_{i})}_{P,i}>g^{\pi}_{P,i}=r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i})-h^{\pi}_{P,i}.$ (92)
On the other hand, since at each state the algorithm selects a Nash equilibrium $\pi(s)$ of the one-stage game with payoffs $Q_{i}(s,a,h^{\pi}_{P,i})$, the NE property implies
$r^{(\pi_{-i},\mu_{i})}_{i}+\sigma_{(\pi_{-i},\mu_{i})}(h^{\pi}_{P,i})\;\leq\;r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i}).$ (93)
Subtracting $h^{\pi}_{P,i}$ from both sides of (93), and comparing with (92), we obtain
$r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i})-h^{\pi}_{P,i}$ $<r^{(\pi_{-i},\mu_{i})}_{i}+\sigma_{(\pi_{-i},\mu_{i})}(h_{i}^{\mu})-h_{i}^{\mu}$
$\leq\sigma_{(\pi_{-i},\mu_{i})}(h_{i}^{\mu})-h_{i}^{\mu}+r^{\pi}_{i}+\sigma_{\pi}(h^{\pi}_{P,i})-\sigma_{(\pi_{-i},\mu_{i})}(h^{\pi}_{P,i}).$ (94)
Rearranging (94) gives
$h_{i}^{\mu}-h^{\pi}_{P,i}<\sigma_{(\pi_{-i},\mu_{i})}(h_{i}^{\mu})-\sigma_{(\pi_{-i},\mu_{i})}(h^{\pi}_{P,i})\;\leq\;P(h_{i}^{\mu}-h^{\pi}_{P,i})$ (95)
for some kernel $P\in\mathcal{P}$, where the last inequality uses the definition of the support function $\sigma_{(\pi_{-i},\mu_{i})}(\cdot)$ as the minimum over transition kernels. Concretely, there exists $P\in\mathcal{P}$ achieving the minimum such that $\sigma_{(\pi_{-i},\mu_{i})}(v)=Pv$ for any $v$, and then the difference of supports is bounded by $P(h_{i}^{\mu}-h^{\pi}_{P,i})$.
Let $x\coloneqq h_{i}^{\mu}-h^{\pi}_{P,i}$. Then (95) says that $x<Px$ componentwise. Since $\mathcal{P}$ satisfies Assumption 1, all kernels in $\mathcal{P}$ (and in particular this $P$) are irreducible. For an irreducible stochastic matrix $P$, any vector $x$ satisfying $x\leq Px$ (or $x\geq Px$) componentwise must be constant, i.e., $x=c\mathbf{1}$ for some scalar $c$. Hence
$h_{i}^{\mu}-h^{\pi}_{P,i}=c_{i}\mathbf{1}.$ (96)
Plugging this into the Bellman equations for $(\pi_{-i},\mu_{i})$ and $\pi$ shows that the corresponding average rewards coincide:
$g^{(\pi_{-i},\mu_{i})}_{P,i}=g^{\pi}_{P,i},$
contradicting (90). Therefore no such improving deviation $\mu_{i}$ exists, and $\pi$ is an average-reward Nash equilibrium.
(ii) Contraction of the operator $T$ and convergence.
Let $h=(h_{1},\dots,h_{N})\in\mathbb{R}^{N|\mathcal{S}|}$ and define
$T_{i}(h)(s)\coloneqq\mathbb{E}_{\pi(s)}\bigl[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i})\bigr],\qquad T(h)\coloneqq(T_{1}(h),\dots,T_{N}(h)).$
We show that $T$ is a contraction in the span seminorm. Let $h,h^{\prime}$ be two vectors and fix a player $i$ and state $s$.
Case 1 (global optimality). If the NE selection at state $s$ is globally optimal in the sense of Assumption 3, then
$T_{i}(h)(s)-T_{i}(h^{\prime})(s)$ $=\max_{\pi(s)}\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i})]-\max_{\pi^{\prime}(s)}\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h^{\prime}_{i})]$
$\leq\mathbb{E}_{\pi(s)}\bigl[\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i})\bigr]\leq\max_{a}\bigl(\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i})\bigr).$ (97)
Case 2 (saddle point). If the NE is a saddle point, Assumption 3 gives, for any joint policy $\pi^{\prime}$ differing only in player $i$’s component,
$T_{i}(h)(s)-T_{i}(h^{\prime})(s)$ $=\mathbb{E}_{\pi(s)}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i})]-\mathbb{E}_{\pi^{\prime}(s)}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h^{\prime}_{i})]$
$\leq\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h_{i})]-\mathbb{E}_{(\pi^{\prime}_{-i}(s),\pi_{i}(s))}[r_{i}(s,a)+\sigma_{P^{a}_{s}}(h^{\prime}_{i})]$
$\leq\max_{a}\bigl(\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i})\bigr).$ (98)
By symmetry (reversing the roles of $h$ and $h^{\prime}$), we obtain the corresponding lower bound
$T_{i}(h)(s)-T_{i}(h^{\prime})(s)\;\geq\;\min_{a}\bigl(\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i})\bigr).$ (99)
Thus for every $s$,
$\min_{a}(\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i}))\;\leq\;T_{i}(h)(s)-T_{i}(h^{\prime})(s)\;\leq\;\max_{a}(\sigma_{P^{a}_{s}}(h_{i})-\sigma_{P^{a}_{s}}(h^{\prime}_{i})).$
By Lemma E.2 (applied to the support-function operator $\sigma_{P}$), this implies a span seminorm bound
$\operatorname{sp}(T_{i}(h)-T_{i}(h^{\prime}))\;\leq\;\operatorname{sp}(\sigma_{P}(h_{i})-\sigma_{P}(h^{\prime}_{i}))=\operatorname{sp}(P(h_{i}-h^{\prime}_{i})),$ (100)
for some irreducible transition matrix $P$ in the uncertainty set. Applying the standard span-contraction property of irreducible Markov chains ( puterman2014markov) and Lemma˜E.1, we get
$\operatorname{sp}(T_{i}(h)-T_{i}(h^{\prime}))\;\leq\;(1-\alpha(P))\,\operatorname{sp}(h_{i}-h^{\prime}_{i})\;\eqqcolon\;\gamma_{i}\,\operatorname{sp}(h_{i}-h^{\prime}_{i}),$ (101)
for some Dobrushin coefficient $\alpha(P)\in(0,1)$ and thus $\gamma_{i}<1$.
Taking the maximum over $i$ yields
$\max_{i}\operatorname{sp}(T_{i}(h)-T_{i}(h^{\prime}))\;\leq\;\gamma\,\max_{i}\operatorname{sp}(h_{i}-h^{\prime}_{i}),\quad\gamma\coloneqq\max_{i}\gamma_{i}<1.$
Hence $T$ is a contraction in the product span seminorm, and the Banach fixed-point theorem implies that the iterates $h^{k+1}=T(h^{k})$ converge (up to an additive constant vector for each player $i$) to a unique fixed point, and the associated policies converge to the unique NE policy $\pi$ described in part (i). This completes the proof. ∎
E.2 Technical Lemmas
Lemma E.1.
Define the ergodicity coefficient of a stochastic matrix $\mathsf{P}$ as
$\alpha(\mathsf{P})=1-\frac{1}{2}\max_{i,j}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|.$
Then for any stochastic matrix $\mathsf{P}$ and vector $V$, $\text{sp}(\mathsf{P}V)\leq(1-\alpha(\mathsf{P}))\cdot\text{sp}(V)$.
Proof.
For any two states $i,j$:
$(\mathsf{P}V)(i)-(\mathsf{P}V)(j)$ $=\sum_{k}(\mathsf{P}(k|i)-\mathsf{P}(k|j))V(k)$
$\leq\left(\frac{1}{2}\sum_{k}|\mathsf{P}(k|i)-\mathsf{P}(k|j)|\right)(\max_{s}V(s)-\min_{s}V(s))$
$\leq\left(\max_{i^{\prime},j^{\prime}}\frac{1}{2}\sum_{k}|\mathsf{P}(k|i^{\prime})-\mathsf{P}(k|j^{\prime})|\right)\text{sp}(V)$
$=(1-\alpha(\mathsf{P}))\text{sp}(V)$
Since this holds for any pair $(i,j)$, it must hold for the pair that defines the span of $\mathsf{P}V$. ∎
Lemma E.2.
For any $V_{1},V_{2}\in\mathbb{R}^{|\mathcal{S}|}$, there exists some transition kernel $\mathsf{P}^{\star}$ from $\mathcal{P}$ such that $\sigma^{\pi}(V_{1})-\sigma^{\pi}(V_{2})=\mathsf{P}^{\star}(V_{1}-V_{2})$.
Proof.
We first show this for a single state-action operator $\sigma_{\mathcal{P}_{s}^{a}}(V)=\min_{\mathsf{P}\in\mathcal{P}_{s}^{a}}\mathsf{P}V$. Let $\mathsf{P}_{1}=\arg\min_{\mathsf{P}}\mathsf{P}V_{1}$ and $\mathsf{P}_{2}=\arg\min_{\mathsf{P}}\mathsf{P}V_{2}$. From the optimality conditions, we have $\mathsf{P}_{1}V_{1}\leq\mathsf{P}_{2}V_{1}$ and $\mathsf{P}_{2}V_{2}\leq\mathsf{P}_{1}V_{2}$. This yields the bounds:
$\mathsf{P}_{1}(V_{1}-V_{2})\leq\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})\leq\mathsf{P}_{2}(V_{1}-V_{2})$
Since $\mathcal{P}_{s}^{a}$ is a convex set and the function $h(\mathsf{P})=\mathsf{P}(V_{1}-V_{2})$ is continuous in $\mathsf{P}$, by the Intermediate Value Theorem there must exist a $\mathsf{P}_{s,a}^{\star}\in\mathcal{P}_{s}^{a}$ such that $\sigma_{\mathcal{P}_{s}^{a}}(V_{1})-\sigma_{\mathcal{P}_{s}^{a}}(V_{2})=\mathsf{P}_{s,a}^{\star}(V_{1}-V_{2})$. We then construct the global kernel $\mathsf{P}^{\star}$ row by row as $\mathsf{P}^{\star}(s^{\prime}|s)=\sum_{a}\pi(a|s)\mathsf{P}_{s,a}^{\star}(s^{\prime}|s,a)$. This leads to the desired vector equality. ∎
E.3 Proximal Descent Algorithm
Fix a player $k\in\mathcal{N}$ and a vector $v\in\mathbb{R}^{|\mathcal{S}|}$. Define the robust one-step return
$F_{k}(s,a;v)\;\triangleq\;r_{k}(s,a)+\min_{q\in U(s,a)}q^{\top}v,\qquad s\in S,\;a\in A(s).$ (102)
For a stationary joint policy $\pi$, let $(g_{k}^{\pi},V_{k}^{\pi})$ be a (relative) robust gain–bias pair for player $k$, anchored at a fixed reference state $s^{\circ}$, solving the robust average-reward Bellman equation (per Theorem 1):
$V_{k}^{\pi}(s)=\sum_{a}\pi(a\mid s)\Bigl(r_{k}(s,a)-g_{k}^{\pi}+\min_{q\in U(s,a)}q^{\top}V_{k}^{\pi}\Bigr),\qquad s\in S.$ (103)
Given $\pi$ and $(g_{k}^{\pi},V_{k}^{\pi})$, define the robust one-step value for a unilateral deviation of player $k$ at state $s$:
$G_{k}\bigl(s,a_{k};\pi_{-k},V_{k}^{\pi}\bigr):=\sum_{a_{-k}}\pi_{-k}(a_{-k}\mid s)\,F_{k}\bigl(s,(a_{k},a_{-k});V_{k}^{\pi}\bigr).$ (104)
The robust temporal-difference error for action $a_{k}$ is
$\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr)\;\triangleq\;g_{k}^{\pi}+V_{k}^{\pi}(s)-G_{k}\bigl(s,a_{k};\pi_{-k},V_{k}^{\pi}\bigr).$ (105)
By plugging $a_{k}\sim\pi_{k}(\cdot\mid s)$ into the Bellman equation, we have, for all $k$ and $s$,
$\sum_{a_{k}}\pi_{k}(a_{k}\mid s)\,\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr)\;=\;0.$ (106)
Definition 3 (Per-state and global robust gaps).
For each player–state pair $(k,s)$ define
$G_{k,s}(\pi)$ $\;\triangleq\;\max_{a_{k}}G_{k}\bigl(s,a_{k};\pi_{-k},V_{k}^{\pi}\bigr)-\bigl(g_{k}^{\pi}+V_{k}^{\pi}(s)\bigr)$ (107)
$=-\,\min_{a_{k}}\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr),$ (108)
and the global gap
$G(\pi)\;\triangleq\;\sum_{k\in\mathcal{N}}\sum_{s\in S}G_{k,s}(\pi).$ (109)
For any $v\in\mathbb{R}^{|\mathcal{S}|}$ define
$F_{k}(s,a;v)\ \triangleq\ r_{k}(s,a)\ +\ \min_{q\in\mathcal{U}(s,a)}q^{\top}v.$ (110)
Fix an anchor $s_{\circ}\in S$; when needed we impose $v(s_{\circ})=0$.
For a fixed $\pi$, a pair $(g_{k}^{\pi},V_{k}^{\pi})$ (unique up to an additive constant fixed by $V_{k}^{\pi}(s_{\circ})=0$) solves the robust average‑reward Bellman equation
$V_{k}^{\pi}(s)\ =\ \sum_{a}\pi(a\mid s)\Big(r_{k}(s,a)-g_{k}^{\pi}+\min_{q\in\mathcal{U}(s,a)}q^{\top}V_{k}^{\pi}\Big),$ (111)
per Theorem˜4.1.
Lemma E.3 (Zero gap $\Leftrightarrow$ robust NE).
For any joint policy $\pi$:
  1. 1.
    For all $k,s$ we have $G_{k,s}(\pi)\geq 0$, hence $G(\pi)\geq 0$.
  2. 2.
    $G(\pi)=0$ if and only if for every $k,s$ the TD errors satisfy
    $\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr)\;\geq\;0\quad\text{for all }a_{k},\qquad\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr)=0\;\text{whenever }\pi_{k}(a_{k}\mid s)>0.$ (112)
  3. 3.
    $G(\pi)=0$ if and only if $\pi$ is a robust Nash equilibrium.
Proof.
(1) Fix $k,s$. By definition
$G_{k,s}(\pi)=-\min_{a_{k}}\Omega^{k}_{R}\bigl(s,a_{k};\pi_{-k}\mid V_{k}^{\pi},g_{k}^{\pi}\bigr).$
From (106), $\sum_{a_{k}}\pi_{k}(a_{k}\mid s)\,\Omega^{k}_{R}(s,a_{k};\cdot)=0$ is a convex combination of the $\Omega^{k}_{R}(s,a_{k};\cdot)$, so $\min_{a_{k}}\Omega^{k}_{R}(s,a_{k};\cdot)\leq 0$. Hence $G_{k,s}(\pi)\geq 0$, and summing over $(k,s)$ yields $G(\pi)\geq 0$.
(2) Since every $G_{k,s}(\pi)\geq 0$, $G(\pi)=0$ is equivalent to $G_{k,s}(\pi)=0$ for all $k,s$. For a fixed $(k,s)$ we have
$G_{k,s}(\pi)=0\quad\Longleftrightarrow\quad\min_{a_{k}}\Omega^{k}_{R}(s,a_{k};\cdot)=0.$
Thus $\Omega^{k}_{R}(s,a_{k};\cdot)\geq 0$ for all $a_{k}$, and there exists at least one action $a_{k}$ with $\Omega^{k}_{R}(s,a_{k};\cdot)=0$.
Now combine this with (106). Because $\Omega^{k}_{R}(s,a_{k};\cdot)\geq 0$ and $\sum_{a_{k}}\pi_{k}(a_{k}\mid s)\Omega^{k}_{R}(s,a_{k};\cdot)=0$, a nonnegative convex combination can vanish only if each term supported by $\pi_{k}(\cdot\mid s)$ is zero. Hence
$\pi_{k}(a_{k}\mid s)>0\quad\Longrightarrow\quad\Omega^{k}_{R}(s,a_{k};\cdot)=0.$
Conversely, if (112) holds, then $\min_{a_{k}}\Omega^{k}_{R}(s,a_{k};\cdot)=0$, so $G_{k,s}(\pi)=0$ and thus $G(\pi)=0$.
(3) $G(\pi)=0\Rightarrow\pi$ robust NE. Assume $G(\pi)=0$. By item (2), for each $k$ and $s$ the policy $\pi_{k}(\cdot\mid s)$ is statewise greedy with respect to $(g_{k}^{\pi},V_{k}^{\pi})$ for the induced single-agent robust MDP where player $k$ controls its own action and the other players follow $\pi_{-k}$. More precisely, the condition $\Omega^{k}_{R}(s,a_{k};\cdot)\geq 0$ for all $a_{k}$ and equality on the support of $\pi_{k}(\cdot\mid s)$ is exactly the Bellman optimality slackness condition for the constant-gain equation in that induced DR-MDP (per Lemma˜D.3). Therefore, $\pi_{k}$ is a robust best response to $\pi_{-k}$ for every $k$, so $\pi$ is a robust Nash equilibrium.
(3) $\pi$ robust NE $\Rightarrow G(\pi)=0$. Conversely, suppose $\pi$ is a robust Nash equilibrium. Then for each $k$, $\pi_{k}$ is an optimal stationary policy for the induced single-agent robust average-reward MDP with transition uncertainty rectangular in $(s,a)$ and reward $r_{k}$. By Theorem˜4.1, there exists a gain–bias pair $(\tilde{g}_{k},\tilde{V}_{k})$ such that
$\tilde{g}_{k}+\tilde{V}_{k}(s)=\max_{a_{k}}G_{k}\bigl(s,a_{k};\pi_{-k},\tilde{V}_{k}\bigr),\qquad s\in S,$ (113)
and the associated TD errors satisfy $\Omega^{k}_{R}(s,a_{k};\pi_{-k}\mid\tilde{V}_{k},\tilde{g}_{k})\geq 0$ with equality on the support of $\pi_{k}(\cdot\mid s)$. Since robust gain–bias pairs are unique up to an additive constant in the bias (and we fix $V_{k}^{\pi}(s^{\circ})=0$), we may identify $(g_{k}^{\pi},V_{k}^{\pi})$ with $(\tilde{g}_{k},\tilde{V}_{k})$. Hence (112) holds, and by item (2) we obtain $G(\pi)=0$. ∎
Hence, finding a robust NE is equivalent to solve $G(\pi)=0$, or equivalently, $G_{k,s}(\pi)=0,\forall k,s$.
E.3.1 Algorithm design
At iterate $\pi^{n}$, compute $(g_{k}^{n},V_{k}^{n})=(g_{k}^{\pi^{n}},V_{k}^{\pi^{n}})$. Define the frozen gap
$\widehat{G}_{n}(\pi)\ \triangleq\ \sum_{k,s}\Big(\max_{a_{k}}\sum_{a_{-k}}\pi_{-k}(a_{-k}\mid s)\,F_{k}\big(s,(a_{k},a_{-k});V_{k}^{n}\big)\ -\ (g_{k}^{n}+V_{k}^{n}(s))\Big).$ (114)
For fixed $(n,k,s)$, the map $\mu\mapsto\widehat{G}_{n}(\pi_{-(k,s)},\mu)$ is a pointwise maximum of affine forms in $\mu\in\Delta(\mathcal{A}_{k})$, hence convex and piecewise linear.
The hybrid (spliced) policy used inside the current block’s subproblem.
At epoch $n$, when we are updating the block $(k,s)$, all blocks before $(k,s)$ have already been updated to their new values $\pi^{n+1}$; the blocks after $(k,s)$ still retain their old values $\pi^{n}$. For the current block we use a candidate vector $\mu\in\Delta(\mathcal{A}_{k})$.
We encode this by the splice (hybrid) policy $\Pi^{(n)}_{k,s}(\mu)\in\Pi\triangleq\times_{(i,t)}\Delta(A^{i}(t))$:
$\big[\Pi^{(n)}_{k,s}(\mu)\big]_{i,t}\;\triangleq\;\begin{cases}\pi^{n+1}_{i}(\cdot\mid t),&\text{if }(i,t)\in\mathcal{B}_{<k,s>},\\[2.0pt] \mu,&\text{if }(i,t)=(k,s),\\[2.0pt] \pi^{n}_{i}(\cdot\mid t),&\text{if }(i,t)\in\mathcal{B}_{>k,s>}.\end{cases}$ (115)
Definition 4.
For any function $F:\Pi\to\mathbb{R}$, and in particular for the frozen gap $\widehat{G}_{n}(\cdot)$, we use the shorthand
$\widehat{G}_{n}\big(\pi^{n+1}_{<k,s>},\,\mu,\,\pi^{n}_{>k,s>}\big)\;\triangleq\;\widehat{G}_{n}\!\big(\,\Pi^{(n)}_{k,s}(\mu)\,\big).$
Recall the frozen gap (evaluation frozen at $(g^{n},V^{n})$):
$\widehat{G}_{n}(\pi)=\sum_{i=1}^{K}\sum_{t\in S}\left(\max_{a_{i}}\ \sum_{a_{-i}}\Big[\times_{j\neq i}\pi_{j}(a_{j}\mid t)\Big]\,F_{i}\big(t,(a_{i},a_{-i});V_{i}^{n}\big)\;-\;(g_{i}^{n}+V_{i}^{n}(t))\right).$ (116)
When $\pi$ is replaced by the splice $\Pi^{(n)}_{k,s}(\mu)$, the block $\mu=\pi_{k}(\cdot\mid s)$ appears:
  • only in the terms with state $t=s$ and player $i\neq k$ (because the product $\times_{j\neq i}\pi_{j}(\cdot\mid t)$ includes $\pi_{k}(\cdot\mid s)$ exactly when $t=s$ and $i\neq k$);
  • not in the terms with $i=k$ (there the product is over $j\neq k$) nor in the terms with $t\neq s$.
We recall the block structure used in Algorithm 2. Let
$\mathcal{B}\;\triangleq\;\mathcal{N}\times S$
be the set of player–state blocks. For $b=(k,s)\in\mathcal{B}$, the associated block variable is the simplex vector $\pi_{k}(\cdot\mid s)\in\Delta(A_{k})$. Fix a total order $\prec$ on $\mathcal{B}$ (e.g., lexicographic).
For a given block $b=(k,s)$, denote the set of earlier and later blocks by
$\mathcal{B}_{<k,s>}:=\{(i,t)\in\mathcal{B}:(i,t)\prec(k,s)\},\qquad\mathcal{B}_{>k,s>}:=\{(i,t)\in\mathcal{B}:(i,t)\succ(k,s)\}.$
At epoch $n$, we evaluate the current policy $\pi^{n}$ to obtain $(g_{k}^{n},V_{k}^{n})_{k\in\mathcal{N}}$. We then freeze these evaluations and define the frozen gap
$G^{b}_{n}(\pi)\;\triangleq\;\sum_{k,s}\widetilde{G}_{k,s}\bigl(\pi;g^{n},V^{n}\bigr),$ (117)
where
$\widetilde{G}_{k,s}(\pi;g,V):=\max_{a_{k}}\sum_{a_{-k}}\pi_{-k}(a_{-k}\mid s)\,F_{k}\bigl(s,(a_{k},a_{-k});V_{k}\bigr)-(g_{k}+V_{k}(s)).$ (118)
Thus $G^{b}_{n}(\cdot)$ is the same as $G(\cdot)$ but with the evaluation pair $(g_{k}^{\pi},V_{k}^{\pi})$ frozen at $(g_{k}^{n},V_{k}^{n})$.
For a block $(k,s)$ and a candidate replacement $\mu\in\Delta(A_{k})$ for the local policy at $(k,s)$, define the hybrid policy
$\Pi^{(n)}_{k,s}(\mu)\;\triangleq\;\bigl(\pi^{n+1}_{<k,s>},\,\mu,\,\pi^{n}_{>k,s>}\bigr),$
i.e., we use the already-updated blocks for earlier coordinates, the new block $\mu$ at $(k,s)$, and the old iterate $\pi^{n}$ for all later blocks. Then the blockwise objective used is the optimistic (frozen) gap
$\mu\;\longmapsto\;G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr).$
Lemma E.4.
Fix $n$ and a block $b=(k,s)$. Let
$\phi_{n,k,s}(\mu)\;\triangleq\;G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr),\qquad\mu\in\Delta(A_{k}).$
Then:
  1. 1.
    $\phi_{n,k,s}(\mu)$ depends on $\mu$ only through the terms $\widetilde{G}_{i,s}(\cdot)$ with $i\neq k$ and the fixed state $s$; all other summands are constant in $\mu$.
  2. 2.
    For each such $(i,s)$, the dependence on $\mu$ is a finite maximum of affine functions of $\mu$. Consequently, $\phi_{n,k,s}(\mu)$ is the sum of a constant and finitely many pointwise maxima of affine forms in $\mu$.
  3. 3.
    In particular, for every $k,s$, the map $\mu\mapsto\phi_{n,k,s}(\mu)$ is convex (indeed piecewise-linear) on $\Delta(A_{k})$.
Proof.
By definition
$G^{b}_{n}(\pi)=\sum_{i\in\mathcal{N}}\sum_{t\in S}\widetilde{G}_{i,t}\bigl(\pi;g^{n},V^{n}\bigr).$
When we form $\pi=\Pi^{(n)}_{k,s}(\mu)$, the only coordinates where $\pi$ depends on $\mu$ are the joint-action probabilities at state $s$ in which player $k$ acts. Thus:
  • For any $t\neq s$, the summand $\widetilde{G}_{i,t}$ is independent of $\mu$.
  • For $t=s$ and $i=k$, $\widetilde{G}_{k,s}$ depends on $\pi_{-k}$ but not on $\pi_{k}(\cdot\mid s)$ inside the max, so changing $\mu$ does not affect $\widetilde{G}_{k,s}$ either.
  • For $t=s$ and $i\neq k$, $\widetilde{G}_{i,s}$ depends on $\pi_{-i}$ through the joint action distribution at state $s$, which in turn depends linearly on $\mu$ because only the $k$-th factor changes.
Hence item (1) holds.
For item (2), fix $i\neq k$ and state $s$. The term $\widetilde{G}_{i,s}$ has the form
$\widetilde{G}_{i,s}(\pi;g^{n},V^{n})=\max_{a_{i}}\sum_{a_{-i}}\pi_{-i}(a_{-i}\mid s)F_{i}\bigl(s,(a_{i},a_{-i});V_{i}^{n}\bigr)-(g_{i}^{n}+V_{i}^{n}(s)).$
For a fixed $a_{i}$, the coefficient of $\mu$ is
$\sum_{a_{-i}}\Bigl(\times_{\ell\neq i}\pi_{\ell}(a_{\ell}\mid s)\Bigr)F_{i}(s,(a_{i},a_{-i});V_{i}^{n}),$
and only the factor $\pi_{k}(\cdot\mid s)$ in this product is replaced by $\mu$. Thus the dependence on $\mu$ is linear for each $a_{i}$, and the outer max over $a_{i}$ produces a pointwise maximum of finitely many affine forms. Summing over all $(i,s)$ and adding the (constant) terms independent of $\mu$ gives the claimed representation for $\phi_{n,k,s}$. A max of affine functions is convex; therefore $\phi_{n,k,s}$ is convex in $\mu$, establishing items (2) and (3). ∎
Proposition E.1.
For any epoch $n$ and block $(k,s)$, the block update
$\pi^{n+1}_{k}(\cdot\mid s)\in\arg\min_{\mu\in\Delta(A_{k})}\Bigl\{\phi_{n,k,s}(\mu)+\frac{1}{2\eta_{n}}\bigl\|\mu-\pi^{n}_{k}(\cdot\mid s)\bigr\|_{2}^{2}\Bigr\}$
has a unique solution. Moreover, letting $\mu^{\star}$ denote this minimizer, we have
$G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu^{\star})\bigr)+\frac{1}{2\eta_{n}}\bigl\|\mu^{\star}-\pi^{n}_{k}(\cdot\mid s)\bigr\|_{2}^{2}\;\leq\;G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\pi^{n}_{k}(\cdot\mid s))\bigr),$
i.e., the block objective at the minimizer is no larger than at the feasible point $\mu=\pi^{n}_{k}(\cdot\mid s)$.
Proof.
By Lemma E.4, $\phi_{n,k,s}$ is convex in $\mu$; adding the quadratic term makes the block objective $(1/\eta_{n})$-strongly convex on the compact convex set $\Delta(\mathcal{A}_{k})$, hence it has a unique minimizer. The displayed inequality is the minimality condition obtained by comparing the minimizer against the feasible competitor $\mu=\pi^{n}_{k}(\cdot\mid s)$. ∎
E.3.2 Convergence analysis
Let
$\Delta_{n}\;\triangleq\;\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}$
denote the per-epoch policy displacement. We first establish a sufficient decrease property for the frozen gap $G^{b}_{n}$, then control the error made by re-evaluating at $\pi^{n+1}$, and finally combine them into a quasi-descent inequality for the true gap $G$.
Lemma E.5.
For any $n$ and $\eta_{n}>0$,
$G^{b}_{n}(\pi^{n+1})\;\leq\;G^{b}_{n}(\pi^{n})-\frac{1}{2\eta_{n}}\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}^{2}.$
Proof.
Index the blocks in the chosen order by $b=1,\dots,B$ and write $\Pi^{(0)}\equiv\pi^{n}$. After finishing block $b$ we denote by $\Pi^{(b)}$ the partially updated policy (so $\Pi^{(B)}=\pi^{n+1}$).
At block $b$, Algorithm solves
$\min_{\mu\in\Delta}\Bigl\{\Phi_{b}(\mu):=G^{b}_{n}\bigl(\Pi^{(b-1)}_{\text{others}},\mu\bigr)+\frac{1}{2\eta_{n}}\bigl\|\mu-\Pi^{(0)}_{b}\bigr\|_{2}^{2}\Bigr\},$
where the dependence on $\mu$ is convex by Lemma˜E.4. By optimality of the chosen update $\Pi^{(b)}_{b}$ and feasibility of $\mu=\Pi^{(b-1)}_{b}$ we have
$G^{b}_{n}(\Pi^{(b)})+\frac{1}{2\eta_{n}}\bigl\|\Pi^{(b)}_{b}-\Pi^{(0)}_{b}\bigr\|_{2}^{2}\;\leq\;G^{b}_{n}(\Pi^{(b-1)})+\frac{1}{2\eta_{n}}\bigl\|\Pi^{(b-1)}_{b}-\Pi^{(0)}_{b}\bigr\|_{2}^{2}.$
Subtract the quadratic term on the right and sum these inequalities over $b=1,\dots,B$. The quadratic terms telescope:
$G^{b}_{n}(\Pi^{(B)})\;\leq\;G^{b}_{n}(\Pi^{(0)})-\frac{1}{2\eta_{n}}\sum_{b=1}^{B}\bigl\|\Pi^{(b)}_{b}-\Pi^{(b-1)}_{b}\bigr\|_{2}^{2}.$
Because the blocks are disjoint groups of coordinates,
$\sum_{b=1}^{B}\bigl\|\Pi^{(b)}_{b}-\Pi^{(b-1)}_{b}\bigr\|_{2}^{2}=\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}^{2}.$
Plugging this in yields the claimed inequality. ∎
Lemma E.6.
There exists a constant $C_{\mathrm{ev}}>0$, depending only on the problem bounds, such that for all $n$,
$\bigl|G(\pi^{n+1})-G^{b}_{n}(\pi^{n+1})\bigr|\;\leq\;C_{\mathrm{ev}}\,\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}.$
Proof.
Write $g^{n+1}_{k}:=g_{k}^{\pi^{n+1}}$ and $V^{n+1}_{k}:=V_{k}^{\pi^{n+1}}$. The difference between the true and frozen gaps at $\pi^{n+1}$ is
$G(\pi^{n+1})$ $-G^{b}_{n}(\pi^{n+1})$
$=$ $\sum_{k,s}\Bigl[\max_{a_{k}}\sum_{a_{-k}}\pi^{n+1}_{-k}(a_{-k}\mid s)F_{k}\bigl(s,(a_{k},a_{-k});V^{n+1}_{k}\bigr)-\bigl(g^{n+1}_{k}+V^{n+1}_{k}(s)\bigr)\Bigr]$
$\quad-\sum_{k,s}\Bigl[\max_{a_{k}}\sum_{a_{-k}}\pi^{n+1}_{-k}(a_{-k}\mid s)F_{k}\bigl(s,(a_{k},a_{-k});V^{n}_{k}\bigr)-\bigl(g^{n}_{k}+V^{n}_{k}(s)\bigr)\Bigr].$
Rearranging gives
$\bigl|G(\pi^{n+1})-G^{b}_{n}(\pi^{n+1})\bigr|\;\leq\;$ $\sum_{k,s}\biggl|\max_{a_{k}}\sum_{a_{-k}}\pi^{n+1}_{-k}(a_{-k}\mid s)\bigl[F_{k}\bigl(s,(a_{k},a_{-k});V^{n+1}_{k}\bigr)-F_{k}\bigl(s,(a_{k},a_{-k});V^{n}_{k}\bigr)\bigr]\biggr|$
$\quad+\sum_{k,s}\bigl|g^{n+1}_{k}-g^{n}_{k}\bigr|+\sum_{k,s}\bigl|V^{n+1}_{k}(s)-V^{n}_{k}(s)\bigr|.$
For fixed $(k,s,a_{k},a_{-k})$, the map $v\mapsto F_{k}(s,(a_{k},a_{-k});v)$ has the form $r_{k}(s,(a_{k},a_{-k}))+\min_{q\in U(s,(a_{k},a_{-k}))}q^{\top}v$ with $q$ in the probability simplex, hence it is 1-Lipschitz in the $\|\cdot\|_{\infty}$ norm:
$\bigl|F_{k}(s,(a_{k},a_{-k});v)-F_{k}(s,(a_{k},a_{-k});v^{\prime})\bigr|\leq\bigl\|v-v^{\prime}\bigr\|_{\infty}.$
Therefore each summand inside the max is bounded in absolute value by $\|V^{n+1}_{k}-V^{n}_{k}\|_{\infty}$, and so the max is bounded by the same quantity. Summing over $s$ and $k$ yields
$\sum_{k,s}\biggl|\max_{a_{k}}\sum_{a_{-k}}\pi^{n+1}_{-k}(a_{-k}\mid s)\bigl[F_{k}(\cdot;V^{n+1}_{k})-F_{k}(\cdot;V^{n}_{k})\bigr]\biggr|\;\leq\;|\mathcal{S}|\sum_{k}\bigl\|V^{n+1}_{k}-V^{n}_{k}\bigr\|_{\infty}.$
For the scalar and bias terms we simply bound
$\sum_{k,s}\bigl|V^{n+1}_{k}(s)-V^{n}_{k}(s)\bigr|\leq|\mathcal{S}|\sum_{k}\bigl\|V^{n+1}_{k}-V^{n}_{k}\bigr\|_{\infty},\qquad\sum_{k,s}\bigl|g^{n+1}_{k}-g^{n}_{k}\bigr|\leq|\mathcal{S}|\sum_{k}\bigl|g^{n+1}_{k}-g^{n}_{k}\bigr|.$
By the Lipschitz continuity of the evaluation map $\pi\mapsto(g^{\pi}_{k},V^{\pi}_{k})$ (see Section˜E.3.3), there exist constants $L_{V},L_{g}>0$ such that for all policies $\pi,\tilde{\pi}$,
$\bigl\|V^{\pi}_{k}-V^{\tilde{\pi}}_{k}\bigr\|_{\infty}\leq L_{V}\bigl\|\pi-\tilde{\pi}\bigr\|_{2},\qquad\bigl|g^{\pi}_{k}-g^{\tilde{\pi}}_{k}\bigr|\leq L_{g}\bigl\|\pi-\tilde{\pi}\bigr\|_{2}.$
Applying this with $(\pi,\tilde{\pi})=(\pi^{n+1},\pi^{n})$ and combining the bounds yields
$\bigl|G(\pi^{n+1})-G^{b}_{n}(\pi^{n+1})\bigr|\leq C_{\mathrm{ev}}\,\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}$
for some constant $C_{\mathrm{ev}}$ depending only on the problem data ($|\mathcal{S}|$, number of players, reward bounds, and the Lipschitz constants). ∎
Proposition E.2.
There exists a constant $C_{\mathrm{ev}}>0$ such that, for all $n$,
$G(\pi^{n+1})\;\leq\;G(\pi^{n})-\frac{1}{4\eta_{n}}\bigl\|\pi^{n+1}-\pi^{n}\bigr\|_{2}^{2}+C_{\mathrm{ev}}^{2}\eta_{n}.$ (119)
As a consequence, if $\sum_{n}\eta_{n}<\infty$, then:
  1. 1.
    The sequence $\{G(\pi^{n})\}$ converges to a finite limit.
  2. 2.
    The step sizes satisfy
    $\sum_{n=0}^{\infty}\frac{\|\pi^{n+1}-\pi^{n}\|_{2}^{2}}{\eta_{n}}<\infty,\qquad\|\pi^{n+1}-\pi^{n}\|_{2}\;\to\;0.$
Proof.
By Lemma E.5 and Lemma E.6,
$G(\pi^{n+1})$ $\leq G^{b}_{n}(\pi^{n+1})+C_{\mathrm{ev}}\|\pi^{n+1}-\pi^{n}\|_{2}$
$\leq G^{b}_{n}(\pi^{n})-\frac{1}{2\eta_{n}}\|\pi^{n+1}-\pi^{n}\|_{2}^{2}+C_{\mathrm{ev}}\|\pi^{n+1}-\pi^{n}\|_{2}$
$=G(\pi^{n})-\frac{1}{2\eta_{n}}\|\pi^{n+1}-\pi^{n}\|_{2}^{2}+C_{\mathrm{ev}}\|\pi^{n+1}-\pi^{n}\|_{2},$
since $G^{b}_{n}(\pi^{n})=G(\pi^{n})$ by construction. Apply Young’s inequality with parameter $2\eta_{n}$:
$C_{\mathrm{ev}}\|\pi^{n+1}-\pi^{n}\|_{2}\leq\frac{1}{4\eta_{n}}\|\pi^{n+1}-\pi^{n}\|_{2}^{2}+C_{\mathrm{ev}}^{2}\eta_{n},$
to obtain
$G(\pi^{n+1})\leq G(\pi^{n})-\frac{1}{4\eta_{n}}\|\pi^{n+1}-\pi^{n}\|_{2}^{2}+C_{\mathrm{ev}}^{2}\eta_{n},$
which is (119).
Now apply the deterministic Robbins–Siegmund lemma (Lemma˜E.17) to the real sequence $x_{n}:=G(\pi^{n})$, with $y_{n}:=\frac{1}{4\eta_{n}}\|\pi^{n+1}-\pi^{n}\|_{2}^{2}$ and $z_{n}:=C_{\mathrm{ev}}^{2}\eta_{n}$. Since $G(\pi^{n})\geq 0$ and $\sum_{n}z_{n}=C_{\mathrm{ev}}^{2}\sum_{n}\eta_{n}<\infty$, the lemma implies that $\{G(\pi^{n})\}$ converges and $\sum_{n}y_{n}<\infty$, i.e.,
$\sum_{n=0}^{\infty}\frac{\|\pi^{n+1}-\pi^{n}\|_{2}^{2}}{\eta_{n}}<\infty.$
Finally, because $\eta_{n}>0$ and $\sum_{n}\eta_{n}<\infty$ forces $\eta_{n}\to 0$, the above square-summability implies $\|\pi^{n+1}-\pi^{n}\|_{2}\to 0$. (Otherwise we could extract a subsequence where $\|\pi^{n+1}-\pi^{n}\|_{2}\geq\varepsilon>0$, which would make $\sum\|\pi^{n+1}-\pi^{n}\|_{2}^{2}/\eta_{n}$ diverge.) ∎
We now derive approximate optimality conditions for each block, which will eventually lead to Clarke stationarity of cluster points for the true gap.
For convenience, collect all block variables into a vector $\pi\in\Pi:=\times_{(k,s)}\Delta(A_{k})$. Let $\|\cdot\|_{2}$ be the Euclidean norm in this product space.
Lemma E.7.
For each epoch $n$ there exist vectors
$\xi^{\,n+1}_{k,s}\in\mathbb{R}^{|\mathcal{A}_{k}|},\qquad(k,s)\in\mathcal{B},$
and normal vectors
$y^{\,n+1}_{k,s}\in N_{\Delta(A_{k})}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr)$
such that:
  1. 1.
    For every block $(k,s)$ and every $\mu\in\Delta(A_{k})$,
    $\bigl\langle\xi^{\,n+1}_{k,s},\mu-\pi^{n+1}_{k}(\cdot\mid s)\bigr\rangle\;\geq\;-\frac{1}{2\eta_{n}}\Bigl(\bigl\|\mu-\pi^{n}_{k}(\cdot\mid s)\bigr\|_{2}^{2}-\bigl\|\pi^{n+1}_{k}(\cdot\mid s)-\pi^{n}_{k}(\cdot\mid s)\bigr\|_{2}^{2}\Bigr).$ (120)
  2. 2.
    Let $\xi^{\,n+1}$ denote the concatenation of all $\xi^{\,n+1}_{k,s}$, and similarly for $\pi^{n}$ and $\mu\in\Pi$. Then for all $\mu\in\Pi$,
    $\bigl\langle\xi^{\,n+1},\,\mu-\pi^{n+1}\bigr\rangle\;\geq\;-\frac{1}{2\eta_{n}}\Bigl(\|\mu-\pi^{n}\|_{2}^{2}-\|\pi^{n+1}-\pi^{n}\|_{2}^{2}\Bigr).$ (121)
  3. 3.
    For every direction $d$ in the tangent cone $T_{\Pi}(\pi^{n+1})$,
    $\bigl\langle\xi^{\,n+1},\,d\bigr\rangle+\frac{1}{\eta_{n}}\bigl\langle\pi^{n+1}-\pi^{n},\,d\bigr\rangle\;\geq\;0.$ (122)
Moreover, each $\xi^{\,n+1}_{k,s}$ belongs to the Clarke subdifferential of the blockwise frozen gap:
$\xi^{\,n+1}_{k,s}\in\partial_{C}\!\Bigl[\mu\mapsto G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr)\Bigr]\Big|_{\mu=\pi^{n+1}_{k}(\cdot\mid s)}.$
Proof.
Fix $(k,s)$ and define the convex block objective
$\varphi_{n,(k,s)}(\mu):=G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr)+\frac{1}{2\eta_{n}}\bigl\|\mu-\pi^{n}_{k}(\cdot\mid s)\bigr\|_{2}^{2},\qquad\mu\in\Delta(A_{k}).$
By Proposition˜E.1, $\pi^{n+1}_{k}(\cdot\mid s)$ is the unique minimizer of $\varphi_{n,(k,s)}$. The first-order optimality condition on the simplex reads
$0\in\partial\varphi_{n,(k,s)}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr)+N_{\Delta(A_{k})}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr).$
Expanding the (Clarke) subgradient of $\varphi_{n,(k,s)}$ at $\pi^{n+1}_{k}(\cdot\mid s)$ gives
$\partial\varphi_{n,(k,s)}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr)=\partial_{\mu}G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr)\Big|_{\mu=\pi^{n+1}_{k}(\cdot\mid s)}+\frac{1}{\eta_{n}}\bigl(\pi^{n+1}_{k}(\cdot\mid s)-\pi^{n}_{k}(\cdot\mid s)\bigr).$
Thus we can choose
$\widetilde{\xi}^{\,n+1}_{k,s}\in\partial_{\mu}G^{b}_{n}\bigl(\Pi^{(n)}_{k,s}(\mu)\bigr)\Big|_{\mu=\pi^{n+1}_{k}(\cdot\mid s)},\qquad y^{\,n+1}_{k,s}\in N_{\Delta(A_{k})}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr)$
such that
$\widetilde{\xi}^{\,n+1}_{k,s}+\frac{1}{\eta_{n}}\bigl(\pi^{n+1}_{k}-\pi^{n}_{k}\bigr)(\cdot\mid s)+y^{\,n+1}_{k,s}=0.$
For any $\mu\in\Delta(A_{k})$, $\mu-\pi^{n+1}_{k}(\cdot\mid s)$ lies in the tangent cone $T_{\Delta(A_{k})}(\pi^{n+1}_{k}(\cdot\mid s))$, and by normal–tangent polarity we have $\langle y^{\,n+1}_{k,s},\mu-\pi^{n+1}_{k}(\cdot\mid s)\rangle\leq 0$. Thus
$\bigl\langle\widetilde{\xi}^{\,n+1}_{k,s},\mu-\pi^{n+1}_{k}(\cdot\mid s)\bigr\rangle+\frac{1}{\eta_{n}}\bigl\langle\pi^{n+1}_{k}(\cdot\mid s)-\pi^{n}_{k}(\cdot\mid s),\,\mu-\pi^{n+1}_{k}(\cdot\mid s)\bigr\rangle\;\geq\;0.$
Expanding the second inner product using
$\bigl\langle a-b,c-b\bigr\rangle=\tfrac{1}{2}\Bigl(\|c-a\|_{2}^{2}-\|b-a\|_{2}^{2}-\|c-b\|_{2}^{2}\Bigr)$
with $a=\pi^{n}_{k}(\cdot\mid s)$, $b=\pi^{n+1}_{k}(\cdot\mid s)$, $c=\mu$, we obtain
$\frac{1}{\eta_{n}}\bigl\langle\pi^{n+1}_{k}-\pi^{n}_{k},\mu-\pi^{n+1}_{k}\bigr\rangle=\frac{1}{2\eta_{n}}\Bigl(\|\mu-\pi^{n}_{k}\|_{2}^{2}-\|\pi^{n+1}_{k}-\pi^{n}_{k}\|_{2}^{2}-\|\mu-\pi^{n+1}_{k}\|_{2}^{2}\Bigr).$
Discarding the nonnegative term $\frac{1}{2\eta_{n}}\|\mu-\pi^{n+1}_{k}\|_{2}^{2}$ yields
$\bigl\langle\widetilde{\xi}^{\,n+1}_{k,s},\mu-\pi^{n+1}_{k}(\cdot\mid s)\bigr\rangle\geq-\frac{1}{2\eta_{n}}\Bigl(\|\mu-\pi^{n}_{k}\|_{2}^{2}-\|\pi^{n+1}_{k}-\pi^{n}_{k}\|_{2}^{2}\Bigr),$
which is (120) with $\xi^{\,n+1}_{k,s}:=\widetilde{\xi}^{\,n+1}_{k,s}$. Summing over all blocks leads to (121).
For item (3), take a direction $d\in T_{\Pi}(\pi^{n+1})$ and consider $\mu=\pi^{n+1}+td$ for small $t>0$ such that $\mu\in\Pi$. Dividing (121) by $t$ and letting $t\downarrow 0$ gives (122). The subgradient membership is exactly the definition of $\xi^{\,n+1}_{k,s}$ as a Clarke subgradient of the blockwise frozen gap at $\pi^{n+1}_{k}(\cdot\mid s)$. ∎
Define the scaled displacement (a “proximal residual”)
$\lambda_{n}\;\triangleq\;\frac{\pi^{n+1}-\pi^{n}}{\eta_{n}}.$
Lemma E.8.
There exists a constant $L_{\mathrm{dir}}>0$ depending only on the problem bounds such that
$\|\lambda_{n}\|_{2}\;\leq\;L_{\mathrm{dir}}\quad\text{for all }n.$
Proof.
From (122), for any $d\in T_{\Pi}(\pi^{n+1})$,
$G^{b}_{n}{}^{\circ}(\pi^{n+1};d)\;\geq\;-\bigl\langle\lambda_{n},d\bigr\rangle,$
where $G^{b}_{n}{}^{\circ}$ denotes the Clarke directional derivative. Take $d=-\Delta_{n}:=\pi^{n}-\pi^{n+1}$. Because $\Pi$ is convex, $d\in T_{\Pi}(\pi^{n+1})$, and thus
$G^{b}_{n}{}^{\circ}(\pi^{n+1};-\Delta_{n})\;\geq\;\langle\lambda_{n},\Delta_{n}\rangle=\frac{\|\Delta_{n}\|_{2}^{2}}{\eta_{n}}.$
On the other hand, $G^{b}_{n}$ is globally Lipschitz as a finite sum of max-of-affine functions with uniformly bounded slopes, so there exists $L_{\mathrm{dir}}>0$ such that
$\bigl|G^{b}_{n}{}^{\circ}(\pi;h)\bigr|\leq L_{\mathrm{dir}}\|h\|_{2}\qquad\text{for all }\pi,h.$
Applying this with $\pi=\pi^{n+1}$ and $h=-\Delta_{n}$ yields
$\frac{\|\Delta_{n}\|_{2}^{2}}{\eta_{n}}\leq L_{\mathrm{dir}}\|\Delta_{n}\|_{2}.$
If $\Delta_{n}\neq 0$, divide by $\|\Delta_{n}\|_{2}$ to obtain $\|\lambda_{n}\|_{2}\leq L_{\mathrm{dir}}$; otherwise the claim holds trivially. ∎
We next pass to cluster points and show Clarke stationarity.
Lemma E.9.
There exists a subsequence (not relabeled) such that
$\pi^{n+1}\to\pi^{\star},\quad\xi^{\,n+1}\to\xi,\quad\lambda_{n}\to\lambda,$
and the blockwise normal vectors $y^{\,n+1}$ from Lemma E.7 satisfy $y^{\,n+1}\to y\in N_{\Pi}(\pi^{\star})$. Moreover, $\xi\in\partial_{C}G(\pi^{\star})$ and
$\xi+\lambda+y=0.$ (123)
Proof.
Compactness of $\Pi$ implies the existence of a convergent subsequence $\pi^{n+1}\to\pi^{\star}$. By Lemma E.8 and boundedness of the slopes of $G^{b}_{n}$, the sequences $\{\lambda_{n}\}$ and $\{\xi^{\,n+1}\}$ are bounded, so (along a subsequence) $\lambda_{n}\to\lambda$ and $\xi^{\,n+1}\to\xi$.
At the block level, the KKT relation from Lemma E.7 reads
$\xi^{\,n+1}_{k,s}+\lambda_{n,k,s}+y^{\,n+1}_{k,s}=0,\quad y^{\,n+1}_{k,s}\in N_{\Delta(A_{k})}\bigl(\pi^{n+1}_{k}(\cdot\mid s)\bigr).$
Summing over $(k,s)$ gives
$\xi^{\,n+1}+\lambda_{n}+y^{\,n+1}=0,\qquad y^{\,n+1}\in N_{\Pi}(\pi^{n+1}).$
Since $\Pi$ is a polyhedron, its normal cone mapping has a closed graph (rockafellar1998variational), so any limit point $y$ of $\{y^{\,n+1}\}$ lies in $N_{\Pi}(\pi^{\star})$. Passing to the limit in the KKT identity yields (123).
It remains to show $\xi\in\partial_{C}G(\pi^{\star})$. Fix a direction $d$ and any sequence $d_{n}\to d$. For each $n$, $\xi^{\,n+1}\in\partial_{C}G^{b}_{n}(\pi^{n+1})$ and $G^{b}_{n}$ is a finite max of affine functions with coefficients that converge to those of $G$ (because $(g^{n},V^{n})\to(g^{\pi^{\star}},V^{\pi^{\star}})$ by the Lipschitz evaluation from Section˜E.3.3). Thus
$\langle\xi^{\,n+1},d_{n}\rangle\leq G^{b}_{n}{}^{\circ}(\pi^{n+1};d_{n}).$
By the directional upper semicontinuity of Clarke directional derivatives for max–affine families (Lemma˜E.18), the right-hand side satisfies
$\limsup_{n\to\infty}G^{b}_{n}{}^{\circ}(\pi^{n+1};d_{n})\leq G^{\circ}(\pi^{\star};d).$
Taking $\limsup$ in the inequality above yields
$\langle\xi,d\rangle\leq G^{\circ}(\pi^{\star};d)\qquad\forall d,$
which is exactly the defining inequality for $\xi\in\partial_{C}G(\pi^{\star})$. ∎
Lemma E.10.
Let $\Pi=\times_{(k,s)}\Delta(A_{k})$. If $\pi_{m}\to\pi$ in $\Pi$ and $d\in T_{\Pi}(\pi)$, then there exist $d_{m}\in T_{\Pi}(\pi_{m})$ with $d_{m}\to d$.
Proof.
It suffices to argue blockwise for a simplex $\Delta_{m}$. Let $p\in\Delta_{m}$ with support $S_{+}:=\{i:p_{i}>0\}$. The tangent cone is $T_{\Delta_{m}}(p)=\{d:\mathbf{1}^{\top}d=0,\;d_{i}\geq 0\;\forall i\notin S_{+}\}$. Given $p_{m}\to p$ and $d\in T_{\Delta_{m}}(p)$, we can choose $t_{m}\downarrow 0$ and set
$d_{m,i}=\begin{cases}d_{i},&i\in S_{+},\\[2.0pt] \max\{d_{i},-p_{m,i}/t_{m}\},&i\notin S_{+},\end{cases}$
then shift the entries on $S_{+}$ by a small common offset so that $\mathbf{1}^{\top}d_{m}=0$. One checks that $d_{m}\in T_{\Delta_{m}}(p_{m})$ and $d_{m}\to d$. Taking the product over blocks gives the claim for $\Pi$. ∎
Theorem E.2.
Every cluster point $\pi^{\star}$ of $\{\pi^{n}\}$ satisfies the constrained Clarke stationarity condition
$0\in\partial_{C}G(\pi^{\star})+N_{\Pi}(\pi^{\star}).$ (124)
Proof.
Let $\pi^{n+1}\to\pi^{\star}$ and $\xi^{\,n+1}\to\xi$, $\lambda_{n}\to\lambda$, $y^{\,n+1}\to y\in N_{\Pi}(\pi^{\star})$ be as in Lemma E.9, so that $\xi\in\partial_{C}G(\pi^{\star})$ and $\xi+\lambda+y=0$.
Take an arbitrary $d\in T_{\Pi}(\pi^{\star})$. By Lemma E.10, we can choose $d_{n}\in T_{\Pi}(\pi^{n+1})$ with $d_{n}\to d$. Applying (122) to each $d_{n}$,
$\bigl\langle\xi^{\,n+1},d_{n}\bigr\rangle+\bigl\langle\lambda_{n},d_{n}\bigr\rangle\geq 0.$
Passing to the limit yields $\langle\xi+\lambda,d\rangle\geq 0$ for all $d\in T_{\Pi}(\pi^{\star})$, which means $\xi+\lambda\in N_{\Pi}(\pi^{\star})$ by polarity. Because $y\in N_{\Pi}(\pi^{\star})$ and $\xi+\lambda+y=0$, we deduce that 0 lies in the Minkowski sum $\partial_{C}G(\pi^{\star})+N_{\Pi}(\pi^{\star})$. ∎
We now show that any Clarke-stationary point with strictly positive gap cannot satisfy the normal cone condition, so cluster points must have zero gap and hence be robust NE.
Lemma E.11.
If $G(\bar{\pi})>0$, then
$0\notin\partial_{C}G(\bar{\pi})+N_{\Pi}(\bar{\pi}).$
Proof.
Suppose $G(\pi^{\star})>0$. Then there exists a block $b=(k,s)$ such that $G_{k,s}(\pi^{\star})>0$. By definition, $G_{k,s}(\pi^{\star})=\max_{a_{k}}\Omega^{k}_{R}(s,a_{k};\pi^{\star}_{-k})=:\delta>0$ (recall $\Omega$ sums to 0, so max must be positive if not all 0).
Let $a^{*}_{k}$ be an action achieving the maximum in the robust Bellman error. Let $e_{a^{*}}$ be the deterministic policy choosing $a^{*}_{k}$. Consider the direction $d\in T_{\Pi}(\pi^{\star})$ where the block $(k,s)$ component is $d_{b}=e_{a^{*}}-\pi^{\star}_{k}(\cdot|s)$ and 0 elsewhere.
The directional derivative of the gap function $G$ along $d$ involves terms related to $\Omega$. Specifically, moving probability mass towards the action with the highest advantages *decreases* the gap (since the gap measures optimality violation). Strictly speaking, the Clarke directional derivative $G^{\circ}(\pi^{\star};d)$ will satisfy:
$G^{\circ}(\pi^{\star};d)\leq-\delta<0.$
However, for stationarity $0\in\partial G+N$, we would require $\langle\xi,d\rangle\geq 0$ for all valid directions (via the normal cone polarity). Here we found a valid direction where the function decreases strictly. Thus $\pi^{\star}$ cannot be stationary. ∎
Theorem E.3 (Convergence to robust Nash equilibria).
Under Algorithm 2 with stepsizes satisfying $\sum_{n}\eta_{n}<\infty$, the sequence $\{G(\pi^{n})\}$ converges and every cluster point $\pi^{\star}$ of $\{\pi^{n}\}$ satisfies $G(\pi^{\star})=0$. Consequently, every cluster point is a robust Nash equilibrium. If the set of robust NE is finite, then the whole sequence $\{\pi^{n}\}$ converges to a robust NE.
Proof.
The convergence of $G(\pi^{n})$ follows from Proposition E.2. Let $\pi^{\star}$ be a cluster point. By Theorem E.2,
$0\in\partial_{C}G(\pi^{\star})+N_{\Pi}(\pi^{\star}).$
If $G(\pi^{\star})>0$, this contradicts Lemma E.11. Thus $G(\pi^{\star})=0$, and by Lemma E.3 $\pi^{\star}$ is a robust NE.
If the set of robust NE is finite, then $G(\pi)$ takes only finitely many values on that set. Since $G(\pi^{n})$ converges and every cluster point is a robust NE, cycling among distinct equilibria with different gap values is impossible, so the entire sequence $\{\pi^{n}\}$ must converge. ∎
E.3.3 Lipschitz
In this section, we derive some Lipschitz smoothness used in our proofs.
We first adopt the following lemma from (xu2025finite).
Lemma E.12.
(xu2025finite) There exists a seminorm $\|\cdot\|_{\mathcal{P}}$ on $\mathbb{R}^{|\mathcal{S}|}$ with kernel $\{c\mathbf{1}:c\in\mathbb{R}\}$ and a constant $\gamma\in(0,1)$ such that: for every fixed $\pi$ and $g\in\mathbb{R}$, the map
$T_{\pi,g}(v)(s)\triangleq\sum_{a}\pi(a\mid s)\Big(r(s,a)-g+\min_{q\in\mathcal{U}(s,a)}q^{\top}v\Big)$
is a $\gamma$‑contraction on the quotient space (mod constants).
On the finite‑dimensional quotient space $\mathbb{R}^{|\mathcal{S}|}/\mathrm{span}\{\mathbf{1}\}$ all norms are equivalent: there exist constants $C_{\infty\leftarrow\mathcal{P}},C_{\mathrm{sp}\leftarrow\mathcal{P}}>0$ such that
$\|x\|_{\infty}\ \leq\ C_{\infty\leftarrow\mathcal{P}}\|x\|_{\mathcal{P}},\qquad\|x\|_{\mathrm{sp}}\triangleq\max_{s}x(s)-\min_{s}x(s)\ \leq\ C_{\mathrm{sp}\leftarrow\mathcal{P}}\|x\|_{\mathcal{P}}.$ (125)
Lemma E.13 (Uniform bound on evaluation biases).
There exists some constant $B_{V}>0$ such that $\|V_{k}^{\pi}\|_{\infty}\leq B_{V}$ for all $k$ and $\pi$ (anchored at $s_{\circ}$).
Proof.
The proof directly follows from (wang2023robust), by noting the continuity of $V_{k}^{\pi}$ and compactness of $\mathcal{P}$. ∎
We prove that both the anchored bias $V^{\pi}$ and the average reward $g^{\pi}$ depend Lipschitz-continuously on the policy profile $\pi$. Throughout, $s_{\circ}\in S$ denotes the anchor state, and $\|\cdot\|_{\mathcal{P}}$ is the semi-norm from Lemma˜E.12; let $C_{\infty\leftarrow\mathcal{P}}>0$ be such that $\|x\|_{\infty}\leq C_{\infty\leftarrow\mathcal{P}}\|x\|_{\mathcal{P}}$ for all zero-mean $x$ (norm equivalence on the quotient space).
Lemma E.14 (Uniform bounds on $V^{\pi}$ and $g^{\pi}$).
Let $R_{\max}\triangleq\max_{k,s,a}|r_{k}(s,a)|$. Then for any player $k$ and policy $\pi$,
$\|V_{k}^{\pi}\|_{\infty}\leq B_{V}\quad\text{and}\quad|g_{k}^{\pi}|\leq R_{\max},$
where $B_{V}$ depends only on $(R_{\max},\gamma,C_{\infty\leftarrow\mathcal{P}})$ and the construction of $\|\cdot\|_{\mathcal{P}}$, but not on $\pi$.
Proof.
The bound on $V_{k}^{\pi}$ follows from the contraction in Lemma˜E.12 exactly as in Lemma E.13 (anchored at $s_{\circ}$): write $V_{k}^{\pi}=T_{\pi,g_{k}^{\pi}}(V_{k}^{\pi})$, subtract $T_{\pi,g_{k}^{\pi}}(0)$, and use the contraction plus norm-equivalence. For the scalar, by definition of average reward (and since $r_{k}$ depends only on current $(s,a)$), $g_{k}^{\pi}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[r_{k}(S_{t},A_{t})]\in[-R_{\max},R_{\max}].$
Lemma E.15 (Product-variation bound for joint actions).
Fix a state $s$. For two profiles $\pi,\tilde{\pi}$, their joint action distributions at state $s$ satisfy
$\sum_{a\in A(s)}\Big|\ \times_{j=1}^{K}\tilde{\pi}_{j}(a_{j}\!\mid s)\ -\ \times_{j=1}^{K}\pi_{j}(a_{j}\!\mid s)\ \Big|\ \leq\ \sum_{j=1}^{K}\|\tilde{\pi}_{j}(\cdot\!\mid s)-\pi_{j}(\cdot\!\mid s)\|_{1}.$
Proof.
Use the telescoping identity for products: $\times_{j}x_{j}-\times_{j}y_{j}=\sum_{m=1}^{K}(x_{m}-y_{m})\times_{j<m}y_{j}\times_{j>m}x_{j}.$ Apply triangle inequality, then sum over $a$ and use that for any fixed $m$, the factors are nonnegative and sum to at most $1$ when marginalizing over $a$. ∎
Lemma E.16 (Operator perturbation w.r.t. the policy).
Fix $k$, and let $B_{V}$ be as in Lemma E.14. For any $v\in\mathbb{R}^{|\mathcal{S}|}$ with $\|v\|_{\infty}\leq B_{V}$ and any scalar $g$ with $|g|\leq R_{\max}$, define the (player-$k$) robust one-step return $F_{k}(s,a;v)=r_{k}(s,a)+\min_{q\in\mathcal{U}(s,a)}q^{\top}v$. Then for any two policies $\pi,\tilde{\pi}$,
$\big\|T_{\tilde{\pi},g}(v)-T_{\pi,g}(v)\big\|_{\infty}\ \leq\ C_{\Pi}\,\|\tilde{\pi}-\pi\|_{2},$
for a constant $C_{\Pi}$ that depends only on $(K,S,A_{\max},R_{\max},B_{V})$, where $A_{\max}\triangleq\max_{k,s}|\mathcal{A}_{k}|$.
Proof.
For each state $s$,
$\big|T_{\tilde{\pi},g}(v)(s)-T_{\pi,g}(v)(s)\big|=\Big|\sum_{a}\big(\tilde{\pi}(a\mid s)-\pi(a\mid s)\big)\big(r_{k}(s,a)-g+\min_{q}q^{\top}v\big)\Big|.$
Since $|r_{k}-g+\min_{q}q^{\top}v|\leq R_{\max}+R_{\max}+\|v\|_{\infty}\leq 2R_{\max}+B_{V}$, we get
$\big|T_{\tilde{\pi},g}(v)(s)-T_{\pi,g}(v)(s)\big|\leq(2R_{\max}+B_{V})\sum_{a}|\tilde{\pi}(a\mid s)-\pi(a\mid s)|.$
Apply Lemma E.15 to bound the last sum by $\sum_{j}\|\tilde{\pi}_{j}(\cdot\mid s)-\pi_{j}(\cdot\mid s)\|_{1}$, then sum over $s$ and use Cauchy–Schwarz to convert the blockwise $\ell_{1}$ sum into the global $\ell_{2}$ norm, absorbing dimension constants in $C_{\Pi}$. ∎
Theorem E.4 (Lipschitz continuity of the evaluation map).
Under Lemma˜E.12, for each player $k$ there exist constants $L_{V},L_{\rho}>0$ depending only on the problem bounds and the contraction factor $\gamma\in(0,1)$ such that, for all policy profiles $\pi,\tilde{\pi}$,
$\|V_{k}^{\tilde{\pi}}-V_{k}^{\pi}\|_{\mathcal{P}}\ \leq\ \frac{L_{V}}{1-\gamma}\,\|\tilde{\pi}-\pi\|_{2},\qquad|g_{k}^{\tilde{\pi}}-g_{k}^{\pi}|\ \leq\ L_{\rho}\,\|\tilde{\pi}-\pi\|_{2}.$
Proof.
[…]
$\Delta v\ =\ \underbrace{T_{\tilde{\pi},g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})}_{\text{policy perturbation}}\ +\ \underbrace{T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\pi})}_{\text{contraction on $v$}}.$
Taking $\|\cdot\|_{\mathcal{P}}$ and applying Lemma˜E.12 to the second difference gives
$\|\Delta v\|_{\mathcal{P}}\ \leq\ \big\|T_{\tilde{\pi},g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})\big\|_{\mathcal{P}}\ +\ \gamma\,\|\Delta v\|_{\mathcal{P}}.$
Rearrange: $(1-\gamma)\|\Delta v\|_{\mathcal{P}}\ \leq\ \|T_{\tilde{\pi},g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})\|_{\mathcal{P}}.$ By norm equivalence and Lemma E.16 (applied with $v=V_{k}^{\tilde{\pi}}$ and $g=g_{k}^{\tilde{\pi}}$, which satisfy the uniform bounds of Lemma E.14), there is a constant $L_{V}>0$ with
$\|T_{\tilde{\pi},g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})\|_{\mathcal{P}}\ \leq\ C_{\mathcal{P}\leftarrow\infty}\,\|T_{\tilde{\pi},g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})-T_{\pi,g_{k}^{\tilde{\pi}}}(V_{k}^{\tilde{\pi}})\|_{\infty}\ \leq\ L_{V}\,\|\tilde{\pi}-\pi\|_{2}.$
Hence $\|V_{k}^{\tilde{\pi}}-V_{k}^{\pi}\|_{\mathcal{P}}\leq\frac{L_{V}}{1-\gamma}\,\|\tilde{\pi}-\pi\|_{2}.$
Use the anchored identity at $s_{\circ}$:
$0=V_{k}^{\pi}(s_{\circ})=\sum_{a}\pi(a\mid s_{\circ})\Big(r_{k}(s_{\circ},a)-g_{k}^{\pi}+\min_{q}q^{\top}V_{k}^{\pi}\Big),$
which rearranges to $g_{k}^{\pi}=\sum_{a}\pi(a\mid s_{\circ})\big(r_{k}(s_{\circ},a)+\min_{q}q^{\top}V_{k}^{\pi}\big).$ Therefore
$g_{k}^{\tilde{\pi}}-g_{k}^{\pi}=\underbrace{\sum_{a}\big(\tilde{\pi}-\pi\big)(a\mid s_{\circ})\big(r_{k}(s_{\circ},a)+\min_{q}q^{\top}V_{k}^{\tilde{\pi}}\big)}_{\text{change in policy at }s_{\circ}}\ +\ \underbrace{\sum_{a}\pi(a\mid s_{\circ})\Big(\min_{q}q^{\top}(V_{k}^{\tilde{\pi}}-V_{k}^{\pi})\Big)}_{\text{change in $v$}}.$
The first term is bounded by $(R_{\max}+B_{V})\sum_{a}|\tilde{\pi}(a\mid s_{\circ})-\pi(a\mid s_{\circ})|$, which by Lemma E.15 is at most $(R_{\max}+B_{V})\sum_{j}\|\tilde{\pi}_{j}(\cdot\mid s_{\circ})-\pi_{j}(\cdot\mid s_{\circ})\|_{1}$. The second term is bounded by $\|V_{k}^{\tilde{\pi}}-V_{k}^{\pi}\|_{\infty}\leq C_{\infty\leftarrow\mathcal{P}}\|V_{k}^{\tilde{\pi}}-V_{k}^{\pi}\|_{\mathcal{P}}$. Combine with Step 1 and Cauchy–Schwarz to obtain
$|g_{k}^{\tilde{\pi}}-g_{k}^{\pi}|\ \leq\ L_{\rho}\,\|\tilde{\pi}-\pi\|_{2},$
for some $L_{\rho}>0$ depending only on $(K,S,A_{\max},R_{\max},B_{V},\gamma,C_{\infty\leftarrow\mathcal{P}})$. ∎
E.3.4 Technical Lemmas
Lemma E.17 (Deterministic Robbins–Siegmund lemma (robbins1971convergence)).
Let $\{x_{n}\}$ be a real sequence bounded below and let $\{y_{n}\}$, $\{z_{n}\}$ be nonnegative sequences with $\sum_{n=0}^{\infty}z_{n}<\infty$. If
$x_{n+1}\ \leq\ x_{n}\ -\ y_{n}\ +\ z_{n}\qquad\text{for all }n,$
then $\{x_{n}\}$ converges (to a finite limit) and $\sum_{n=0}^{\infty}y_{n}<\infty$.
Lemma E.18.
Let $L$ be a finite index set, $\{a_{n,\ell}\}\in\mathbb{R}^{d}$ and $\{c_{n,\ell}\}\in\mathbb{R}$. Define the convex piecewise-linear function
$f_{n}(x)\triangleq\max_{\ell\in L}\{\langle a_{n,\ell},x\rangle+c_{n,\ell}\}.$
Assume $a_{n,\ell}\to a_{\ell}$ and $c_{n,\ell}\to c_{\ell}$, and define that
$f(x)\triangleq\max_{\ell\in L}\{\langle a_{\ell},x\rangle+c_{\ell}\}.$
Then, for $x_{n}\to x$ and $d_{n}\to d$, it holds that
$\limsup_{n\to\infty}\ f_{n}^{\circ}(x_{n};\,d_{n})\ \leq\ f^{\circ}(x;\,d),$
where
$f^{\circ}(x;d)\triangleq\limsup_{y\to x,t\downarrow 0}\frac{f(y+td)-f(y)}{t}$
is the Clarke directional derivative.
Proof.
For a function $g(x)=\max_{\ell\in L}\{\langle b_{\ell},x\rangle+\gamma_{\ell}\}$ with $L$ finite, and any $y,d$ and $t>0$, we have that
$g(y+td)=\max_{\ell\in L}\{\langle b_{\ell},y+td\rangle+\gamma_{\ell}\}=\max_{\ell\in L}\{\langle b_{\ell},y\rangle+\gamma_{\ell}+t\langle b_{\ell},d\rangle\}.$
Subtract $g(y)=\max_{\ell\in L}\{\langle b_{\ell},y\rangle+\gamma_{\ell}\}$ and we have that
$g(y+td)-g(y)$ $=\max_{\ell\in L}\{\langle b_{\ell},y\rangle+\gamma_{\ell}+t\langle b_{\ell},d\rangle\}-g(y)\}.$
Divide by $t>0$ and we get
$\frac{g(y+td)-g(y)}{t}=\max_{\ell\in L}\left\{\langle b_{\ell},d\rangle+\frac{\langle b_{\ell},y\rangle+\gamma_{\ell}-g(y)}{t}\right\}.$
Now let $A_{g}(y)\triangleq\arg\max_{\ell\in L}\{\langle b_{\ell},y\rangle+\gamma_{\ell}\}$. If $\ell\notin A_{g}(y)$, then the second term equals a negative constant divided by $t$, hence tends to -$\infty$ as $t\downarrow 0$. Therefore
$g^{\circ}(y;d)\ =\ \limsup_{t\downarrow 0}\frac{g(y+td)-g(y)}{t}\ =\ \max_{\ell\in A_{g}(y)}\ \langle b_{\ell},d\rangle.$ (126)
Now we define
$A_{n}(x_{n})\triangleq\arg\max_{\ell\in L}\{\langle a_{n,\ell},x_{n}\rangle+c_{n,\ell}\}.$
Then by (126),
$f_{n}^{\circ}(x_{n};d_{n})\ =\ \max_{\ell\in A_{n}(x_{n})}\ \langle a_{n,\ell},d_{n}\rangle.$ (127)
Pick, for each $n$, an index $\ell_{n}\in A_{n}(x_{n})$ such that
$f_{n}^{\circ}(x_{n};d_{n})\ =\ \langle a_{n,\ell_{n}},d_{n}\rangle.$
Because $L$ is finite, there exists a some subsequence $\{n_{t}\}$, such that $\ell_{n_{t}}\equiv\bar{\ell}$ is a constant. Then
$\lim_{t\to\infty}\ \langle a_{n_{t},\ell_{n_{t}}},d_{n_{t}}\rangle=\lim_{t\to\infty}\ \langle a_{n_{t},\bar{\ell}},d_{n_{t}}\rangle=\ \langle a_{\bar{\ell}},d\rangle,$ (128)
due to convergence of $a_{n,\ell}$ and $d_{n}$.
Note that the LHS of (128) is
$\lim_{t\to\infty}\ \langle a_{n_{t},\ell_{n_{t}}},d_{n_{t}}\rangle=\lim_{t\to\infty}f^{\circ}_{n_{t}}(x_{n_{t}};d_{n_{t}})=\limsup_{n\to\infty}f^{\circ}_{n}(x_{n};d_{n}),$ (129)
and thus $\limsup_{n\to\infty}f^{\circ}_{n}(x_{n};d_{n}),=\langle a_{\bar{\ell}},d\rangle$.
Define
$u_{n,\ell}\triangleq\langle a_{n,\ell},x_{n}\rangle+c_{n,\ell},\qquad u_{\ell}\triangleq\langle a_{\ell},x\rangle+c_{\ell}.$
Since $a_{n,\ell}\to a_{\ell}$, $c_{n,\ell}\to c_{\ell}$, and $x_{n}\to x$, we have $u_{n,\ell}\to u_{\ell}$ for all $\ell$. Since $L$ is a finite set, it holds that
$f_{n}(x_{n})=\max_{\ell}u_{n,\ell}\ \to\ \max_{\ell}u_{\ell}\ =\ f(x).$
Since $\ell_{n}\in A_{n}(x_{n})$, we have $u_{n,\ell_{n}}=f_{n}(x_{n})\to f(x)$; taking the limit along the constant subsequence $\ell_{n_{t}}\equiv\bar{\ell}$ implies that
$u_{n_{t},\ell_{n_{t}}}=f_{n_{t}}(x_{n_{t}})\to_{t\to\infty}f(x).$ (130)
On the other hand, $u_{n_{t},\ell_{n_{t}}}\to u_{\bar{\ell}}$m thus $u_{\bar{\ell}}=f(x)$, i.e., $\bar{\ell}\in A(x)\triangleq\arg\max_{\ell}\{\langle a_{\ell},x\rangle+c_{\ell}\}$.
Finally, using (128) and $\bar{\ell}\in A(x)$,
$\limsup_{n\to\infty}\ f_{n}^{\circ}(x_{n};d_{n})\ =\ \langle a_{\bar{\ell}},d\rangle\ \leq\ \max_{\ell\in A(x)}\ \langle a_{\ell},d\rangle\ =\ f^{\circ}(x;d).$
This hence completes the proof. ∎
Appendix F Proof of Section 7
Theorem F.1.
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted DR-MG by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward DR-MG.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted DR-MG is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
(3). The discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2)
Proof.
(1). By the uniform convergence of the robust discounted value function derived in (wang2023robust), we have that
$\lim_{\gamma\to 1}(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}=g^{\pi}_{\mathcal{P},i},\text{ uniformly on }\Pi.$ (131)
Thus for any $\epsilon$, there exists some $\gamma_{e}$, such that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,$ (132)
for any $\pi\in\Pi$ and $\gamma>\gamma_{e}$. Moreover, there exists some constant $T$ such that for any $t>T$,
$\gamma_{t}>\gamma_{e},\|\pi_{t}-\pi\|\leq\epsilon.$ (133)
Since $\pi_{t}$ is the Nash Equilibrium of the discounted robust MG, it holds that
$V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}\geq V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i},\forall i\in\mathcal{N},\forall\mu^{i}.$ (134)
Thus we have that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$
$=-g^{\pi}_{\mathcal{P},i}+(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}$
$\quad-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}+g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}$. (135)
Consider any $t>T$, it first holds that due to (132):
$(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}$ $\leq\epsilon,$ (136)
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}$ $\leq\epsilon.$ (137)
Moreover,
$(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{\pi}_{\gamma_{t},\mathcal{P},i}\leq(1-\gamma_{t})(V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}-V^{\pi}_{\gamma_{t},\mathcal{P},i})\leq\epsilon,$ (138)
$(1-\gamma_{t})V^{(\pi^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}-(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq\epsilon,$ (139)
which is due to the $\frac{1}{1-\gamma}$-Lipschitz of the robust discounted value function in $\pi$, and the closeness of $\pi_{t}$ and $\pi$ in (133). Finally,
$-(1-\gamma_{t})V^{\pi_{t}}_{\gamma_{t},\mathcal{P},i}+(1-\gamma_{t})V^{(\pi_{t}^{-i},\mu^{i})}_{\gamma_{t},\mathcal{P},i}\leq 0$ (140)
due to (134). Combining all results hence implies that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq\epsilon,$ (141)
for any $i$ and $\mu^{i}$ uniformly. Thus let $\epsilon\to 0$ we complete the proof.
(2). Set $\gamma$ to be any factor that is larger than $\gamma_{e}$. Then
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}=g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}+(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}.$ (142)
Note that for $\gamma>\gamma_{e}$, it holds that
$g^{(\pi^{-i},\mu^{i})}_{\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\leq 2\epsilon+(1-\gamma)V^{(\pi^{-i},\mu^{i})}_{\gamma,\mathcal{P},i}-(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}\leq 3\epsilon,$ (143)
due to $\pi$ is an $\epsilon$-NE for the discounted robust MG, which hence completes the proof.
(3). The proof is a direct corollary of Theorem 7.1 and results from (Roch2025reduction). Specifically, it is shown in Theorem 3.4 in (Roch2025reduction) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^3] The results in (Roch2025reduction) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ (Roch2025reduction)., it holds that
$\|(1-\gamma)V^{\pi}_{\gamma,\mathcal{P},i}-g^{\pi}_{\mathcal{P},i}\|\leq\epsilon,\forall\pi,\forall i.$ (144)
Combining with Theorem 7.1 further completes the proof. ∎